A Complete Bibliography of Publications in Natural Resource Modeling

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.14

Title word cross-reference

1 [880]. 2 [880]. 3 [900]. _{2} [848, 918]. _{2.5} [961]. C^{1} [297]. L [860]. N [764]. R_{0} [769]. θ [932].

-based [1029]. -logistic [932]. -moments [860]. -patch [764].

1997 [335].

2011 [643]. 21st [441].

3.0 [219]. 386i [227].

4.0 [207].

abalone [685, 832]. abatement [604]. absorb [848]. abundance [432, 651, 702, 138]. Acarapis [194]. access [1023, 486, 267, 465]. account [566]. Accounting [537, 889, 925, 625]. accuracy [458, 187, 188]. accurate [619]. achieve [887]. acid [657]. acidification [974]. across [336, 861]. Act [113]. action [608]. actions [855]. activities [698, 802, 1050, 462, 916]. activity [412]. actual [993]. adaptation [597, 646]. adaptations [1027]. adaption [702]. Adaptive [20, 600, 615, 938, 549, 574, 734]. adjustment [19, 777]. adoption [687, 647]. adult [639]. adulticide [837]. advantage [708]. advection [606]. adventure [437]. aerial [880]. affect [1030, 787]. affecting [819]. affects [501]. Africa [720, 678]. African [344, 342]. after [666, 506]. again [440]. against [720, 25]. Age [177, 116, 1052, 878, 141, 67, 766, 483, 276, 321, 402, 388, 194, 441, 770, 513, 511, 560, 726]. age- [766, 321]. age-class [513]. age-dependent [67]. age-specific [513]. age-structured [116, 1052, 878, 141, 483, 770, 560, 726]. aged [872, 576]. agent [397, 27, 937, 951, 965, 493, 817]. agent-based [397, 937, 951, 965, 493, 817]. ages [439, 818]. agglomeration [1051]. aggregate [803]. aggregated [369]. aggregation [448]. agreement [425]. agreements [705, 431, 496, 815, 664]. agricultural [866, 595, 712, 457, 920]. agriculturally [751]. agriculture [996, 785, 646, 582, 667, 691, 299, 813, 403, 692]. agriculture-industry [299]. Agroecological [794]. AHP [805]. AICME [364, 373]. AIDS [984]. air [584, 590]. airborne [852]. Ajora [1014]. Ajora-Woybo [1014]. Alaska [754, 965]. Alaskan [570]. Alberta [291]. Alcalá [364]. Alcalæ [373]. alfalfa [780, 820]. Algeria [944]. algorithm [651]. Algorithmic [813]. algorithms [728, 390]. Allee [725, 95, 793, 843, 762, 989]. allelochemical [972]. Alligator [822]. Allocation [833, 840, 1024, 183, 476, 867, 115, 362, 279]. allocations [600, 268]. Allocative [36]. allometric [546, 852]. allowable [280]. allowance [716]. allowing [648]. Alluvial [982]. almost [506]. along [1038]. altered [844]. Alternative [533, 685, 809, 651, 287, 471, 740, 1008, 697]. alternatives [174]. Amazon [962, 919]. ambiguity [34]. amenity [535]. America [486]. American [660, 749]. americana [472]. amphibian [495]. analyses [980, 782]. Analysis [244, 303, 718, 861, 846, 170, 409, 898, 991, 596, 27, 760, 599, 475, 547, 1029, 608, 301, 819, 592, 183, 628, 331, 26, 459, 942, 440, 860, 700, 994, 35, 343, 273, 398, 255, 150, 531, 310, 754, 657, 601, 455, 574, 392, 947, 282, 783, 1009]. analytic [583, 732]. Analytical [868, 248, 606]. analyze [533]. Analyzing [118, 395, 830, 386, 472, 928]. anchovy [911, 670, 376]. Andalusia [847]. animal [690, 612]. annual [749]. ant [552]. anthrax [842]. anthropogenic [903, 1032]. anticipate [210]. antimicrobial [642]. Any [405]. Apis [194, 303]. app [782]. Appalachians [225]. apparent [740]. applicability [1039]. Application [921, 941, 549, 302, 568, 175, 1031, 627, 681, 399, 263, 293, 683, 355, 1045, 589, 564, 613, 499, 981, 1053, 442]. applications [297, 495, 559, 349, 354, 383]. applied [674, 658, 368, 255, 146]. Applying [897]. approach [360, 75, 396, 705, 1044, 875, 650, 517, 372, 140, 507, 1022, 163, 634, 1024, 261, 503, 16, 173, 504, 183, 124, 312, 461, 902, 35, 46, 277, 607, 371, 426, 830, 506, 386, 524, 381, 309, 382, 709, 509, 403, 404, 561, 378, 723, 493, 781]. approaches [850, 470, 368, 264, 248]. Approximately [1036]. approximation [187, 188, 427]. AquaCrop [818]. aquaculture [862, 991, 783]. aquatic [629, 681]. Aquifer [982, 940, 920]. Araguaia [919]. aratio [538]. aratio-dependent [538]. arching [819]. Arctic [514]. Ardennes [302]. area [413, 566, 899, 414, 897]. areas [596, 410, 411, 336, 122, 610, 664, 406, 1026]. Argentine [552]. arid [898, 1024, 940, 512, 641, 1002, 1026]. arising [38, 868]. arrangements [1011]. Array [158]. Artificial [585, 890, 906, 400]. ash [950]. Asia [26]. aspects [347]. assess [939, 794]. Assessing [812, 651, 726, 522, 328, 187, 613]. Assessment [962, 917, 805, 1026, 878, 892, 411, 440, 450, 726]. assumption [875]. assumptions [726]. asymmetric [545, 371]. Asymptotic [355]. Atlantic [708, 166, 150]. atmosphere [590]. atmospheric [963]. attract [850]. attraction [989]. attributes [404]. audit [139]. augmenting [222]. Australia [520, 507, 832, 512, 614]. Australian [149]. Austria [951]. AUTODIF [158]. automata [816]. automated [592]. automaton [781]. availability [650, 681, 695]. average [722]. averse [647, 992, 20]. aversion [599, 555, 511]. avian [695, 837]. avoidance [698]. awareness [557]. axiomatic [503]. axisymmetric [356].

Back [892, 440]. bacteria [700]. baited [761]. balance [108]. Balancing [988]. Baltic [1007, 702, 566, 524, 668, 995]. Barau [602]. Barents [523, 729]. bargaining [387]. Barred [796]. Barrier [510, 509, 939]. barriers [881]. based [397, 1012, 898, 396, 892, 530, 541, 323, 593, 937, 1029, 540, 821, 890, 328, 651, 308, 504, 542, 1039, 401, 393, 636, 767, 285, 906, 961, 718, 900, 863, 332, 386, 398, 394, 717, 951, 676, 916, 583, 732, 1025, 965, 493, 783, 722, 781, 816, 817]. Basin [1038, 1004, 326, 1029, 1014, 627, 860, 404]. basins [713, 989]. Bay [645, 271, 670, 615]. Baye [288]. Bayesian [78, 58, 824, 440, 979, 984, 847, 549, 954]. be [29, 607, 425]. beach [992]. bears [477]. beaver [176]. Beddington [379]. bee [303, 816]. been [425]. bees [194]. beetle [954]. beetles [559]. behaved [613]. behavior [204, 865, 875, 205, 656, 542, 508, 768, 9, 975, 448, 416, 117, 704, 454, 306, 916, 955]. Beijiang [855]. Beijing [994]. beliefs [800]. benefit [331, 678, 255]. Benefits [605, 921, 348, 312, 550, 8, 256, 416, 509]. Bengal [653]. Bering [515, 246]. Bertalanffy [383]. best [206, 7]. better [964]. Betula [900]. between [944, 866, 983, 1017, 5, 819, 812, 780, 368, 993, 1045, 303, 833, 899, 546, 576, 783]. Beverton [630, 927]. Beyond [1052, 458]. biased [636]. bifurcation [581, 353]. bifurcations [299]. big [665, 834]. bilevel [277]. bio [466, 783]. bio-economic [466, 783]. biocapacity [1030]. biodiversity [519, 480, 597, 866, 431, 773, 980, 1049, 855, 844, 594]. Bioeconomic [460, 668, 409, 413, 909, 596, 543, 733, 830, 389, 841, 601, 561, 671]. bioeconomics [633, 739, 377]. bioenergy [647, 578]. biogeography [480]. Biological [552, 725, 14, 793, 564, 528, 94, 1053]. biology [375, 303]. biomass [918, 621, 634, 647, 900, 560, 583, 732, 852]. biopolitical [463]. bioreactor [263]. biotechnological [656]. birch [900]. birds [323, 120]. birth [267]. Biscay [670]. black [913, 749, 477, 911]. black-tailed [913]. block [361, 765]. block-rate [361]. blowing [593]. blue [488]. blue-winged [488]. bluefin [149]. body [983]. bole [585]. bole-volume [585]. Book [199, 61, 192, 112, 142, 221, 53, 79, 101, 11, 125, 91, 62, 133, 80, 152, 102, 171, 159, 100, 247, 10, 31, 160, 144, 252, 111, 126, 191, 70, 265, 89, 119, 143, 153, 71, 200, 208, 134, 213, 239, 90, 21, 178, 238, 258, 135, 42, 110, 220, 228, 161, 43, 179]. booming [391]. bootstrap [275]. borer [950]. bottom [395]. bottom-up [395]. boundaries [801, 363]. Bounded [827, 856]. Brazil [919]. BRI [1033]. bridge [859]. bridging [960]. brief [725, 636, 84, 490]. brood [771]. Bt [479]. budget [840]. budworm [305, 363, 196]. buffer [124]. Bus [1037]. bushfires [834]. butterflies [784]. Buying [595]. bycatch [498].

C [158, 849]. c-Si [849]. calculating [811]. Calendar [12, 22, 32, 44, 54, 63, 72, 92, 103, 81]. calibrated [857]. Calibrating [787]. California [410]. call [660]. Campaign [302]. Can [796, 991, 964, 688, 881, 953]. Canada [494, 643, 671]. Cannibalism [975, 741, 771]. capability [848]. capacities [1025]. Capacity [274, 821, 1040, 863, 938]. capital [211, 889, 469, 691, 115, 217]. capture [624, 783]. capture-based [783]. Carbon [786, 939, 848, 716, 1046, 942, 256, 455, 852, 963]. Caribbean [809]. Carinthia [951]. Carlo [588, 660]. Carolina [271, 310]. carrying [821, 1040, 863, 1025]. Case [899, 291, 929, 1007, 909, 856, 713, 523, 570, 609, 621, 940, 16, 173, 922, 851, 805, 589, 860, 699, 994, 216, 533, 879, 510, 1037, 951, 960, 333, 799, 1002, 573]. cases [352]. Castellammare [460]. Catastrophic [1025, 455, 799]. Catch [856, 983, 621, 894, 50, 280, 18, 945]. catch-per-unit-effort [18]. catch-to-biomass [621]. Catch-to-stock [856]. catchability [18]. Catchment [508, 1024, 897]. catchments [600]. Catherine [807]. caused [593, 954]. cellular [781, 816, 219]. Central [751, 678]. century [441]. chain [996, 298, 379, 915, 973, 911, 891, 785, 581, 949, 574, 816]. chains [939, 432]. chalco [515]. chalco-gramma [515]. chalcogrammus [965]. challenge [73]. chance [95]. change [522, 1031, 597, 627, 892, 147, 214, 514, 521, 628, 644, 556, 699, 566, 524, 847, 754, 467, 671, 829, 816]. changed [118]. changes [878, 683, 901, 370, 391]. changing [885, 656, 679, 51]. Chaohu [1004]. Chaos [132]. characteristic [312]. characteristics [901, 181]. characterize [547]. Characterizing [616]. Chem [961]. China [875, 935, 905, 1051, 1018, 855, 993, 942, 900, 863, 879, 904, 1004, 876]. Chinese [224]. Chinook [611, 321, 404]. choice [576]. choices [716, 992]. Chromosome [910]. Cienfuegos [645]. cities [224]. class [38, 402, 513]. classes [511]. classical [563]. classification [906, 879]. Cleaning [608, 850]. cleanup [943, 550, 882]. Climate [522, 1031, 627, 467, 214, 118, 514, 521, 1015, 1040, 628, 527, 688, 644, 556, 901, 699, 847, 754, 671]. climatic [1032]. climax [117, 306]. close [665]. closed [122]. Co [738, 848, 918]. Co-evolutionary [738]. coal [186]. Coalition [501, 426]. Coalitions [554, 604, 445]. Coastal [27, 157, 260, 894, 940, 497, 527, 805]. coastline [805]. coasts [683]. cod [934, 708, 1007, 514, 566, 524, 668, 995]. coefficients [495, 261]. Coexistence [973, 891, 338, 740, 257]. coffee [589]. Coho [527]. cold [785, 345]. collapsed [670]. Collective [758]. Colony [771, 194, 816]. Colorado [954]. Columbia [404]. combat [302]. combination [1050]. combined [28]. Combining [40, 691, 68, 964]. Come [841]. comity [859]. commercial [1010, 851, 150]. commercial-scale [1010]. Commission [442]. commodities [251]. Common [907, 193, 307, 710, 249, 419, 294, 338, 531, 339]. common-pool [338]. commons [131]. communicating [395]. community [518, 844]. Comparative [918, 1039, 459]. Comparing [932, 679, 852, 952, 569, 196]. Comparison [50, 39, 910, 264, 880]. compartmental [810]. compatible [553, 185]. competing [525, 453, 306, 972]. Competition [591, 737, 359, 420, 706, 371, 740]. competitive [338]. complementarity [40]. Complex [915, 539, 505]. compliance [223]. Component [725]. components [570, 651]. composition [1022]. compounds [704]. Comprehending [886]. computable [754]. Computation [357]. computational [969]. Computer [198]. computers [184]. Computing [294, 268, 1042]. conceptual [1044, 941]. concerned [113]. concessions [788]. condition [941]. conditions [630, 901, 472, 995]. Conference [364, 373, 643, 335]. Confidence [824]. Conflict [307]. conflicting [157, 702]. conflicts [586]. conifer [1040]. conifers [732]. Connectivity [956, 953]. consensus [463]. consequences [412]. Conservation [602, 113, 466, 553, 982, 811, 139, 1017, 106, 431, 73, 775, 725, 980, 998, 502, 1049, 567, 822, 284, 1050, 855, 236, 594, 343, 853, 620, 887, 1011, 723]. conserve [991, 665, 235]. conserving [342]. Consideration [14]. considerations [801, 127, 394]. considering [996]. consistent [873]. conspiracy [51]. Constant [374, 306]. constraints [291, 804, 524]. construction [803, 392]. consumer [517, 748, 557]. consumers [210]. consumption [609, 731, 899]. contact [680]. contagion [613]. containment [952]. contamination [326, 981]. content [76]. context [412, 123, 602, 712]. contingency [952]. Contingent [410]. continuous [485, 425, 810]. continuous-time [485, 810]. continuously [487]. contract [249, 660, 1011]. contribution [597, 676]. contributions [497]. Control [477, 674, 757, 4, 40, 165, 326, 689, 474, 372, 764, 189, 725, 921, 148, 891, 913, 987, 971, 263, 308, 183, 478, 724, 795, 26, 337, 174, 370, 533, 933, 528, 575, 506, 498, 510, 224, 999, 146, 333, 378, 723, 963, 781]. Controlled [124, 376]. Controlling [329, 316, 881]. controls [272, 573]. convergence [368]. converse [358]. conversion [902, 829]. convert [829]. convolutional [906]. cooperation [307, 627, 591, 312, 51]. Cooperative [801, 843, 262, 315]. copula [898, 912]. copula-based [898]. coral [809, 864]. core [595]. cormorants [734]. correct [425]. corrective [309]. correlated [611]. Correlation [994]. Cost [1046, 255, 956, 567, 331, 216, 264, 250, 676, 269, 98]. Cost-benefit [255, 331]. cost-effective [567, 264]. Cost-effectiveness [1046]. costate [337, 492]. costly [19]. Costs [774, 348, 566, 964, 250, 777, 403, 186]. could [980]. counter [508]. counter-cyclical [508]. countries [705, 1007, 1030]. counts [695]. County [879]. coupled [903, 638]. Coupling [899]. cover [494, 917, 1038, 1014]. coverages [568]. cowbirds [399]. crabs [615]. creating [412]. Creation [594, 783]. credibility [941]. crested [734]. crimes [720]. crisp [918]. criteria [868, 589]. critical [95, 728]. crop [245, 780, 478]. cropping [671]. crops [476, 847, 945]. crowd [881]. Crowley [936]. Cuba [645]. culling [796]. cultivated [935]. cultural [603]. Cumulative [919]. Current [984]. curvature [454]. curve [897]. curved [683]. cut [374, 290]. cut-offs [374]. cutthroat [398]. cutting [164]. cycle [749, 138, 957]. cycle-jumping [957]. cyclical [508].

D [880, 900]. DADiSP [170]. daily [919, 676]. Dakota [902]. damage [799, 884]. damages [6, 450]. Data [169, 34, 568, 170, 962, 218, 275, 890, 559, 824, 50, 533, 941, 709, 834, 726, 175]. data-limited [824]. datasets [917]. dealing [695]. DeAngelis [379]. death [267]. debate [595]. debugging [395]. decay [557, 816]. decision [589, 673, 476, 332, 235, 960, 549, 758]. decision-making [332, 960]. decisions [444, 658, 148, 531, 671, 828]. declining [796]. decomposition [831, 942]. decreasing [267]. Dedication [807, 836]. deep [906]. deer [800]. defense [891]. deficit [476, 672]. defined [313]. definition [494]. degradation [1033, 593]. Degree [714]. Delaware [615]. delay [538, 810]. delayed [584, 38, 67]. delayed-recruitment [38]. delays [1023]. Delta [899]. DeLury [515]. demand [222, 130, 420, 803, 197, 696]. demersal [128]. demographic [1012, 868]. demography [14]. density [885, 47, 95, 49, 695, 458, 881, 305, 888]. density-dependent [885, 47]. dependence [856, 621, 700]. dependent [970, 885, 67, 530, 47, 746, 49, 802, 538, 305, 606, 625]. depleted [998]. Depletion [250, 515, 266, 452, 700, 620]. deposits [532, 274]. Derivation [613, 75]. Derivatives [666]. derived [256]. describing [402, 303]. deserts [860]. design [996, 585, 501, 14, 930]. Designing [437]. Desk [169]. destruction [466]. detail [948]. Details [539]. detect [512]. detection [648]. determinants [224]. Determination [60, 280, 897]. determine [4]. determined [536]. determining [937]. deterministicand [495]. deterrence [648, 380]. developed [705, 941]. developing [705, 1039]. Development [476, 751, 935, 442, 106, 266, 387, 823, 870, 84, 436, 999, 916, 315]. developments [368]. difference [303]. differences [1017]. different [918, 562, 652, 215, 849, 377, 818, 692]. differential [627, 69, 744, 331, 180].

Differentiated [993]. diffusing [176]. Diffusion [581, 797, 304, 353, 352, 222, 552, 305, 363]. Diffusion-driven [304]. diffusive [298, 379, 86]. dimensional [935, 660, 863, 305, 606]. dioxide [939, 848, 942, 963]. dipsaci [780]. disciplinary [232]. Discontinuous [297]. discount [563, 48]. discounting [422, 474, 550, 499, 428]. discovery [187, 188]. discrete [495, 29, 971, 724, 639, 747, 194, 482, 484, 268, 842, 989]. discrete-time [495, 724, 639, 484, 268, 842, 989]. discretization [865, 838]. Disease [929, 764, 760, 536, 753, 166, 989]. diseases [980, 881]. disequilibrium [129]. disinfectant [753]. dispersal [839, 488, 739]. dispersion [584, 590, 929, 606]. display [170]. dissolved [653, 700]. distant [27, 157]. distortions [658]. distributed [423, 183]. distribution [323, 1047, 917, 693, 348, 919, 532, 339, 246, 427]. distributional [596]. distributions [934, 276, 616]. disturbance [846, 832, 388, 844, 225, 828]. disturbances [872]. ditonicity [255]. Ditylenchus [780]. dive [832]. diversions [193]. diversity [530, 613]. diversity-dependent [530]. division [771]. Do [619, 787, 829]. documents [886]. Does [753, 362, 1030, 1051, 393, 560]. dog [913]. doing [465]. domain [379, 370, 378]. domains [353]. domestic [279]. dominant [585, 479]. don’t [800]. Dorée [929]. double [734]. double-crested [734]. douglas [260, 510]. douglas-fir [260]. downs [696]. drainage [148]. driven [397, 680, 304, 751]. drivers [1014]. Drought [582, 898, 909, 667, 691]. drought* [883]. drown [853]. drum [310]. ducks [749]. due [657]. during [912, 476]. Dynamic [193, 309, 648, 360, 705, 249, 140, 464, 821, 905, 308, 577, 294, 459, 273, 19, 531, 74, 181]. Dynamical [865, 916, 591]. Dynamics [848, 864, 483, 97, 728, 381, 737, 956, 885, 611, 517, 372, 748, 140, 325, 56, 85, 690, 485, 915, 630, 612, 432, 500, 507, 1029, 559, 736, 293, 624, 738, 324, 287, 569, 402, 1045, 653, 1038, 463, 680, 659, 345, 484, 436, 582, 771, 310, 450, 472, 732, 1014, 751, 928, 972]. dynamite [84].

early [441]. Eastern [515, 940, 566, 524, 668, 692, 995]. eco [736]. eco-evolutionary [736]. Ecological [791, 520, 841, 799, 1044, 536, 595, 792, 636, 826, 863, 1009, 1032, 638, 951, 1030, 561]. Ecological-economic [520]. ecologically [418, 800, 804]. Ecology [364, 373, 395, 375, 756, 863]. econometric [905]. Economic [956, 523, 656, 642, 9, 406, 522, 827, 982, 412, 520, 921, 536, 621, 59, 521, 502, 812, 348, 497, 683, 857, 714, 46, 754, 949, 466, 887, 299, 455, 783, 870]. Economically [730, 680]. Economics [637, 407, 702, 77, 564, 273, 786, 672, 248]. economies [1033, 516]. economy [165, 943, 675, 994, 863, 314]. ecosystem [519, 597, 147, 909, 748, 773, 774, 980, 624, 1050, 548, 15, 581, 759, 887, 697]. ecosystems [400, 831, 746, 410, 267, 446, 1008, 516]. Edge [320]. Editor [99]. Editorial [565, 162, 967, 924, 13, 45, 489, 978, 23, 456, 505, 341, 434, 473, 481, 33, 1, 55, 65, 145, 1006, 167, 947]. educators [969]. Eel [182]. Effect [650, 205, 963, 291, 903, 710, 423, 452, 86, 656, 819, 556, 459, 462, 267, 77, 833, 314, 620, 657, 762, 671, 1004, 989, 914, 625]. effective [567, 264, 981]. effectiveness [1046]. Effects [379, 839, 322, 848, 584, 590, 360, 222, 530, 541, 725, 95, 918, 640, 612, 562, 618, 832, 793, 514, 521, 49, 843, 542, 767, 993, 662, 433, 250, 557, 421, 726, 649, 1002, 972, 1043]. efficacy [968]. Efficiency [510, 36, 365, 215, 1051, 822, 471]. Efficient [279, 155, 687, 69, 688, 60, 115]. Effort [380, 840, 271, 374, 856, 1047, 50, 18]. egg [975]. egg-laying [975]. eigenvalues [868]. elasticity [211, 130, 882]. Eldana [817]. electric [942]. element [553]. elephants [347, 720, 346]. elk [932]. elliptic [356, 722]. emerald [950]. emerging [647]. Emission [156, 865, 607]. emission-reduction [865]. emissions [588, 942, 704]. Empirical [628, 280, 703, 991, 792, 150]. employment [904]. encompassing [140]. Endangered [257, 520, 602, 853, 549]. endemic [602]. endogenous [271, 329, 804]. energetic [676]. Energy [139, 716, 905, 996, 808, 865, 994, 197]. energy-saving [865]. Energy-use [716]. enforcement [365, 698, 114]. engineering [394]. England [692]. enhancement [855]. enhancing [344]. enrichment [668]. ensemble [733]. entire [917]. Entry [469, 271, 236, 289, 380, 408, 282]. environment [116, 468, 703, 885, 374, 1036, 679, 994, 345, 867, 638, 493]. Environmental [292, 165, 597, 6, 609, 223, 1003, 449, 846, 1033, 553, 203, 903, 705, 185, 501, 496, 365, 502, 545, 712, 497, 992, 542, 453, 313, 925, 815, 566, 964, 524, 224, 557, 949, 704, 472, 299, 876]. environments [430, 950, 918, 327, 48]. enzootiology [166]. epidemic [495, 769]. epidemics [447]. epipelagic [1010]. equation [630, 744, 180, 552, 305, 363, 427, 606]. equations [38, 303, 356, 383]. equilibria [262, 69, 294, 314]. Equilibrium [138, 121, 862, 905, 129, 754, 1019, 697, 351, 98]. equity [387, 887]. Equivalencies [663]. eradication [154, 910]. ergodic [85]. erosion [593, 1029, 805, 1038, 902]. Errata [30, 41, 52, 82, 93, 201, 253, 443, 83]. Erratum [895, 958, 1020, 534]. error [1043]. escapement [289]. established [460]. estimate [676]. estimates [681, 651, 569, 619, 471, 1043]. Estimating [293, 322, 622, 728, 932, 902]. Estimation [289, 803, 175, 585, 458, 50, 8, 583]. estimator [515]. estuarine [653]. Ethiopia [1029, 1038, 1014]. EU [788]. Eurasia [852]. eutrophication [791, 974]. Evaluating [685, 983, 795, 855, 1045, 678, 880, 704, 288, 847]. Evaluation [849, 413, 761, 656, 280, 1003, 509, 751]. evapotranspiration [861]. even [872, 576]. even- [576]. even-aged [872]. event [397]. events [12, 22, 32, 44, 54, 63, 72, 92, 103, 81]. Evidence [903, 1018, 703, 24, 1030]. evolution [970, 635, 493]. Evolutionarily [740]. evolutionary [737, 682, 741, 518, 736, 738, 771]. evolve [1051]. evolves [739]. exact [838]. examine [518]. examining [525]. example [260, 513]. excessive [823]. exchange [352]. exclosures [649]. exclusion [460, 338]. exertion [272]. Exhaustible [7, 359, 36, 108, 675, 425, 499]. Existence [367, 567]. exit [469]. Exogenous [623]. exotic [520, 979]. expectations [420, 19, 37, 87]. expected [280, 288]. experiences [1030]. explicit [572, 700]. Exploitation [339, 57, 682, 611, 419, 873, 651, 223, 381, 257, 762]. exploited [617, 639, 377]. exploration [658, 475]. Exploring [123]. expressions [868]. extended [27]. extending [592]. extensification [595]. Extension [158]. extent [880]. externalities [203, 195, 974, 109, 426, 759, 309]. externality [571]. extinct [937]. extinction [330, 486, 630, 95, 495, 868, 321, 257]. extraction [658, 1015, 17, 874, 98]. extractions [459]. extreme [990, 912].

face [1052, 799]. facing [892]. factors [903, 819, 1032, 1026]. facultative [738]. fair [873]. FAO [506]. far [881]. Farm [875, 904]. farmed [629]. farmers [1018, 829]. farming [827, 690, 556, 601, 829]. farms [622]. farsighted [604]. farsightedly [815]. fast [881]. Faustmann [776, 561]. feasibility [1010]. Features [377, 383]. fecundity [844]. federal [186]. feedback [1045, 506, 281, 309, 382, 625]. feedback-control [506]. feel [841]. feral [745, 717]. fertile [452]. fertility [531]. fertilization [945]. fertilizer [571]. Fever [674]. field [396, 890]. field-of-neighborhood [396]. fields [297]. filling [962]. filter [733]. financial [1033, 292]. findings [778]. finite [130]. fir [260]. Fire [458, 388, 118, 174, 25, 952]. fires [812]. firm [568]. firms [215]. First [364, 373, 708]. Fish [412, 983, 1007, 991, 666, 205, 470, 621, 525, 526, 453, 744, 569, 1010, 285, 679, 855, 331, 587, 447, 107, 246]. Fisheries [596, 105, 995, 522, 703, 128, 775, 523, 562, 599, 988, 570, 521, 544, 635, 504, 435, 580, 312, 788, 1010, 445, 554, 893, 440, 267, 286, 127, 619, 84, 426, 729, 506, 380, 436, 930, 389, 777, 408, 279, 709, 1019, 734, 282, 51, 334, 783, 1053]. fishers [412, 855]. Fishery [113, 693, 416, 19, 648, 934, 271, 862, 1023, 708, 892, 195, 244, 204, 58, 78, 663, 140, 423, 856, 873, 469, 911, 572, 570, 927, 832, 894, 328, 411, 605, 702, 698, 579, 284, 527, 670, 639, 925, 288, 9, 463, 338, 833, 471, 441, 377, 415, 37, 217, 437, 310, 450, 754, 770, 414, 787, 378, 965, 573]. fishes [939]. Fishing [454, 553, 271, 27, 157, 1047, 262, 788, 460, 673, 564, 664, 759, 150, 787, 625]. fishing-up [625]. fishmeal [587]. fishy [666]. fittest [451]. fitting [559]. flammable [467]. Flavescence [929]. fleet [205]. fleets [377]. Flexible [269]. Flood [1044, 922, 860, 880]. Flood-oriented [1044]. Flooding [182]. flow [634, 901, 309]. flows [920]. fluctuation [542, 453]. fluency [969]. Flushing [645]. Focus [863]. follows [1047]. food [996, 298, 379, 915, 973, 911, 891, 432, 634, 857, 581, 574]. food-chain [379, 581]. footprint [1030]. forage [66, 146, 333]. forage-ungulate-predator [66]. foraging [397, 542]. forcing [297, 306]. forecast [984]. Forecasting [218, 427, 600, 722]. forecasts [182, 527]. Foreign [468]. Forest [444, 234, 647, 57, 535, 292, 1012, 75, 260, 831, 347, 889, 494, 323, 214, 547, 1036, 987, 39, 1040, 812, 948, 308, 1046, 802, 402, 336, 644, 388, 368, 369, 174, 96, 613, 256, 180, 3, 25, 88, 235, 87, 951, 949, 455, 467, 578, 511, 916, 561, 346, 225, 947, 897, 904, 953, 876, 884]. forest-wide [75]. Forestland [993, 650, 955]. forestry [998, 767, 724, 129, 236, 576]. forests [776, 954, 1011, 852, 719]. form [563]. formation [501, 366, 657]. formulas [776]. formulation [316]. forward [427]. foundations [8, 841]. FP [880]. fractal [301]. fractional [865]. fractional-order [865]. fracturing [874]. fragmented [323, 595]. framework [911, 976, 600, 1039, 747, 749, 999, 726, 346, 828]. free [1047, 490]. French [892]. Frequency [970, 648, 860]. Frequency-dependent [970]. frontiers [589]. fuel [491, 951, 583]. full [749]. fully [416]. function [563, 261, 108, 182, 312, 747, 35, 919, 20, 210]. functional [379, 563, 915, 973, 537, 936]. functions [592]. fundamental [549]. future [892, 808, 157, 105, 1040, 435, 670, 436, 190, 1014, 995, 210]. futures [666, 77]. fuzzy [918].

G20 [1030]. Gadus [965]. gains [420, 129]. Game [286, 209, 114, 316, 604, 741, 501, 545, 348, 69, 312, 313, 127, 607, 610, 51]. Game-theoretic [286, 209, 741, 607]. games [627, 562, 294, 331]. Ganges [850, 943, 1031, 882]. gaps [962]. gas [848, 223]. Gauss [227]. GEM [1003]. gender [904]. General [697, 915, 973, 905, 251, 129, 754]. generalist [659]. Generalized [792]. generated [590, 617]. generational [530]. generations [474]. genetic [728, 129]. genetically [479, 963]. genotypes [818]. geolocation [813]. geological [683]. Geometrical [928]. Getting [539, 386]. Gibe [1014]. Gierer [304]. GIS [805]. glaucescens [485, 484]. glaucous [485, 484]. glaucous-winged [485, 484]. Global [214, 526, 646, 417, 892, 991, 983, 123, 523, 501, 974, 587, 1003]. goals [857, 887]. Golden [439, 488, 441]. golden-winged [488]. Golestan [960]. good [1027]. Gorgan [880]. gorges [821]. gorillas [678]. governance [938]. Gower [916]. Gower-based [916]. gramma [515]. graphene [1004]. grassland [147, 902, 1002]. Gray [706, 702]. grazing [512, 614, 641, 121]. great [193, 745, 717, 909, 795, 510, 509]. greater [676]. Greece [461]. Greed [1027]. green [1051, 1037, 899, 999]. greenhouse [314]. Greenland [522]. grey [1007]. Groundwater [883, 316, 982, 746, 940, 287, 329, 920, 765, 874, 269, 155]. groundwater-dependent [746]. groups [833, 1000]. Growth [108, 579, 629, 203, 991, 244, 234, 95, 675, 502, 932, 644, 465, 96, 425, 818, 235, 732, 383]. growths [918]. Guangdong [905]. Guiana [892]. guide [811, 795]. guiding [40]. Gulf [529, 460, 965]. gull [586]. gull/human [586]. gulls [485, 484]. gully [1029]. gypsum [945].

Habitat [485, 484, 956, 260, 811, 323, 988, 504, 665, 833, 416, 466, 320, 953]. habitat-quality [811]. hake [471]. Halibut [442, 543, 272, 217]. hand [824, 785, 813]. hand-picked [785, 813]. handling [840]. haphazard [823]. harbor [684]. Hardin [859]. Hartman [561]. Hartwick [358]. Harvest [330, 535, 1052, 75, 885, 520, 5, 927, 971, 577, 660, 619, 833, 25, 235, 310, 626, 813, 872, 20, 719, 625]. harvested [324]. Harvesting [67, 47, 621, 285, 340, 560, 629, 141, 682, 430, 839, 856, 973, 572, 891, 29, 782, 49, 453, 744, 149, 925, 910, 538, 28, 68, 581, 447, 107, 867, 180, 3, 477, 466, 770, 513, 511, 752, 1016, 723, 1053]. hatchery [138]. haulout [684]. Hawai’i [589, 787, 857]. Hawassa [1029]. health [1044, 1032]. heat [590]. HEC [880]. HEC-RAS [880]. height [367]. help [991, 953]. herbivore [541, 97, 842]. heterogeneity [322, 730, 1053]. heterogeneous [629]. Heuristic [1027, 948]. hierarchical [316, 367, 954]. high [452, 284, 312, 660, 426, 664]. high-dimensional [660]. highly [811]. Himachal [917]. historic [644]. Historical [1014]. history [766]. HIV [984]. HIV/AIDS [984]. hog [795]. hogs [745, 717]. holders [291]. holes [834]. Holling [973]. Holt [630, 927]. honey [194, 303]. Hooghly [653]. horizons [638]. horseshoe [615]. host [4, 243, 738, 837]. host-parasitoid [4, 243]. host-vector [837]. hostile [363]. hosts [972]. hotspots [804]. household [875, 904]. households [993]. housing [391]. Hubbell [480]. Hubbert [957]. Hulunbuir [1002]. Human [531, 903, 586, 979, 842]. human-transported [979]. humile [552]. hunting [650, 348, 843, 707, 822]. Hurricane [799]. Hurst [301]. hybrid [996]. hydraulic [874, 880]. hydrocarbon [187, 188]. Hydrogeochemical [940]. hydrograph [922]. hydrographic [919]. hydrologic [1032]. hydrological [976, 1024, 990, 854, 941]. hyperbolic [499]. hyperparasitoids [972]. hyperplanes [428]. hyperspectral [890].

Iceland [522]. Icelandic [673]. ideal [1047]. identifiability [1042]. identification [34, 246]. identifying [261, 979]. ignition [458]. ii [485, 184, 566]. illegal [987, 348]. illustration [827]. image [1029]. imagery [890, 880]. Impact [987, 555, 846, 522, 432, 321, 628, 820, 447, 646, 664, 938, 704, 887, 876]. impacted [939]. Impacts [687, 844, 150, 982, 909, 596, 523, 118, 504, 457, 855, 603, 408, 754, 578]. implement [688]. implementation [360, 854]. implementing [506]. Implications [777, 525, 857, 450, 454]. Importance [831, 498]. important [886, 416]. improve [761]. improving [188, 269, 1004]. impulsive [630]. Imputing [275]. In-season [58, 78]. Inaccessible [487]. Incentive [185]. incentives [139, 529, 594]. including [832]. inclusion [1033]. income [175]. incomplete [525]. incomplete-information [525]. inconstant [648]. Incorporating [793, 840, 689, 611, 589, 671]. Incorporation [369]. increase [619]. increased [605, 768]. independent [28]. Index [64, 104, 137, 172, 229, 259, 278, 295, 317, 898, 202, 897]. India [653]. Indian [1045]. indicator [494, 759]. indicators [507]. indices [613]. Individual [195, 540, 401, 396, 530, 541, 322, 518, 328, 542, 393, 636, 662, 718, 386, 398, 394, 717, 676, 132]. Individual-based [540, 401, 396, 530, 541, 328, 542, 393, 636, 718, 386, 398, 394, 717, 676]. individual-level [132]. Induced [617, 802, 581, 371, 740]. inducements [139]. Industrial [224, 441]. industrialization [620]. industries [767]. industry [556, 272, 673, 299, 787, 98]. inefficiency [361]. infectious [980, 753, 989]. infer [644]. inference [695, 984, 549]. Inferring [463]. infestation [194, 626]. Influence [955, 920, 369, 901, 659, 847, 668, 904]. Influences [681]. influencing [1026]. inform [811, 734]. Information [654, 661, 845, 858, 869, 877, 896, 908, 923, 931, 946, 959, 966, 977, 986, 997, 1001, 1005, 1013, 1021, 1028, 1034, 1041, 1048, 359, 669, 677, 686, 694, 701, 711, 715, 721, 727, 735, 742, 750, 755, 763, 772, 779, 789, 798, 806, 814, 825, 835, 423, 525, 365, 404]. information-theoretic [404]. informational [233]. infrastructure [999]. infrastructures [1037]. inheritance [611]. initio [400]. injured [884]. Inner [1002]. Innovation [871, 1033]. inorganic [653]. input [889, 222, 990, 573]. insect [689, 154]. insects [771]. inshore [624]. insight [725]. insights [286]. instability [304]. instar [4, 243]. institutional [291, 938]. instruments [329]. insurance [519]. Integer [781]. integral [38]. integrated [59, 683, 1039, 508, 440, 830, 516]. Integrating [411, 196, 948]. integrative [509]. intelligence [906]. intended [852]. intensity [905]. inter [571]. inter-sectoral [571]. interacting [640, 800, 974, 138, 743]. interaction [929, 417, 530, 541, 738, 303]. interactions [1052, 4, 460, 66, 844, 446, 498, 299, 561, 783]. interactive [227, 226]. interconnected [655]. interdependence [568]. interdependency [444]. interdependent [418]. interests [702]. interior [323, 567, 860]. intermediary [914]. International [364, 373, 705, 442, 431, 496, 545, 886, 554, 286, 815, 664, 758]. interplay [983]. interpretation [392]. interpretive [999]. Intertemporal [558, 544, 250, 87]. intertemporally [688]. intertidal [410]. intervention [764, 888]. intraspecific [706]. intrinsic [563]. Introduction [364, 373, 773, 296, 300, 521, 736, 435, 349, 354, 231, 241, 318, 393, 539, 311, 342, 643, 826, 870, 631, 283, 335, 516, 325, 350, 490, 406]. invaders [479]. invading [308]. invariant [563, 378]. invasion [479, 564, 797, 575, 552]. invasions [528]. invasive [840, 939, 761, 921, 652, 533, 769, 781]. inventory [1012, 211, 275]. inventory-based [1012]. investigate [612]. Investigating [658, 461, 823, 542, 462, 751]. investigation [588]. investment [57, 468, 424, 94, 904]. investments [699, 558]. Iran [898, 1044, 922, 912, 1003, 880, 960]. Iranian [860]. irreversibilities [315]. Irreversible [57, 433]. irrigated [600, 818, 692]. irrigation [687, 148, 507, 622, 273, 672]. Isfahan [912]. Island [653, 602, 461]. islands [590]. isolated [320]. Isometric [546]. Issue [654, 661, 669, 677, 686, 694, 701, 711, 715, 721, 727, 735, 742, 750, 755, 763, 772, 779, 789, 798, 806, 814, 825, 835, 845, 858, 869, 877, 896, 908, 923, 931, 946, 959, 966, 977, 986, 997, 1001, 1005, 1013, 1021, 1028, 1034, 1041, 1048, 773, 335, 364, 373, 924, 713, 296, 300, 489, 521, 456, 505, 736, 435, 318, 393, 539, 473, 481, 311, 342, 643, 826, 870, 631, 516, 947]. Issues [438, 233, 725, 349, 354, 342, 279]. itch [933]. iterative [47, 173]. ITQ [360, 693]. IV [226].

jacobsoni [303]. Japanese [149]. Jasper [552]. Java [384]. Jeffara [940]. Jiangxi [993, 904]. Joint [1007, 873, 608, 146]. Jointly [536]. Jointly-determined [536]. jumping [957]. jurisdiction [27]. juvenile [488, 639]. juvenile-adult [639].

Kalman [733]. kangaroos [293]. Kanpur [1031]. Kappa [747]. Kenya [343]. kernel [458, 886]. Khingan [900]. Khorasan [1044]. Khoshk [922]. KINEROS2 [1024]. kisutch [527]. Kızılcasu [851]. knots [615]. Kolmogorov [427]. Kona [589].

L. [194, 303, 818]. Labor [904, 469, 771, 955]. Lactarius [851]. laden [828]. lag [1002]. Lake [1004, 909, 528, 575, 903, 1029]. lakes [193]. lamprey [761]. Lanchester [302]. Land [106, 821, 960, 390, 816, 291, 935, 234, 917, 1051, 1038, 879, 582, 786, 1014, 751, 351, 186]. land-use [879, 186]. landings [580]. landlord [712]. landowner [647]. landscape [323, 399, 457, 458, 902, 613, 404, 1004]. landscape-level [404]. landscapes [953]. landslide [819]. Language [158, 227, 226]. Lanka [805]. large [1012, 326, 16, 336, 548, 941, 546]. large-scale [1012, 16, 941]. larus [485, 484]. larval [376]. larvicide [837]. laser [900, 852]. late [376]. latent [886]. lateral [920]. law [284]. laying [975]. lead [768]. learn [699]. Learning [465, 464, 7, 948, 824, 600]. least [660]. least-squares [660]. legacy [945]. length [878]. Lenhart [836]. Leslie [916]. Lesser [900]. lessons [778, 334]. Lesvos [461, 457]. lettuce [861]. level [517, 793, 458, 194, 398, 132, 404]. levels [60, 722]. liberalization [609]. license [650]. life [400, 766, 138]. light [1004]. likelihood [618]. limitations [644, 726]. limited [824, 408, 282]. limited-entry [408]. limpet [244]. linear [356, 1043]. linearization [769]. linefishing [513]. Linepithema [552]. linkage [713]. linkages [415]. LISFLOOD [880]. LISFLOOD-FP [880]. literature [810]. livelihoods [855]. livestock [885]. LMDI [942]. loading [1029]. loads [583]. loblolly [164, 628]. local [857, 339, 378]. localized [118]. Locational [351]. locations [979]. log [391]. Logging [884, 987]. logistic [932]. Long [211, 314, 361, 684, 336, 849, 853]. Long-run [211, 314, 361]. long-term [684, 849, 853]. longline [787]. longterm [719]. look [105, 636, 88]. Loot [860]. losses [343]. Lotka [357]. Louisiana [982]. Low [881]. Low-density [881]. lower [964]. Lyme [166].

Macintosh [184]. mackerel [150]. magic [439]. Magnuson [113]. major [476, 829]. make [393]. makes [266]. making [501, 589, 332, 960, 709, 758]. mallard [5]. mammals [1007]. manage [913]. managed [775, 992, 561, 953]. Management [130, 761, 968, 113, 539, 87, 945, 568, 757, 883, 292, 1052, 685, 260, 708, 892, 831, 1007, 950, 983, 58, 78, 655, 47, 617, 911, 652, 800, 214, 746, 205, 988, 536, 976, 656, 14, 540, 778, 59, 940, 399, 544, 586, 635, 411, 605, 124, 284, 1039, 714, 670, 445, 554, 679, 893, 242, 288, 236, 286, 463, 615, 127, 1027, 46, 84, 426, 256, 122, 332, 506, 641, 377, 3, 930, 524, 777, 1037, 281, 382, 837, 949, 549, 578, 888, 734, 995, 928, 20, 282, 74, 120, 196, 965, 334, 904, 1053, 828]. management [155, 884]. managers [969]. Managing [418, 780, 688, 447, 853, 743, 713, 505, 504, 527, 679, 638]. mangrove [581]. manure [568]. many [362]. mapping [879, 880]. marginal [662].

MARGOT [1012]. Maria [799]. Marine [664, 498, 414, 783, 409, 413, 1007, 412, 596, 988, 410, 407, 544, 545, 635, 411, 624, 623, 925, 739, 730, 446, 594, 610, 566, 416, 415, 408, 454, 406]. market [139, 647, 272, 951, 391]. markets [491, 359, 693, 77, 587, 578, 186]. Markov [939, 816]. marshallian [282]. Martin [936]. Matenchose [1038]. materials [108]. Mathematical [632, 364, 373, 689, 998, 667, 598, 586, 802, 823, 753, 756, 933, 787]. mating [618]. Matla [653]. MATLAB [207]. Matrix [392, 227, 226, 56, 85, 254, 766, 868, 784]. matter [810, 493]. matters [800]. maturity [810]. Maximin [422, 428, 926]. Maximum [1047, 770, 912, 404]. may [29, 768, 425]. MCDM [1039]. MCDM-SWOT [1039]. McKelvey [232]. mean [878, 599]. mean-variance [599]. measure [498]. Measurement [1043, 203, 579, 331, 129]. measures [648, 938]. Measuring [909, 570, 403, 471]. meat [348]. mechanism [580]. mechanisms [584]. Medenine [940]. media [606]. mediated [839, 844]. Mediterranean [128]. meets [1049]. megaherbivore [520]. Meinhardt [304]. meir [121]. Mekong [713]. mellifera [194, 303]. Members [198]. memory [69]. Mesopelagic [1010]. mesoscale [590]. metaheuristic [875, 949]. metaphysiological [244, 327]. metapopulation [968, 483, 321, 745, 482, 477, 381, 340]. metapopulations [605]. method [684, 592, 559, 478, 548, 860, 575, 949, 957]. methodological [976]. methodology [494, 263, 999]. Methods [731, 837, 40, 913, 355, 349, 354, 458, 50, 673, 533, 951, 928, 74]. metrics [811]. metropolis [999]. Mexico [529]. Michigan [734]. micro [517]. micro-level [517]. microfoundation [500]. microprototype [889]. middle [1051]. might [1025]. migration [955]. migrations [526]. migratory [285, 120]. MIKE11 [1004]. mineral [218, 532]. minimize [447]. minimum [882]. mining [802]. Missed [242]. missing [275]. Mississippi [982]. mite [194, 303]. mitigation [688]. mitra [724]. mitra-wan [724]. Mixed [389, 757, 614, 1000]. mixed-grazing [614]. mixed-species [1000]. Mixed-stock [389]. mobile [801, 811, 577]. Model [980, 1003, 451, 553, 397, 203, 1012, 417, 1035, 943, 413, 878, 271, 862, 326, 689, 591, 658, 892, 147, 611, 244, 379, 58, 78, 935, 530, 541, 442, 764, 760, 748, 543, 243, 741, 485, 915, 973, 323, 387, 572, 891, 968, 794, 507, 937, 483, 571, 327, 976, 1024, 821, 182, 905, 592, 328, 293, 784, 622, 683, 802, 823, 69, 542, 738, 922, 857, 508, 724, 285, 745, 639, 886, 961, 15, 369, 26, 492, 194, 303, 718, 482, 979, 512, 66, 912, 730, 581, 446, 797, 941, 933, 484, 166, 575, 180, 398, 551]. model [771, 415, 389, 304, 477, 717, 87, 281, 837, 34, 117, 472, 676, 1019, 466, 601, 770, 859, 299, 513, 1008, 132, 414, 403, 138, 787, 302, 888, 583, 732, 574, 668, 561, 936, 734, 751, 758, 726, 965, 429, 534, 671, 334, 121, 1004, 783, 842, 989, 722, 781, 1026, 816, 817]. model-based [892]. modeled [744]. Modeling [590, 568, 291, 1012, 809, 249, 808, 148, 773, 640, 913, 987, 399, 344, 348, 504, 712, 851, 539, 336, 767, 765, 556, 662, 653, 462, 479, 900, 700, 174, 643, 587, 471, 345, 938, 310, 450, 421, 620, 657, 532, 455, 578, 246, 762, 649, 120, 493, 722, 719, 1000, 648, 400, 944, 632, 791, 175, 233, 396, 637, 1044, 140, 911, 585, 598, 470, 105, 118, 163, 998, 940, 890, 435, 624, 642, 287, 457, 792, 368, 747, 731, 2, 461, 615, 209, 342, 854, 870, 371, 441, 667, 88, 631, 749, 37, 1032, 999]. modeling [638, 552, 509, 454, 834, 697, 916, 404, 516, 957, 963, 248, 790, 829]. Modelling [831, 832, 385, 35, 335]. Models [4, 452, 51, 648, 737, 520, 38, 67, 297, 298, 76, 633, 47, 56, 85, 254, 766, 495, 864, 475, 29, 621, 634, 540, 681, 971, 395, 39, 164, 544, 586, 559, 411, 698, 355, 366, 569, 932, 402, 367, 733, 838, 401, 393, 636, 548, 50, 910, 662, 267, 830, 616, 289, 436, 377, 880, 386, 3, 810, 394, 769, 268, 499, 114, 437, 960, 841, 438, 1042, 954, 406, 306, 852, 351, 392, 953, 1043]. MODFLOW [940]. modified [479, 963]. MODIS [917]. moments [860, 439]. monarch [784]. Mongolia [1002]. Mongolian [987]. monitoring [684, 365, 114, 346]. Monopoly [359, 156]. monopsony [710]. monotonicity [235]. Monte [588, 660]. month [1002]. monthly [944, 722]. morphogenesis [304]. morphological [246]. morphology [906]. mortalities [5, 513]. mortality [67, 651, 662, 616, 289, 954, 884]. mountain [559, 678]. Mountains [795, 745, 717]. movement [939, 612, 205]. movements [399]. mover [708]. MRAA [982]. MT3DMS [940]. Multi [1002, 412, 602, 232, 589, 471, 403, 334]. multi-activity [412]. multi-criteria [589]. multi-disciplinary [232]. multi-model [403]. Multi-month [1002]. multi-predator [602]. multi-species [412, 471, 334]. multiattribute [332]. multicriteria [926, 960]. multicyclic [957]. multidimensional [1022, 733, 1018]. Multifunctional [879, 719]. multifunctionality [935]. multilevel [16, 173]. Multiobjective [949, 1017, 960]. Multiple [276, 474, 1017, 760, 748, 774, 154, 60, 610, 1019, 333]. multiple-species [1019]. multiple-use [333]. multiplicative [927]. Multiscale [1003]. Multispecies [682, 569, 615, 260, 163, 9, 729, 616, 498, 438, 726]. multiuse [988]. multivariate [981]. mushrooms [851]. mutual [891]. mutualisms [914]. myopically [815]. myoporum [512].

Namibian [513]. Nash [387, 331]. National [795, 568, 595, 745, 717]. nations [27, 157]. Natural [648, 419, 643, 77, 335, 1033, 757, 903, 591, 924, 5, 598, 969, 275, 832, 778, 173, 402, 48, 242, 337, 2, 433, 870, 871, 386, 631, 268, 362, 450, 549, 1042, 872, 138, 257, 74, 493, 828, 790]. Nature [1018, 720]. near [362]. near-optimal [362]. Nearly [838]. needed [938]. Negotiating [109]. neighborhood [396, 816]. nematode [820]. nematodes [780]. Neoclassical [8]. nesting [323, 462]. Net [151, 212, 254, 424]. network [996, 585, 547, 890, 906, 854, 938, 953]. network-based [890]. networks [528, 490, 847, 532]. Neural [854, 585, 890, 906, 532]. neurodynamic [575]. neutral [480]. newly [941]. nexus [996, 1010]. Nile [301, 837]. nitrogen [206, 571, 588, 653, 704, 945]. Noah [567]. noise [35, 603, 841]. non [262, 767, 107, 282, 98]. non-cooperative [262]. non-marshallian [282]. non-renewable [98]. non-stationary [107]. non-wood [767]. nonautonomous [936]. nonconcavity [567, 696]. noncooperating [713]. Noncooperative [708, 562, 315]. nondelayed [584]. nondetection [695]. nonharvesting [723]. nonindustrial [578]. Nonlinear [56, 629, 175, 141, 140, 85, 823, 744, 303, 268, 916, 1043]. Nonmalleable [115, 217]. Nonmarket [277, 8]. nonnative [910]. nonpoint [316]. Nonrenewable [459, 216, 115, 250, 274]. Nonspatial [633]. nonstationary [1036]. nonstorable [666]. nontrivial [567]. Nordic [1007, 601]. normative [264]. North [413, 271, 486, 310, 751]. northeast [708, 166, 514]. Northern [796, 520, 702, 471]. Northwest [1040, 942, 391]. Norwegian [934, 556]. nose [968]. note [99, 240, 489, 456, 505, 17, 473, 481, 230]. notes [151, 212]. Notices [81, 136]. novel [559]. noy [121]. noy-meir [121]. nuisance [176]. nuisance-beaver [176]. number [897]. numbers [349, 354]. Numerical [475, 305, 363, 427]. nutrient [653, 945, 668]. nutrients [497]. NW [460].

OBIA [1029]. Object [1029]. Object-based [1029]. Objective [1003, 60]. objectives [518, 1049, 949]. obligate [738]. observations [880]. Observed [912]. obtaining [769]. occupancy [485, 681, 484]. occur [1025]. occurrence [458]. ocean [974]. off [935, 863, 601, 904]. off-farm [904]. offs [374, 536, 812]. offsets [1049]. oil [24, 223, 529, 587, 499, 696]. old [234, 235]. old-growth [234, 235]. oligopoly [518]. olive [847]. Omo [1014]. Omo-Gibe [1014]. Oncorhynchus [527]. One [606]. One-dimensional [606]. Open [486, 165, 1023, 267, 465]. open-access [1023, 267]. opportunities [790, 242]. Opportunity [186]. Optimal [674, 629, 840, 1052, 474, 266, 764, 243, 911, 245, 921, 572, 652, 800, 154, 988, 894, 971, 49, 176, 683, 577, 69, 744, 149, 925, 550, 785, 433, 216, 107, 867, 256, 528, 641, 3, 94, 557, 626, 513, 146, 333, 1011, 872, 511, 576, 888, 752, 334, 181, 315, 1053, 625, 57, 4, 326, 563, 760, 234, 973, 214, 1036, 365, 555, 164, 276, 263, 97, 580, 223, 478, 1050, 17, 724, 731, 26, 337, 28, 68, 699, 730, 370, 127, 476, 575, 180, 377, 25, 268, 362, 281, 382, 467, 378, 1016, 723, 995, 928]. Optimal-sustainable [334]. optimality [121]. Optimization [819, 875, 1017, 518, 464, 261, 16, 173, 614, 277, 729, 37, 1037, 960, 390, 74, 719, 816]. optimize [965]. Optimizing [950, 308, 15, 692, 260]. Optimum [96]. option [123, 6, 660, 255]. Options [128, 666, 582]. orbit [722]. order [865]. ordered [533]. ordinal [551]. ordinal-valued [551]. organic [589, 700, 704, 493, 829]. organizations [445]. Organized [720]. organizing [390]. oriented [1044, 386]. Oryza [818]. oscillation [357]. oscillations [617]. Ostrom [859]. other [1042]. Ottawa [643]. oudemans [303]. outbreaks [372]. outcome [707]. outcomes [964]. outdoor [939]. outflow [301]. outpacing [1035]. Output [573, 889, 222]. overgrazing [339]. owl [391, 796]. Owls [796]. owned [351]. owner [867]. ownership [444, 36, 249, 650]. oxides [588, 704]. oxygen [700].

Pacific [442, 703, 543, 1040, 272, 217, 391]. package [184]. Panda [1018]. Panxi [935]. parabolic [744]. Paradigm [440]. Parameter [1024, 681, 183, 569, 1042, 574, 897]. parameters [990]. parametric [598]. parasite [738]. parasitic [303]. parasitism [738]. parasitization [4, 243]. parasitoid [4, 243]. parasitoids [972]. park [595, 745, 717, 795]. parrotfish [864]. part [187, 188, 566]. Partial [251, 744, 180]. Participation [496]. particles [593]. parties [603]. passenger [486, 937]. Past [670, 435]. pastoral [1050]. Pasture [885]. patch [116, 764, 485, 665, 730, 484, 340]. patch-type [116]. patches [320]. patchy [590]. path [358]. pathogens [979, 881]. Paths [969, 362]. pattern [366, 371, 386, 834]. pattern-oriented [386]. patterns [939, 458, 388, 25, 752]. Paying [1046]. payments [887]. payoff [873]. payoffs [313]. PC [226]. PC-DOS [226]. peak [372]. peak-to-peak [372]. pelagic [856, 621]. penalties [580]. perceived [993]. perception [582]. perfect [420]. perfectly [675]. performance [15, 414, 692]. period [476, 357]. Periodic [430, 297, 495, 784, 117, 306]. periods [780, 336]. periphery [595]. perishable [785]. permeability [320]. permits [156]. Persistence [116, 29, 538, 914, 298, 379, 322, 642, 318, 320]. personal [716, 437]. perspective [904]. perspectives [631]. perturbation [349, 354]. perturbations [893, 350]. perturbed [352, 353]. PES [964]. Pest [372, 4, 40, 689, 59, 780, 174]. pests [308, 479]. petrel [602]. Petroleum [46, 658]. pheromone [761]. pheromone-baited [761]. Phoca [684]. phosphorus [945]. photocatalysis [1004]. picked [785, 813]. piece [313]. piece-wise [313]. pig [324]. pigeon [486, 937]. Pigs [603]. piles [819]. pine [164, 559, 628]. pioneer [117, 306]. pioneer-climax [117, 306]. placement [413]. plague [913]. plain [880]. plankton [653]. Planning [610, 1037, 330, 948, 567, 851, 60, 719]. plant [396, 518, 97, 881, 657]. plant/herbivore [97]. plantation [963]. plantations [628]. planting [245]. plants [848, 479]. platycarpum [512]. platyphylla [900]. play [753]. plover [649]. PM [961]. poachers [853]. poaching [1045]. Point [1009, 889, 1047, 695, 581, 928, 121]. point-input [889]. point-output [889]. points [458, 995]. policies [468, 637, 374, 106, 617, 688, 857, 25, 787]. policy [316, 234, 47, 973, 891, 571, 609, 251, 457, 28, 68, 343, 871, 224, 777, 449, 704, 709, 1053]. political [943]. Politics [703]. pollination [914]. pollination-mutualisms [914]. pollock [515, 570, 965]. pollutant [545]. pollutants [584, 189, 700]. polluted [867]. Polluting [1015]. pollution [590, 165, 943, 1031, 1023, 316, 604, 148, 571, 609, 823, 124, 459, 510, 557, 758, 181]. polygenic [611]. ponds [939]. pool [907, 710, 338, 852]. Population [141, 495, 375, 398, 451, 846, 116, 417, 1035, 796, 38, 67, 611, 297, 517, 47, 684, 325, 56, 254, 630, 95, 612, 432, 29, 327, 634, 782, 618, 971, 868, 293, 706, 308, 784, 802, 744, 569, 932, 367, 838, 747, 1045, 303, 659, 345, 436, 552, 620, 472, 676, 770, 560, 320, 649, 392]. population-based [308]. Population-level [398]. population-resource [417]. populations [629, 396, 347, 379, 655, 322, 85, 617, 766, 640, 495, 154, 540, 176, 707, 366, 324, 318, 461, 107, 810, 743, 842]. porous [606]. Portfolio [614]. positional [458]. positive [264, 787]. possibilistic [996]. Possibilities [644]. possibility [425]. possible [514]. Potential [982, 359, 793, 393]. poverty [1018]. power [710, 272, 942]. Practical [298]. practices [292, 603]. Pradesh [917]. pragmatic [16]. prairie [913]. precautionary [506, 930]. precipitable [961]. precipitation [912, 1003, 657, 1002]. predation [725, 659]. predator [970, 1052, 530, 602, 500, 86, 843, 537, 538, 267, 68, 66, 797, 132, 340, 574, 936, 649, 334]. predator-prey [530, 500, 538, 267, 132, 340, 334]. predators [345]. predict [766]. Predicting [878, 586, 684, 187, 188]. predictions [1035]. predictive [734]. preference [697]. preferences [795]. preliminary [524]. premiums [622]. prescription [75]. presence [864, 1047, 802, 823, 453, 465, 174, 433, 426, 874, 916, 20]. present [435]. preservation [48, 255]. Preserve [552]. Preserving [804]. pressure [1044, 802, 466]. prevention [571, 812]. preventive [196]. prey [970, 1052, 530, 891, 500, 86, 843, 538, 267, 68, 797, 132, 340, 936, 334]. price [211, 555, 164, 448, 87, 872, 511, 274, 210, 391]. prices [24, 210, 696]. pricing [361, 692, 625]. primary [421]. principal [27]. principal-agent [27]. principle [40]. principles [282]. Private [775, 359, 578, 225]. proactive [679]. Probabilistic [850]. probability [648, 321, 954]. problem [535, 757, 943, 563, 567, 478, 724, 370, 867, 217, 549, 378]. problems [474, 464, 731, 337, 498, 949, 339, 353]. procedure [769]. procedures [47]. process [976]. processes [471, 981]. produced [479]. producer [541]. product [876]. Production [98, 211, 195, 866, 875, 108, 529, 639, 589, 471, 899, 146, 333, 957]. productivity [570, 851, 336, 616, 995]. products [962, 333]. profitability [1049, 370]. program [964, 429, 534]. programming [996, 227, 226, 598, 575, 787, 781]. programs [139]. progress [656, 129]. progressed [506]. projected [912]. projecting [39]. projection [706]. projects [976]. proliferation [921]. prolonged [909]. promise [788]. promote [887]. propagating [1012]. properties [563, 123, 890, 66, 613, 1042, 828]. property [801, 193, 307, 249, 419, 294, 332, 531]. prophylactic [837]. prorationing [710]. prospects [344, 678]. protect [470]. protected [413, 596, 410, 411, 610, 664, 566, 416, 406, 414]. Protecting [25, 953, 678]. protection [553, 871, 25, 94, 391]. protective [617]. protocols [501]. provenance [644]. Province [1044, 960, 912]. provision [519]. PSR [1044]. Public [225, 291]. publicly [351]. publisher [240, 230]. publishing [790]. Puerto [799]. Puge [879]. pulse [888]. purse [204]. Pv [190, 849]. Pv-wave [190].

qualitative [76, 563]. quality [811, 222, 329, 925, 902, 269, 155]. Quantitative [673, 509, 635]. quantitatively [939]. quantities [868]. quantity [902]. quasi [123]. quasi-option [123]. quasilinear [451]. queuing [996]. quota [291, 1019, 282]. quota-regulated [1019]. quotafishing [374]. quotas [195, 206].

R&D [699, 216]. radial [356]. rain [345, 657]. rainfall [944, 962, 919, 641]. rainwater [1026]. RAMAS [177]. Ramsar [886]. Ramsey [203]. Random [580, 236, 374, 49, 662, 952]. Range [900, 60]. rangelands [512, 614, 641]. rank [533]. rapid [187, 188, 607]. rarity [762]. RAS [880]. rate [361, 1029, 651, 48, 187, 188, 306]. rate-based [651]. rates [662, 680, 616, 20]. rational [420, 19, 1019]. rationalization [696]. reach [580]. reached [425]. reaction [352, 552, 305, 363, 210]. reaction-diffusion [352, 305, 363]. reactive [679]. reality [800]. realize [393]. reallocating [403]. reallocations [310]. reauthorization [113]. Rebuilding [566, 524]. recessive [479]. record [949]. record-to-record [949]. recovery [846, 69, 450, 299]. recreational [564, 150]. recruitment [38, 67, 35, 376, 107, 676, 404]. red [615, 310]. reduce [980, 622, 619]. reducing [618, 216]. reduction [865, 355, 607]. Reef [509, 864, 510]. reefs [809]. reemergence [780]. reference [1047]. reforestation [585]. regeneration [174]. regime [995]. regimes [844, 225]. region [935, 962, 919, 509]. Regional [754, 568, 148, 59, 445, 961, 1002]. regions [898, 940]. regression [662, 1043]. regulated [1019]. Regulation [378, 710, 128, 206, 496, 765, 1018, 9, 433, 804, 19, 339, 692, 181]. regulations [185, 876]. regulators [688]. Regulatory [272, 329]. reinforcement [464]. reintroduction [937]. Relationship [866, 780, 899]. Relationships [961, 5, 107, 546]. relative [215, 497, 1018]. relaxation [357]. remedial [196]. REMO [912]. removal [584, 840, 793]. remuneration [1046]. renewable [76, 130, 640, 666, 16, 580, 577, 714, 492, 465, 1027, 867, 362, 281, 421, 620, 1016, 74, 98]. Rennie [194]. rent [218, 688]. rents [693]. repeated [828]. replacement [488]. reproduction [652, 96]. reproductive [254]. research [497, 280, 288]. researchers [969]. Reserve [1018, 412, 415, 466, 783]. reserve-fishery [415]. reserves [409, 470, 14, 407, 595, 707, 623, 739, 730, 594, 264, 853, 498, 416, 408, 454, 1018]. reservoir [821, 46]. residential [351]. Resilience [828, 924, 1022, 778, 756, 1027, 826, 582, 667, 691, 799]. resistance [820, 642]. resistant [820, 479]. resolution [605]. Resource [790, 643, 335, 648, 359, 801, 417, 193, 307, 36, 233, 76, 419, 222, 130, 266, 748, 387, 598, 640, 969, 475, 464, 86, 976, 275, 778, 1015, 16, 173, 577, 802, 69, 242, 26, 337, 459, 492, 2, 465, 209, 77, 1027, 277, 871, 867, 94, 631, 115, 250, 268, 362, 499, 382, 450, 620, 549, 1042, 888, 257, 957, 1016, 74, 274, 315, 98, 914, 1026]. resource-exploration [475]. resource-quality-augmenting [222]. resources [1033, 757, 907, 141, 591, 637, 924, 666, 7, 998, 108, 675, 580, 251, 284, 714, 767, 433, 216, 870, 425, 499, 531, 938, 281, 421, 947, 1053]. response [379, 1044, 915, 973, 902, 626, 936]. responses [683, 644, 537, 1002]. restoration [794]. restoring [956]. Restricted [445]. restrictions [186]. result [964]. results [494, 164, 129, 386]. retrofits [139]. retrospective [892]. returns [280, 288]. Réunion [602]. Review [219, 227, 226, 207, 184, 177, 74, 791, 199, 61, 192, 112, 142, 221, 53, 79, 101, 11, 125, 91, 62, 133, 80, 152, 102, 171, 159, 100, 247, 31, 160, 144, 252, 111, 126, 728, 191, 70, 265, 89, 119, 143, 153, 71, 200, 208, 756, 134, 213, 239, 90, 854, 84, 178, 238, 638, 258, 135, 42, 110, 220, 228, 161, 237, 384, 43, 179]. reviews [10, 21, 270]. revisionist [338]. Revisions [78]. revisited [301, 217]. Rhea [472]. rheas [676]. rhinos [853]. rice [875, 818]. Richards [383]. Ricker [932]. Rico [799]. Ridge [552]. Rift [1038, 674]. Right [539]. rights [419, 504]. rights-based [504]. Risk [927, 992, 332, 599, 555, 647, 1015, 979, 174, 830, 582, 455, 872, 511, 20]. risk-averse [647]. risk-aversion [599]. Risk-based [332]. Risks [778, 525, 605]. risky [768]. River [627, 982, 922, 855, 901, 860, 899, 404, 1031, 326, 608, 728, 758, 1004, 182, 1051]. RMA [198]. Robert [232]. Roberts [807]. Robust [598, 996, 827, 985]. robustness [603]. rocky [410]. role [713, 746, 496, 488, 753, 990, 582, 235, 930]. root [588]. rotation [535, 1036, 555, 780, 467]. rotations [444]. Rotifers [480]. roundwood [175]. route [636]. Royalty [529]. rule [358, 625]. rules [189, 641, 51]. Ruminations [436]. run [291, 175, 361, 211, 527, 314]. runoff [326, 1024, 999]. rural [821, 993, 955, 904].

saccharina [817]. safe [882]. Sagar [653]. sales [650]. saline [273]. salmon [703, 611, 321, 702, 527, 556, 289, 138, 404]. Salt [909]. salvage [626]. sample [981]. sampling [1012, 347, 613]. sand [593]. satellite [962]. satisfactory [313]. sativa [818]. saturation [731]. save [796]. saving [865]. Scale [387, 1012, 399, 16, 1010, 855, 902, 941, 490, 546, 743]. scale-free [490]. scales [624]. scallop [271]. scanning [900]. scarce [941]. scarcity [432]. scenario [614]. scenarios [892, 1040, 849, 1025, 952]. schedules [766, 96]. scheduling [719]. scheme [716, 1046]. schemes [678]. Schnute [383]. scholar [232]. school [246]. schools [246]. science [105, 435, 549]. scientists [113]. sea [761, 284, 462, 515, 413, 1007, 911, 523, 729, 246]. seabird [975]. seabirds [602]. seal [684, 702]. seals [1007, 412]. Search [204, 2]. seas [284, 312, 426, 664]. season [58, 78]. seasonal [600, 462]. Second [206, 7]. secondary [421]. sector [942, 786, 150, 87]. sectoral [571]. security [1044, 993, 1032]. sediment [1029, 728]. sedimentation [728]. seedling [818]. Seeking [995]. segmentation [36]. seine [204]. selected [778]. selection [774, 976]. selective [611, 453]. selectivity [1052, 319]. self [553, 396, 390]. self-organizing [390]. self-protection [553]. self-thinning [396]. semantic [886]. semelparous [766]. semi [898, 641, 356, 1002]. semi-arid [898, 641, 1002]. semi-linear [356]. semiarid [940, 922, 752, 1026]. sensing [852]. sensitive [804]. Sensitivity [846, 927, 592, 718, 512]. separability [875]. separating [428]. separation [743]. sequence [17]. Sequential [262, 75]. sequestration [455]. Serengeti [348, 1050]. series [182, 35, 551]. services [519, 773, 774, 595, 964, 887]. set [827, 985]. settlements [821]. sex [141]. share [693]. shared [69]. Sharghonj [1044]. Sharing [514, 312, 627, 873, 678]. sheep [690, 601]. shift [581]. shifts [440]. Shiraz [922]. Shire [510]. shocks [623]. Shooting [853]. shorebird [540]. shoreline [683]. short [291, 175, 527, 18, 853]. short- [853]. short-run [291, 175, 527]. short-term [18]. shortage [849]. should [607]. show [980]. shrimp [892]. SI [194, 849]. Sicily [460]. side [459]. silver [849]. simple [730, 769, 859]. simplified [470, 382]. simplifying [548]. simulated [952]. Simulating [264, 951, 941, 709]. Simulation [818, 1004, 475, 184, 1024, 395, 922, 15, 660, 506, 817]. simulation-tested [506]. Simulations [612, 990]. single [1052, 1017, 569, 838, 964, 732, 852]. single- [1017]. single-species [1052, 569]. single-tree [852]. singletree [39]. singular [350]. Singularly [352, 353]. sink [545]. sinks [319]. SIR [680]. site [774, 681]. sites [550]. size [451, 629, 413, 604, 611, 983, 379, 665, 482, 267, 370, 981, 427, 625]. size-dependent [625]. size-heterogeneous [629]. size-selective [611]. size-structured [451]. slope [624]. sluggish [696]. small [322, 856, 621, 855, 546]. small-scale [855, 546]. smart [707]. Smoky [795, 745, 717]. snow [890, 917]. snowy [649]. social [738, 457, 792, 691, 1009, 771, 938, 638, 887]. social-ecological [792, 1009, 638]. socially [828]. socio [826, 951]. socio-ecological [826, 951]. socioecological [1025]. soft [999]. Software [177, 394, 270, 237, 384, 170, 184, 405]. Soil [593, 1038, 794, 819, 861, 786, 981]. sole [375, 867]. Solow [534, 429]. solution [313, 575, 606]. solutions [297, 567, 1016, 356, 315]. Solving [464, 478, 949]. some [85]. Soummam [944]. source [316]. Sources [319, 760, 497]. South [1044, 940, 512, 902]. South-Eastern [940]. Southeast [26, 476]. Southeastern [578]. Southern [805, 1038, 149, 410, 628, 720, 225]. southwest [863]. southwestern [935]. Space [177, 543, 321, 886, 879, 551]. space-and [321]. spacing [665, 546]. Spain [471, 847]. span [474]. Spatial [934, 541, 1022, 324, 371, 887, 568, 929, 596, 650, 633, 205, 988, 905, 948, 308, 348, 605, 624, 636, 902, 730, 381, 951, 532, 454, 834, 752, 981, 181]. spatially [423, 572]. spatio [559]. spatio-temporal [559]. spatiotemporal [840]. Speakeasy [226]. Special [924, 773, 335, 947, 364, 373, 296, 300, 489, 521, 456, 505, 736, 435, 349, 354, 318, 393, 539, 473, 481, 311, 342, 643, 826, 870, 631, 516]. specialist [659]. species [840, 1052, 811, 1007, 412, 418, 1017, 517, 748, 973, 921, 432, 652, 800, 453, 569, 838, 644, 910, 665, 718, 236, 28, 471, 533, 616, 657, 1019, 549, 306, 732, 257, 723, 762, 334, 781, 1000]. specific [725, 513]. specification [287]. specifications [697]. speed [452, 961]. split [525]. split-stream [525]. Spotted [796, 391]. spread [780, 753, 881, 781]. spruce [305, 363, 954]. squares [660]. squirrel [461]. Sri [805]. stabilities [352]. Stability [760, 66, 299, 574, 210, 254, 328, 538]. stabilization [197]. stabilizing [819]. Stable [862, 117, 306, 487, 604, 809, 518, 815, 1008]. Stackelberg [627, 331]. Stage [177, 243, 725, 766, 39, 868, 376, 771, 837, 472, 404]. stage-specific [725]. stage-structured [766, 39, 837, 472]. stand [444, 75, 244, 1040, 256, 290, 1043]. stand-growth [244]. Standard [1046, 396, 882]. Standardization [18]. standards [926, 215]. stands [39, 872, 732]. state [584, 1044, 543, 544, 276, 857, 313, 551, 527]. state-space [551]. Statement [113]. states [862, 809, 27, 157, 740, 1008, 628, 476, 166, 578, 751]. stationary [107]. Statistical [944, 347, 549, 346]. Steady [544, 276, 313]. steady-state [313]. Stella [184]. stem [780]. steps [472]. sterile [689, 154]. stirred [263]. Stochastic [164, 893, 575, 729, 197, 535, 553, 627, 423, 495, 555, 14, 293, 321, 492, 589, 463, 235, 281, 626, 841, 1008, 888, 181, 493]. stochastic-structure [321]. stochasticity [289, 603]. stock [359, 878, 266, 189, 234, 856, 69, 285, 679, 855, 440, 280, 619, 35, 107, 566, 524, 389, 309, 450]. stock-recruitment [35, 107]. stocking [97, 897]. stockpile [731]. stocks [991, 824, 447, 138]. Stopping [189, 563]. storage [687, 327, 251, 197]. stores [945]. stormwater [1037]. stragedy [131]. strategic [562]. strategies [487, 685, 682, 983, 764, 245, 652, 308, 176, 285, 795, 930, 692, 952, 196]. strategy [761, 154, 910, 797, 646, 871, 872]. stream [525]. streamflow [941]. Strong [989, 85, 502]. Structural [37, 182, 999, 1042]. structurally [583, 732]. structure [444, 547, 595, 321, 402, 636, 388, 482, 280, 844, 1016]. Structured [366, 451, 116, 1052, 878, 141, 243, 325, 85, 766, 483, 39, 367, 318, 194, 810, 837, 472, 770, 560, 726]. structuring [114]. studies [546]. study [291, 689, 918, 1029, 940, 922, 851, 753, 805, 272, 589, 860, 370, 994, 879, 510, 217, 951, 960, 305, 363, 916, 936, 963, 1002]. Studying [954, 802]. stylized [1008]. Sub [1029]. Sub-basin [1029]. subjected [29, 762]. submersed [681]. subsidies [380]. substitutable [675]. substitution [882]. substocks [205, 138]. subtropical [345]. success [501, 224]. successful [937]. successional [147]. supplies [250]. supply [291, 175, 996, 75, 193, 687, 211, 222, 218, 261, 420, 849, 949, 248]. support [907, 673, 828]. supportive [753]. suppression [812]. surface [326, 917, 462]. surveillance [950]. Survey [175, 198, 3]. Survival [451, 652, 344, 802, 797]. Sustainability [926, 292, 773, 507, 424, 502, 503, 823]. Sustainable [996, 260, 599, 520, 358, 935, 911, 985, 1047, 675, 465, 870, 930, 938, 770, 1025, 429, 534, 334]. Suzanne [836]. Swedish [175]. swimmer [933]. switching [682]. SWMM [922]. SWOT [1039]. sylvatic [913]. synchrony [975]. syndrome [968]. synergy [994]. synfuels [558]. synthesis [106, 248]. Systat [168]. system [970, 996, 865, 690, 507, 86, 118, 821, 182, 712, 843, 823, 857, 765, 538, 653, 267, 900, 174, 994, 575, 951, 601, 916, 751, 752, 1025]. Systems [539, 903, 827, 352, 140, 507, 184, 693, 505, 183, 97, 792, 639, 854, 826, 506, 386, 1009, 114, 999, 638, 740, 357, 947, 828].

tailed [913]. Tailoring [695]. take [393]. taking [566]. target [60]. targeting [9, 887]. tariff [765]. tax [658]. taxation [365, 925, 786, 852]. taxes [175, 206, 309, 557, 181]. Taxing [974]. technical [471]. technique [689, 154]. techniques [761, 1039, 805, 614, 854]. Technological [1033, 203, 849, 582]. technologies [222]. technology [687, 545, 216, 691, 620, 269]. Tehran [999, 1037]. telemetry [684]. telescoping [624]. temperature [917, 537, 462, 912, 1003]. Temporal [448, 934, 1022, 559, 948]. temporally [606]. tenant [712]. tenure [993]. term [684, 849, 18, 853]. terminal [389, 1037]. terrestrial [900, 852]. test [16, 173]. tested [935, 506]. testing [941, 448]. tests [644]. their [349, 354, 393, 613, 664]. theorem [85, 288]. theoretic [741, 182, 286, 209, 127, 607, 404]. theoretical [368, 613]. theory [480, 141, 518, 24, 7, 2, 912, 549, 146, 991]. Theragra [515]. there [588]. thinning [396]. threatened [665, 718]. threats [853]. three [935, 973, 821, 868, 863]. three-dimensional [935, 863]. three-stage [868]. threshold [617, 59, 433, 888]. thresholds [536, 985]. tigers [1045]. Tilabad [960]. Timber [491, 420, 291, 211, 261, 26, 236, 660, 448, 146, 333, 248, 196]. Time [1023, 846, 629, 840, 330, 873, 485, 495, 182, 724, 639, 236, 645, 35, 484, 551, 810, 268, 499, 472, 743, 1002, 842, 989]. time-consistent [873]. time-lag [1002]. time-steps [472]. timing [444, 496, 971, 449, 626]. tip [1009]. Tipping [1009, 995, 928]. Tocantins [919]. tolerance [844]. tool [586, 344, 476, 122]. tools [405]. topic [886]. topsoil [452]. total [280]. tourism [988, 1025]. tourism-based [1025]. tourists [850, 412]. towns [913]. toxicants [421]. toxins [479]. tracheal [194]. trade [935, 536, 609, 812, 508, 863, 601, 876]. trade-off [935, 863, 601]. trade-offs [536, 812]. trading [716]. traditional [542]. tragedy [859]. traits [611, 839]. Transboundary [874, 627, 609, 679, 463, 127, 380, 965]. transcends [801]. transfer [182, 993, 35, 955, 972]. transferable [195, 788, 282]. transformation [457, 751]. transition [794, 282]. transport [593, 785]. transported [979]. trapping [761, 176]. travel [949]. traveling [356]. Treatment [571, 75, 945]. treatments [491]. treaty [501]. treaty-making [501]. tree [164, 366, 662, 546, 852, 427, 884]. trees [585, 96, 583, 963, 290]. trembling [824]. trend [704]. trends [684, 832, 917, 462, 912, 18]. tribal [916]. tribal-forest [916]. tritrophic [915, 574]. Trojan [910]. trophic [446]. tropical [84, 415, 884]. trout [398]. Truth [385]. tsembaga [603]. tuna [204, 149]. Tunisia [940]. turf [775]. turf-managed [775]. Turkey [851]. turtle [462]. tutorial [288]. Two [233, 640, 800, 453, 324, 50, 730, 613, 377, 880, 964, 771, 472, 305, 340, 404]. two-dimensional [305]. two-patch [730, 340]. two-stage [771, 404]. type [116, 650, 973, 879]. Typha [921].

U.S. [491, 186]. uncertain [330, 647, 641, 1053]. uncertainties [990]. Uncertainty [834, 1012, 827, 687, 106, 598, 599, 420, 651, 605, 992, 579, 285, 731, 768, 718, 699, 77, 255, 930, 449, 34, 872, 511, 726, 995, 696, 573, 98]. uncoordinated [774]. Understanding [1017, 748, 395, 505, 497]. undertake [139]. uneven [576]. uneven-aged [576]. ungulate [66, 345]. unified [480, 163]. unifying [747]. unilateral [165]. uniqueness [358, 367]. unit [588, 851, 18]. United [628, 476, 578, 751, 166]. unitization [710]. univariate [551]. unknowingly [425]. unmanaged [339]. unregulated [579]. unstable [862]. Unsteady [584]. unsustainability [714]. unused [266]. ups [696]. urban [590, 950, 1051, 823, 994, 1004]. USA [982]. Use [766, 198, 211, 716, 276, 1051, 728, 624, 1038, 879, 610, 582, 786, 960, 509, 333, 390, 1014, 186, 816]. user [833, 250]. Using [939, 507, 518, 470, 547, 568, 175, 996, 244, 1044, 764, 748, 154, 1022, 464, 1024, 940, 890, 917, 651, 922, 644, 886, 662, 303, 288, 860, 912, 984, 524, 1032, 837, 532, 390, 954, 952, 1004]. usually [362]. Utah [909]. utility [723, 20]. utilization [987, 26, 1026]. utilizing [307].

v3.1 [227]. vaccination [760, 837]. validation [398]. Valley [674, 1038]. valuable [266]. valuation [410, 277, 804]. value [535, 519, 123, 254, 497, 527, 920, 879, 255, 197, 762, 828]. value-laden [828]. valued [551, 828]. values [6, 276]. Valuing [660]. vapor [961]. Varanasi [850, 882]. variability [944, 369, 619]. variable [327, 337, 492, 345, 472]. variables [562]. variance [599]. variation [87]. variations [1038]. varieties [820]. various [818]. Varroa [303]. vary [215, 1025]. varying [848, 901]. VCS [1046]. vector [297, 837]. vegetation [690, 918, 681, 917, 906, 601, 752, 1002]. Verified [1046]. Version [219, 207]. versus [139, 518, 571, 251, 738, 196, 573]. via [660]. Viability [446, 827, 924, 926, 794, 782, 618, 1010, 461]. view [463, 338]. views [157]. vigilance [1000]. violation [648]. viral [632]. virus [837]. visible [1004]. visible-light [1004]. Visual [395]. vitulina [684]. volatile [704]. Vole [659]. Volta [627]. Volterra [38, 357]. Volume [104, 137, 172, 229, 259, 278, 295, 317, 202, 64, 585, 349, 354]. voluntary [608]. VPA [651]. vs [479, 389, 257]. vulnerability [805, 826]. vulnerable [94].

Waiting [883]. Walker [817]. walleye [515, 965]. wan [724]. war [525]. warblers [488]. warming [983, 123, 523, 501, 526, 974, 646]. Washington [527]. waste [124, 550]. wastewater [224]. water [996, 1031, 193, 627, 687, 637, 27, 157, 760, 387, 976, 183, 600, 508, 753, 961, 902, 476, 861, 273, 510, 938, 509, 333, 403, 692, 722, 719]. waterborne [764, 760]. waterfowl [671]. waters [497]. Watershed [1032, 960, 944, 1044, 1022, 922, 1039, 1038, 332, 1014]. wave [190]. waves [356]. way [395]. Weak [502]. web [782, 401]. webs [634]. Weight [376]. Weight-controlled [376]. Welfare [203, 331, 420, 129]. well [263, 613]. well-behaved [613]. well-stirred [263]. West [653, 1037, 837]. western [649, 491, 671]. wetland [886, 906]. wetlands [264, 671]. which [29]. while [853]. white [968, 900]. white-nose [968]. Whither [24]. whole [88, 1043]. whole-forest [88]. whole-stand [1043]. wide [75]. widths [992]. wild [991, 324, 795, 447]. wilderness [106, 255]. Wildlife [707, 343, 632, 260, 655, 536, 470, 464, 868, 344, 802, 1050, 718, 342, 853, 466, 762, 953]. wind [590, 452, 593, 961]. Window [1009]. winged [485, 488, 484]. wise [313]. Within [530, 648, 602]. Within-generational [530]. without [422, 684, 730]. wolf [706]. wood [831, 767]. woodi [194]. woodland [583]. work [560]. working [1011]. World [1035, 643, 335, 442, 499]. Woybo [1014]. WRF [961].

Yangtze [1051, 899]. Year [122]. years [442, 506, 631]. Yield [820, 1029, 1047, 818, 770, 454, 719, 1043]. yields [599, 728, 619].

zone [124, 922, 460]. zoning [1022]. zoonoses [632]. zoonotic [989]. Zürich [903].

References

McKelvey:1986:E

[1] Robert McKelvey and Roland H. Lamberson. Editorial. Natural Resource Modeling, 1(1):1–2, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mangel:1986:STN

[2] Marc Mangel and Colin Clark. Search theory in natural resource modeling. Natural Resource Modeling, 1(1):3–54, Fall 1986. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Reed:1986:OHM

[3] William J. Reed. Optimal harvesting models in forest management — a survey. Natural Resource Modeling, 1(1):55–79, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Barclay:1986:MHP

[4] Hugh J. Barclay. Models of host-parasitoid interactions to determine the optimal instar of parasitization for pest control. Natural Resource Modeling, 1(1):81–103, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cohen:1986:RBN

[5] Yosef Cohen. On the relationships between natural and harvest mortalities of the mallard. Natural Resource Modeling, 1(1):105–110, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fisher:1986:EDO

[6] A. C. Fisher and W. M. Hanemann. Environmental damages and option values. Natural Resource Modeling, 1(1):111–124, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fuller:1986:ERL

[7] David Fuller and Serge Dupont. Exhaustible resources, learning, and the theory of the second best. Natural Resource Modeling, 1(1):125– 150, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McKean:1986:NFN

[8] John R. McKean and Richard G. Walsh. Neoclassical foundations for nonmarket benefits estimation. Natural Resource Modeling, 1(1):153– 170, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McKelvey:1986:ERT

[9] Robert McKelvey. Economic regulation of targeting behavior in multispecies fishery. Natural Resource Modeling, 1(1):171–189, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gulland:1986:BR

[10] John Gulland. Book reviews. Natural Resource Modeling, 1(1):191– 193, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Brauer:1986:BR

[11] Fred Brauer. Book review. Natural Resource Modeling, 1(1):195–197, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1986:CE

[12] Anonymous. Calendar of events. Natural Resource Modeling, 1(1):199– 200, Fall 1986. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Clark:1987:E

[13] Colin W. Clark. Editorial. Natural Resource Modeling, 1(2):201–203, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Goodman:1987:CSD

[14] Daniel Goodman. Consideration of stochastic demography in the design and management of biological reserves. Natural Resource Modeling, 1 (2):205–234, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Loehle:1987:OES

[15] Craig Loehle. Optimizing ecosystem simulation model performance. Natural Resource Modeling, 1(2):235–243, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hof:1987:PMA

[16] John G. Hof and James B. Pickens. A pragmatic multilevel approach to large-scale renewable resource optimization: a test case. Natural Resource Modeling, 1(2):245–264, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Krautkraemer:1987:NOS

[17] Jeffrey A. Krautkraemer. A note on the optimal sequence of extraction. Natural Resource Modeling, 1(2):265–278, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Quinn:1987:SCP

[18] Terrance J. Quinn II. Standardization of catch-per-unit-effort for shortterm trends in catchability. Natural Resource Modeling, 1(2):279–296, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rosenman:1987:FRU

[19] Robert E. Rosenman and Charles H. Whiteman. Fishery regulation under rational expectations and costly dynamic adjustment. Natural Resource Modeling, 1(2):297–320, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Walters:1987:AMH

[20] Carl Walters and Don Ludwig. Adaptive management of harvest rates in the presence of a risk averse utility function. Natural Resource Modeling, 1(2):321–337, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Plant:1987:BR

[21] Richard E. Plant. Book reviews. Natural Resource Modeling, 1(2):339– 340, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1987:CEa

[22] Anonymous. Calendar of events. Natural Resource Modeling, 1(2):341– 342, Spring 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hallam:1987:E

[23] Thomas G. Hallam. Editorial. Natural Resource Modeling, 2(1):1–3, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fisher:1987:WOP

[24] Anthony C. Fisher. Whither oil prices: the evidence from theory. Natural Resource Modeling, 2(1):5–22, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Reed:1987:PFA

[25] William J. Reed. Protecting a forest against fire: optimal protection patterns and harvest policies. Natural Resource Modeling, 2(1):23–53, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lyon:1987:OCM

[26] Kenneth S. Lyon, Roger A. Sedjo, and Bambang P. Adiwiyoto. An optimal control model for analysis of timber resource utilization in Southeast Asia. Natural Resource Modeling, 2(1):55–80, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Clarke:1987:CSD

[27] F. H. Clarke and G. R. Munro. Coastal states, distant water fishing nations and extended jurisdiction: a principal-agent analysis. Natural Resource Modeling, 2(1):81–107, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mesterton-Gibbons:1987:OPC

[28] Michael Mesterton-Gibbons. On the optimal policy for combined harvesting of independent species. Natural Resource Modeling, 2(1):109–134, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Freedman:1987:PDM

[29] H. I. Freedman and J. W.-H. So. Persistence in discrete models of a population which may be subjected to harvesting. Natural Resource Modeling, 2(1):135–145, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1987:Ea

[30] Anonymous. Errata. Natural Resource Modeling, 2(1):147, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hastings:1987:BR

[31] Alan Hastings. Book review. Natural Resource Modeling, 2(1):149–151, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1987:CEb

[32] Anonymous. Calendar of events. Natural Resource Modeling, 2(1):153– 156, Summmer 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ludwig:1987:E

[33] Donald Ludwig. Editorial. Natural Resource Modeling, 2(2):157–158, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Schnute:1987:DUM

[34] Jon Schnute. Data uncertainty, model ambiguity, and model identification. Natural Resource Modeling, 2(2):159–212, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Noakes:1987:TSA

[35] D. Noakes, D. W. Welch, and M. Stocker. A time series approach to stock-recruitment analysis: transfer function noise modelling. Natural Resource Modeling, 2(2):213–233, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Berck:1987:AES

[36] Peter Berck and Stephen G. Cecchetti. Allocative efficiency and the segmentation of exhaustible resource ownership. Natural Resource Modeling, 2(2):235–243, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rosenman:1987:SME

[37] Robert E. Rosenman. Structural modeling of expectations and optimization in a fishery. Natural Resource Modeling, 2(2):245–258, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Brauer:1987:CVI

[38] Fred Brauer. A class of Volterra integral equations arising in delayedrecruitment population models. Natural Resource Modeling, 2(2):259– 278, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Haight:1987:CSS

[39] Robert G. Haight and Wayne M. Getz. A comparison of stage-structured and singletree models for projecting forest stands. Natural Resource Modeling, 2(2):279–298, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Barclay:1987:CMP

[40] Hugh J. Barclay. Combining methods of pest control: complementarity of methods and a guiding principle. Natural Resource Modeling, 2(2): 299–323, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1987:Eb

[41] Anonymous. Errata. Natural Resource Modeling, 2(2):325, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Silvert:1987:BR

[42] William Silvert. Book review. Natural Resource Modeling, 2(2):327– 330, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swierzbinski:1987:BR

[43] Joseph E. Swierzbinski. Book review. Natural Resource Modeling, 2(2): 331–335, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1987:CEc

[44] Anonymous. Calendar of events. Natural Resource Modeling, 2(2):337– 340, Fall 1987. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dobson:1988:E

[45] Andrew Dobson. Editorial. Natural Resource Modeling, 2(3):341–344, Winter 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Nystad:1988:PRM

[46] Arild N. Nystad. Petroleum reservoir management: a reservoir economic approach. Natural Resource Modeling, 2(3):345–382, Winter 1988. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cooke:1988:HPM

[47] Kenneth L. Cooke, Richard Elderkin, and Matthew Witten. Harvesting procedures with management policy in iterative density-dependent population models. Natural Resource Modeling, 2(3):383–420, Winter 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Krautkraemer:1988:RDP

[48] Jeffrey A. Krautkraemer. The rate of discount and the preservation of natural environments. Natural Resource Modeling, 2(3):421–437, Winter 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hanson:1988:OHD

[49] Floyd Hanson and Dennis Ryan. Optimal harvesting with density dependent random effects. Natural Resource Modeling, 2(3):439–455, Winter 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ludwig:1988:CTM

[50] D. Ludwig, C. J. Walters, and Justin Cooke. Comparison of two models and two estimation methods for catch and effort data. Natural Resource Modeling, 2(3):457–498, Winter 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

White:1988:MCC

[51] G. N. White III and P. Mace. Models for cooperation and conspiracy in fisheries: changing the rules of the game. Natural Resource Modeling, 2 (3):499–530, Winter 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1988:E

[52] Anonymous. Errata. Natural Resource Modeling, 2(3):531, Winter 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bohi:1988:BR

[53] Douglas R. Bohi. Book review. Natural Resource Modeling, 2(3):533–536, Winter 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1988:CEa

[54] Anonymous. Calendar of events. Natural Resource Modeling, 2(3):537– 540, Winter 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McKelvey:1988:Ea

[55] Robert McKelvey. Editorial. Natural Resource Modeling, 2(4):535–537, Spring 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cushing:1988:NMM

[56] J. M. Cushing. Nonlinear matrix models and population dynamics. Natural Resource Modeling, 2(4):539–580, Spring 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Allard:1988:IIO

[57] Jacques Allard, Darrel Errico, and William J. Reed. Irreversible investment and optimal forest exploitation. Natural Resource Modeling, 2(4): 581–597, Spring 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Charles:1988:SFM

[58] Anthony T. Charles. In-season fishery management: a Bayesian model. Natural Resource Modeling, 2(4):599–629, Spring 1988. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic). See revisions [78].

Hall:1988:RET

[59] Darwin C. Hall. The regional economic threshold for integrated pest management. Natural Resource Modeling, 2(4):631–652, Spring 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mendoza:1988:DER

[60] Guillermo A. Mendoza. Determination of efficient range of target levels in multiple objective planning. Natural Resource Modeling, 2(4):653–667, Spring 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Baskerville:1988:BR

[61] G. L. Baskerville. Book review. Natural Resource Modeling, 2(4):669–671, Spring 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Collie:1988:BR

[62] Jeremy Collie. Book review. Natural Resource Modeling, 2(4):673–677, Spring 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1988:CEb

[63] Anonymous. Calendar of events. Natural Resource Modeling, 2(4):679– 682, Spring 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1988:IV

[64] Anonymous. Index to volume 2. Natural Resource Modeling, 2(4):683, Spring 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McKelvey:1988:Eb

[65] Robert McKelvey. Editorial. Natural Resource Modeling, 3(1):1, Summmer 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Metzgar:1988:SPM

[66] Lee H. Metzgar and Ernest Boyd. Stability properties in a model of forage-ungulate-predator interactions. Natural Resource Modeling, 3(1): 3–43, Summmer 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Brauer:1988:HPM

[67] Fred Brauer, David Rollins, and A. C. Soudack. Harvesting in population models with delayed recruitment and age-dependent mortality. Natural Resource Modeling, 3(1):45–62, Summmer 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mesterton-Gibbons:1988:OPC

[68] Michael Mesterton-Gibbons. On the optimal policy for combining harvesting of predator and prey. Natural Resource Modeling, 3(1):63–90, Summmer 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kaitala:1988:ORS

[69] V. Kaitala and M. Pohjola. Optimal recovery of a shared resource stock: a differential game model with efficient memory equilibria. Natural Resource Modeling, 3(1):91–119, Summmer 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:1988:BR

[70] Roland H. Lamberson. Book review. Natural Resource Modeling, 3(1): 121–122, Summmer 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McKee:1988:BR

[71] Mac McKee. Book review. Natural Resource Modeling, 3(1):123–127, Summmer 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1988:CEc

[72] Anonymous. Calendar of events. Natural Resource Modeling, 3(1):129– 130, Summmer 1988. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cohen:1989:CC

[73] Yosef Cohen. The challenge of conservation. Natural Resource Modeling, 3(2):131–135, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Williams:1989:RDO

[74] Byron K. Williams. Review of dynamic optimization methods in renewable natural resource management. Natural Resource Modeling, 3 (2):137–216, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Barber:1989:SAD

[75] Richard L. Barber and J. Douglas Brodie. A sequential approach to derivation of stand treatment and forest-wide harvest prescription or supply. Natural Resource Modeling, 3(2):217–239, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Caputo:1989:QCR

[76] Michael R. Caputo. The qualitative content of renewable resource models. Natural Resource Modeling, 3(2):241–259, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mueller:1989:NRE

[77] Michael J. Mueller. Natural resource economics under uncertainty: effect of futures markets. Natural Resource Modeling, 3(2):261–288, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Charles:1989:RSF

[78] Anthony T. Charles. Revisions to “In-season fishery management: a Bayesian model”. Natural Resource Modeling, 3(2):289–290, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic). See [58].

Botsford:1989:BR

[79] Louis W. Botsford. Book review. Natural Resource Modeling, 3(2):291– 292, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cushing:1989:BR

[80] J. M. Cushing. Book review. Natural Resource Modeling, 3(2):293–296, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1989:NCE

[81] Anonymous. Notices and calendar of events. Natural Resource Modeling, 3(2):297–302, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1989:Ea

[82] Anonymous. Errata. Natural Resource Modeling, 3(2):303, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mesterton-Gibbons:1989:E

[83] Michael Mesterton-Gibbons. Errata. Natural Resource Modeling, 3(2): 305, Spring 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Pauly:1989:DFD

[84] Daniel Pauly, Geronimo Silvestre, and Ian R. Smith. On development, fisheries and dynamite: a brief review of tropical fisheries management. Natural Resource Modeling, 3(3):307–329, Summmer 1989. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cushing:1989:SET

[85] J. M. Cushing. A strong ergodic theorem for some nonlinear matrix models for the dynamics of structured populations. Natural Resource Modeling, 3(3):331–357, Summmer 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Freedman:1989:EPR

[86] H. I. Freedman and J. B. Shukla. The effect of a predator resource on a diffusive predator prey system. Natural Resource Modeling, 3(3):359–383, Summmer 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sallnas:1989:MVP

[87] Ola Sallnäs and Ljusk Ola Eriksson. Management variation and price expectations in an intertemporal forest sector model. Natural Resource Modeling, 3(3):385–398, Summmer 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Reed:1989:NLW

[88] William J. Reed and Darrell Errico. A new look at whole-forest modeling. Natural Resource Modeling, 3(3):399–427, Summmer 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mangel:1989:BR

[89] Marc Mangel. Book review. Natural Resource Modeling, 3(3):429–445, Summmer 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Noon:1989:BR

[90] Barry R. Noon. Book review. Natural Resource Modeling, 3(3):447–450, Summmer 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Clark:1989:BR

[91] Colin W. Clark. Book review. Natural Resource Modeling, 3(3):451–453, Summmer 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1989:CEa

[92] Anonymous. Calendar of events. Natural Resource Modeling, 3(3):455– 460, Summmer 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1989:Eb

[93] Anonymous. Errata. Natural Resource Modeling, 3(3):461, Summmer 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Reed:1989:OIP

[94] William J. Reed. Optimal investment in the protection of a vulnerable biological resource. Natural Resource Modeling, 3(4):463–480, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dennis:1989:AEP

[95] Brian Dennis. Allee effects: population growth, critical density, and the chance of extinction. Natural Resource Modeling, 3(4):481–538, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Namkoong:1989:OGR

[96] G. Namkoong and J. Rodriguez. Optimum growth and reproduction schedules for forest trees. Natural Resource Modeling, 3(4):539–551, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Huffaker:1989:DOS

[97] Ray G. Huffaker and James E. Wilen. Dynamics of optimal stocking in plant/herbivore systems. Natural Resource Modeling, 3(4):553–575, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yeung:1989:PCU

[98] David Yeung and John Hartwick. Production cost uncertainty and industry equilibrium for non-renewable resource extraction. Natural Resource Modeling, 3(4):577–588, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1989:EN

[99] Anonymous. Editor’s note. Natural Resource Modeling, 3(4):589, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Frauenthal:1989:BR

[100] James C. Frauenthal. Book review. Natural Resource Modeling, 3(4):591– 593, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Boyce:1989:BR

[101] Mark S. Boyce. Book review. Natural Resource Modeling, 3(4):595– 598, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Derrick:1989:BR

[102] Bill Derrick. Book review. Natural Resource Modeling, 3(4):599–602, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1989:CEb

[103] Anonymous. Calendar of events. Natural Resource Modeling, 3(4):603– 610, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1989:VI

[104] Anonymous. Volume index. Natural Resource Modeling, 3(4):611–612, Fall 1989. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Francis:1990:FSM

[105] Robert C. Francis. Fisheries science and modeling: a look to the future. Natural Resource Modeling, 4(1):1–9, Winter 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Clarke:1990:LDW

[106] Harry R. Clarke and William J. Reed. Land development and wilderness conservation policies under uncertainty: a synthesis. Natural Resource Modeling, 4(1):11–37, Winter 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Parma:1990:OHF

[107] Ana M. Parma. Optimal harvesting of fish populations with nonstationary stock-recruitment relationships. Natural Resource Modeling, 4(1):39–76, Winter 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gross:1990:GER

[108] L. S. Gross and E. C. H. Veendorp. Growth with exhaustible resources and a materials — balance production function. Natural Resource Modeling, 4(1):77–94, Winter 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Howe:1990:NE

[109] Eric C. Howe and Kevin F. Forbes. Negotiating over externalities. Natural Resource Modeling, 4(1):95–119, Winter 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Simberloff:1990:BR

[110] Daniel Simberloff and Robert O. Lawton. Book review. Natural Resource Modeling, 4(1):121–128, Winter 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hood:1990:BR

[111] J. Myron Hood. Book review. Natural Resource Modeling, 4(1):129–130, Winter 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Biles:1990:BR

[112] Charles M. Biles. Book review. Natural Resource Modeling, 4(1):131–132, Winter 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fox:1990:SCS

[113] William W. Fox, Jr. Statement of concerned scientists on the reauthorization of the Magnuson Fishery Conservation and Management Act. Natural Resource Modeling, 4(2):133–142, Spring 1990. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Russell:1990:GMS

[114] Clifford S. Russell. Game models for structuring monitoring and enforcement systems. Natural Resource Modeling, 4(2):143–173, Spring 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rowse:1990:NCE

[115] John Rowse and David Yeung. Nonmalleable capital and efficient allocation of a nonrenewable resource. Natural Resource Modeling, 4(2): 175–196, Spring 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Allen:1990:PAS

[116] Linda J. S. Allen, Michael P. Moulton, and Francis L. Rose. Persistence in an age-structured population for a patch-type environment. Natural Resource Modeling, 4(2):197–214, Spring 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Selgrade:1990:SPB

[117] James F. Selgrade and Gene Namkoong. Stable periodic behavior in a pioneer-climax model. Natural Resource Modeling, 4(2):215–227, Spring 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fried:1990:ALC

[118] Jeremy S. Fried and Margaret S. Torn. Analyzing localized climate impacts with the changed climate fire modeling system. Natural Resource Modeling, 4(2):229–253, Spring 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mangel:1990:BR

[119] Marc Mangel. Book review. Natural Resource Modeling, 4(2):255–271, Spring 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Williams:1990:MMM

[120] Byron K. Williams and James D. Nichols. Modeling and the management of migratory birds. Natural Resource Modeling, 4(3):273–311, Summmer 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xaba:1990:EPO

[121] A. B. Xaba. Equilibrium point optimality for the noy-meir grazing model. Natural Resource Modeling, 4(3):313–325, Summmer 1990. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Polacheck:1990:YAC

[122] Tom Polacheck. Year around closed areas as a management tool. Natural Resource Modeling, 4(3):327–354, Summmer 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cochrane:1990:EPQ

[123] Hal Cochrane. Exploring the properties of quasi-option value in the context of global warming. Natural Resource Modeling, 4(3):355–380, Summmer 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kahn:1990:CPB

[124] Peter B. Kahn and Yair Zarmi. Controlled pollution buffer zone — an approach to waste management. Natural Resource Modeling, 4(3): 381–387, Summmer 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Caraco:1990:BR

[125] Thomas Caraco and Jonathan A. Newman. Book review. Natural Resource Modeling, 4(3):389–393, Summmer 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Horn:1990:BR

[126] Carole L. Horn. Book review. Natural Resource Modeling, 4(3):395–401, Summmer 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Munro:1990:OMT

[127] Gordon R. Munro. The optimal management of transboundary fisheries: game theoretic considerations. Natural Resource Modeling, 4(4):403–426, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Caddy:1990:ORM

[128] J. F. Caddy. Options for the regulation of Mediterranean demersal fisheries. Natural Resource Modeling, 4(4):427–475, Fall 1990. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lufgren:1990:MWG

[129] Karl-Gustaf Lüfgren. The measurement of welfare gains from genetic progress in forestry: general equilibrium and disequilibrium results. Natural Resource Modeling, 4(4):477–494, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Clarke:1990:MRR

[130] Harry R. Clarke. Management of a renewable resource with a finite demand elasticity. Natural Resource Modeling, 4(4):495–519, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Negri:1990:SC

[131] Donald H. Negri. “stragedy” of the commons. Natural Resource Modeling, 4(4):521–537, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Stone:1990:CIL

[132] Nicholas D. Stone. Chaos in an individual-level predator-prey model. Natural Resource Modeling, 4(4):539–553, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Conrad:1990:BR

[133] Jon M. Conrad. Book review. Natural Resource Modeling, 4(4):555– 561, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mullen:1990:BR

[134] Ashley Mullen. Book review. Natural Resource Modeling, 4(4):563–565, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Shukla:1990:BR

[135] J. B. Shukla. Book review. Natural Resource Modeling, 4(4):567, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1990:N

[136] Anonymous. Notices. Natural Resource Modeling, 4(4):569–571, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1990:VI

[137] Anonymous. Volume index. Natural Resource Modeling, 4(4):573–574, Fall 1990. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:1991:EAS

[138] Gordon L. Swartzman. Equilibrium abundance of salmon stocks in a life cycle model with interacting hatchery and natural substocks. Natural Resource Modeling, 5(1):1–18, Winter 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cameron:1991:EAP

[139] Trudy Ann Cameron and Matthew B. Wright. Energy audit programs versus market incentives as inducements to undertake energy conservation retrofits. Natural Resource Modeling, 5(1):19–53, Winter 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Criddle:1991:EAM

[140] Keith R. Criddle and Arthur M. Havenner. An encompassing approach to modeling fishery dynamics: modeling dynamic nonlinear systems. Natural Resource Modeling, 5(1):55–90, Winter 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bergh:1991:PHT

[141] M. O. Bergh. Population and harvesting theory for nonlinear sex and age-structured resources. Natural Resource Modeling, 5(1):91–134, Winter 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Biles:1991:BR

[142] Charles M. Biles. Book review. Natural Resource Modeling, 5(1):135–139, Winter 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mangel:1991:BRa

[143] Marc Mangel. Book review. Natural Resource Modeling, 5(1):141–142, Winter 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hilborn:1991:BR

[144] Ray Hilborn. Book review. Natural Resource Modeling, 5(1):143–144, Winter 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McKelvey:1991:E

[145] Robert McKelvey. Editorial. Natural Resource Modeling, 5(2):145–146, Spring 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Steinkamp:1991:OCT

[146] Eric A. Steinkamp and David R. Betters. Optimal control theory applied to joint production of timber and forage. Natural Resource Modeling, 5 (2):147–160, Spring 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Boyd:1991:MSC

[147] Ernest Boyd. A model for successional change in a grassland ecosystem. Natural Resource Modeling, 5(2):161–189, Spring 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dinar:1991:MRI

[148] Ariel Dinar, Stephen A. Hatchett, and Edna T. Loehman. Modeling regional irrigation decisions and drainage pollution control. Natural Resource Modeling, 5(2):191–212, Spring 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kennedy:1991:OAJ

[149] John O. S. Kennedy and Hanoch Pasternak. Optimal Australian and Japanese harvesting of southern bluefin tuna. Natural Resource Modeling, 5(2):213–238, Spring 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rosenman:1991:IRF

[150] Robert Rosenman. Impacts of recreational fishing on the commercial sector: an empirical analysis of Atlantic mackerel. Natural Resource Modeling, 5(2):239–257, Spring 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:1991:NN

[151] Gordon L. Swartzman. Net notes. Natural Resource Modeling, 5(2): 259–263, Spring 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cushing:1991:BR

[152] J. M. Cushing. Book review. Natural Resource Modeling, 5(2):265–267, Spring 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mangel:1991:BRb

[153] Marc Mangel. Book review. Natural Resource Modeling, 5(2):269–271, Spring 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Erickson:1991:OSE

[154] Cindy L. Erickson and Richard E. Plant. Optimal strategy for eradication of multiple populations using the sterile insect technique. Natural Resource Modeling, 5(3):273–291, Summmer 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Zeitouni:1991:EMG

[155] Naomi Zeitouni. Efficient management of groundwater quality. Natural Resource Modeling, 5(3):293–320, Summmer 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Innes:1991:EPU

[156] Robert Innes, Catherine Kling, and Jonathan Rubin. Emission permits under monopoly. Natural Resource Modeling, 5(3):321–343, Summmer 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Clarke:1991:CSD

[157] F. H. Clarke and G. R. Munro. Coastal states and distant water fishing nations: conflicting views of the future. Natural Resource Modeling, 5(3): 345–369, Summmer 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kleiber:1991:PAC

[158] Pierre Kleiber. AUTODIF — a C++ array language extension. Natural Resource Modeling, 5(3):371–375, Summmer 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Francis:1991:BR

[159] Robert C. Francis. Book review. Natural Resource Modeling, 5(3):377– 379, Summmer 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hastings:1991:BR

[160] Alan Hastings. Book review. Natural Resource Modeling, 5(3):381–383, Summmer 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Starfield:1991:BR

[161] Anthony M. Starfield. Book review. Natural Resource Modeling, 5(3): 385–387, Summmer 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Berck:1991:E

[162] Peter Berck. Editorial. Natural Resource Modeling, 5(4):389–392, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Getz:1991:UAM

[163] Wayne M. Getz. A unified approach to multispecies modeling. Natural Resource Modeling, 5(4):393–421, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Haight:1991:SPM

[164] Robert G. Haight and Thomas P. Holmes. Stochastic price models and optimal tree cutting: results for loblolly pine. Natural Resource Modeling, 5(4):423–443, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Batabyal:1991:EPU

[165] Amitrajeet A. Batabyal. Environmental pollution and unilateral control in an open economy. Natural Resource Modeling, 5(4):445–467, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Porco:1991:MEL

[166] Travis C. Porco. A model of the enzootiology of Lyme disease in the Atlantic northeast of the United States. Natural Resource Modeling, 5 (4):469–505, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:1991:E

[167] Gordon L. Swartzman. Editorial. Natural Resource Modeling, 5(4):507– 508, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dupont:1991:S

[168] Diane Dupont. Systat. Natural Resource Modeling, 5(4):509–513, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Roberts:1991:DD

[169] Jerry Roberts. Data desk. Natural Resource Modeling, 5(4):515–517, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anderson:1991:PDD

[170] James Anderson. DADiSP: a data analysis and display software. Natural Resource Modeling, 5(4):519, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dorfman:1991:BR

[171] Jeffrey H. Dorfman. Book review. Natural Resource Modeling, 5(4):521– 527, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1991:VI

[172] Anonymous. Volume index. Natural Resource Modeling, 5(4):529–530, Fall 1991. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hof:1992:IMA

[173] John Hof, Brian Kent, and Tony Baltic. An iterative multilevel approach to natural resource optimization: a test case. Natural Resource Modeling, 6(1):1–22, Winter 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Moll:1992:MRP

[174] Richard H. H. Moll and John W. Chinneck. Modeling regeneration and pest control alternatives for a forest system in the presence of fire risk. Natural Resource Modeling, 6(1):23–49, Winter 1992. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Aronsson:1992:MES

[175] Thomas Aronsson. the modeling and estimation of the short-run supply of roundwood under nonlinear income taxes: an application using Swedish Survey Data. Natural Resource Modeling, 6(1):51–70, Winter 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Huffaker:1992:OTS

[176] R. G. Huffaker, M. G. Bhat, and S. M. Lenhart. Optimal trapping strategies for diffusing nuisance-beaver populations. Natural Resource Modeling, 6(1):71–97, Winter 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:1992:SRP

[177] Roland Lamberson. Software review of RAMAS — age, stage, space. Natural Resource Modeling, 6(1):99–102, Winter 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Quinn:1992:BR

[178] Terrance J. Quinn II. Book review. Natural Resource Modeling, 6(1): 103–107, Winter 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tuljapurkar:1992:BR

[179] Shripad Tuljapurkar. Book review. Natural Resource Modeling, 6(1): 109–110, Winter 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Quinn:1992:PDE

[180] John Quinn. A partial differential equation model of optimal forest harvesting. Natural Resource Modeling, 6(2):111–138, Spring 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xepapadeas:1992:OTP

[181] A. Xepapadeas. Optimal taxes for pollution regulation: dynamic, spatial and stochastic characteristics. Natural Resource Modeling, 6(2):139–170, Spring 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Havenner:1992:FER

[182] Arthur Havenner and John Tracy. Flooding on the Eel River: system theoretic time series, transfer function time series, and structural model forecasts. Natural Resource Modeling, 6(2):171–190, Spring 1992. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hsu:1992:DPC

[183] Shih-Hsun Hsu and Ronald C. Griffin. A distributed parameter control systems approach to the analysis of water allocation. Natural Resource Modeling, 6(2):191–219, Spring 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Getz:1992:RSI

[184] Wayne Getz. Review of Stella II: a systems simulation software package for Macintosh computers. Natural Resource Modeling, 6(2):221–224, Spring 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ellis:1992:ICE

[185] Gregory M. Ellis. Incentive compatible environmental regulations. Natural Resource Modeling, 6(3):225–256, Summmer 1992. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Watson:1992:OCF

[186] W. D. Watson. Opportunity costs of federal land-use restrictions for U.S. coal markets. Natural Resource Modeling, 6(3):257–284, Summmer 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ninpong:1992:RAPa

[187] R. Ninpong, M. Power, and J. D. Fuller. A rapid approximation for predicting the hydrocarbon discovery rate: part 1. assessing the accuracy. Natural Resource Modeling, 6(3):285–303, Summmer 1992. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ninpong:1992:RAPb

[188] R. Ninpong, M. Power, and J. D. Fuller. A rapid approximation for predicting the hydrocarbon discovery rate: part 2. improving the accuracy. Natural Resource Modeling, 6(3):305–313, Summmer 1992. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Conrad:1992:SRC

[189] Jon M. Conrad. Stopping rules and the control of stock pollutants. Natural Resource Modeling, 6(3):315–327, Summmer 1992. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:1992:PWW

[190] Gordon L. Swartzman. Pv-wave: the wave of the future? Natural Resource Modeling, 6(3):329–332, Summmer 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kaitala:1992:BR

[191] Veijo Kaitala. Book review. Natural Resource Modeling, 6(3):333–335, Summmer 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Batabyal:1992:BR

[192] Amitrajeet A. Batabyal. Book review. Natural Resource Modeling, 6(3): 337–342, Summmer 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Becker:1992:DSC

[193] Nir Becker and K. William Easter. Dynamic supply from common property resource: water diversions from the great lakes. Natural Resource Modeling, 6(4):343–371, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mangum:1992:DAS

[194] W. A. Mangum. A discrete age structured SI model of tracheal mite (Acarapis woodi (Rennie)) infestation of honey bees (Apis mellifera L.) at the colony level. Natural Resource Modeling, 6(4):373–384, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Boyce:1992:ITQ

[195] John R. Boyce. Individual transferable quotas and production externalities in a fishery. Natural Resource Modeling, 6(4):385–408, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Williams:1992:IBT

[196] Jeremy S. Williams and J. C. Nautiyal. Integrating budworm and timber management: comparing preventive versus remedial strategies. Natural Resource Modeling, 6(4):409–433, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tsur:1992:SED

[197] Yacov Tsur and Amos Zemel. Stochastic energy demand and the stabilization value of energy storage. Natural Resource Modeling, 6(4):435– 447, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:1992:RMS

[198] Gordon L. Swartzman. RMA Members’ survey on computer use. Natural Resource Modeling, 6(4):449–450, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Barbier:1992:BR

[199] Edward B. Barbier. Book review. Natural Resource Modeling, 6(4):451– 453, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mesterson-Gibbons:1992:BR

[200] Michael Mesterson-Gibbons. Book review. Natural Resource Modeling, 6(4):455–457, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1992:E

[201] Anonymous. Errata. Natural Resource Modeling, 6(4):459, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Porco:1992:VI

[202] Travis C. Porco. Volume index. Natural Resource Modeling, 6(4):461– 462, Fall 1992. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Aronsson:1993:WMT

[203] Thomas Aronsson and Karl-Gustaf Löfgren. Welfare measurement of technological and environmental externalities in the Ramsey growth model. Natural Resource Modeling, 7(1):1–14, Winter 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Campbell:1993:SBP

[204] H. F. Campbell, G. Meyer, and R. B. Nicholl. Search behavior in the purse seine tuna fishery. Natural Resource Modeling, 7(1):15–35, Winter 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fahrig:1993:EFM

[205] Lenore Fahrig. Effect of fish movement and fleet spatial behavior on management of fish substocks. Natural Resource Modeling, 7(1):37–56, Winter 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Feinerman:1993:SBT

[206] Eli Feinerman and E. Kwan Choi. Second best taxes and quotas in nitrogen regulation. Natural Resource Modeling, 7(1):57–84, Winter 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gatto:1993:RMV

[207] Marino Gatto and Andrea Rizzoli. Review of MATLAB, Version 4.0. Natural Resource Modeling, 7(1):85–88, Winter 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mesterton-Gibbons:1993:BR

[208] Michael Mesterton-Gibbons. Book review. Natural Resource Modeling, 7(1):89–91, Winter 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mesterton-Gibbons:1993:GTR

[209] Michael Mesterton-Gibbons. Game-theoretic resource modeling. Natural Resource Modeling, 7(2):93–147, Spring 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wirl:1993:SPR

[210] Franz Wirl. Stability of the price reaction function when the consumers anticipate future prices. Natural Resource Modeling, 7(2):149–161, Spring 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Binkley:1993:LRT

[211] Clark S. Binkley. Long-run timber supply: price elasticity, inventory elasticity, and the use of capital in timber production. Natural Resource Modeling, 7(2):163–181, Spring 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:1993:NN

[212] Gordon L. Swartzman. Net notes. Natural Resource Modeling, 7(2): 183–185, Spring 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Narain:1993:BR

[213] Urvashi Narain. Book review. Natural Resource Modeling, 7(2):187–189, Spring 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Englin:1993:GCC

[214] Jeffrey Englin and John M. Callaway. Global climate change and optimal forest management. Natural Resource Modeling, 7(3):191–202, Summmer 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Helfand:1993:RED

[215] Gloria E. Helfand. The relative efficiency of different standards when firms vary. Natural Resource Modeling, 7(3):203–217, Summmer 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Norquist:1993:ORN

[216] Suzanne M. Norquist. Optimal R&D for nonrenewable resources: the case of cost reducing technology. Natural Resource Modeling, 7(3):219– 243, Summmer 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Schellberg:1993:PNC

[217] Thomas Schellberg. The problem of nonmalleable capital revisited; a study of the Pacific halibut fishery. Natural Resource Modeling, 7(3): 245–276, Summmer 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dale:1993:FMS

[218] Larry L. Dale and Nancy A. Williams. Forecasting mineral supply from mineral rent data. Natural Resource Modeling, 7(3):277–294, Summmer 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Burke:1993:RCV

[219] Matthew Michael Burke. Review of Cellular, Version 3.0. Natural Resource Modeling, 7(3):295–297, Summmer 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Smith:1993:BR

[220] Paul E. Smith. Book review. Natural Resource Modeling, 7(3):299–300, Summmer 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Biles:1993:BR

[221] Charles M. Biles. Book review. Natural Resource Modeling, 7(3):301–303, Summmer 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Caswell:1993:DRQ

[222] Margriet F. Caswell, David Zilberman, and Gary Casterline. The diffusion of resource-quality-augmenting technologies: output supply and input demand effects. Natural Resource Modeling, 7(4):305–329, Fall 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jin:1993:ECO

[223] Di Jin and Thomas A. Grigalunas. Environmental compliance and optimal oil and gas exploitation. Natural Resource Modeling, 7(4):331–352, Fall 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rozelle:1993:IWC

[224] Scott Rozelle, Xiaoying Ma, and Leonard Ortolano. Industrial wastewater control in Chinese cities: determinants of success in environmental policy. Natural Resource Modeling, 7(4):353–378, Fall 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wear:1993:PPF

[225] David N. Wear and Richard Flamm. Public and private forest disturbance regimes in the Southern Appalachians. Natural Resource Modeling, 7(4):379–397, Fall 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Criddle:1993:RPS

[226] Keith R. Criddle. Review of PC-DOS Speakeasy IV: an interactive matrix programming language. Natural Resource Modeling, 7(4):399–401, Fall 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Criddle:1993:RGV

[227] Keith Criddle. Review of Gauss 386i v3.1: an interactive matrix programming language. Natural Resource Modeling, 7(4):403–405, Fall 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

StMary:1993:BR

[228] Colette M. St. Mary. Book review. Natural Resource Modeling, 7(4):407– 411, Fall 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1993:VI

[229] Anonymous. Volume index. Natural Resource Modeling, 7(4):413–414, Fall 1993. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sherman:1994:NP

[230] Tom Sherman. A note from the publisher. Natural Resource Modeling, 8(1):1–2, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:1994:Ia

[231] Roland H. Lamberson. Introduction. Natural Resource Modeling, 8(1): 3, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:1994:RMM

[232] Roland H. Lamberson. Robert McKelvey, a multi-disciplinary scholar. Natural Resource Modeling, 8(1):5–12, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Berck:1994:TII

[233] Peter Berck. Two informational issues in resource modeling. Natural Resource Modeling, 8(1):13–25, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Conrad:1994:FLP

[234] Jon Conrad and Don Ludwig. Forest land policy: the optimal stock of old-growth forest. Natural Resource Modeling, 8(1):27–45, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Reed:1994:RSM

[235] William J. Reed and Jane J. Ye. The role of stochastic monotonicity in the decision to conserve or harvest old-growth forest. Natural Resource Modeling, 8(1):47–79, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McKelvey:1994:REF

[236] Kevin McKelvey and Roland H. Lamberson. Random entry forestry: timber management in a time of species conservation. Natural Resource Modeling, 8(1):81–93, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:1994:SR

[237] Gordon L. Swartzman. Software review. Natural Resource Modeling, 8 (1):95–97, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rodriguez:1994:BR

[238] Jon Paul Rodriguez. Book review. Natural Resource Modeling, 8(1):99– 102, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Noble:1994:BR

[239] Chris Noble. Book review. Natural Resource Modeling, 8(1):103–105, Winter 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1994:NP

[240] Anonymous. A note from the publisher. Natural Resource Modeling, 8 (2):107–108, Spring 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:1994:Ib

[241] Roland H. Lamberson. Introduction. Natural Resource Modeling, 8(2): 109, Spring 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ludwig:1994:MON

[242] Donald Ludwig. Missed opportunities in natural resource management. Natural Resource Modeling, 8(2):111–117, Spring 1994. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Crowe:1994:OIP

[243] Kathleen M. Crowe and J. M. Cushing. Optimal instar parasitization in a stage structured host-parasitoid model. Natural Resource Modeling, 8 (2):119–138, Spring 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bustamante:1994:ALF

[244] Rodrigo H. Bustamante, Wayne M. Getz, and George M. Branch. Analysis of a limpet fishery using a metaphysiological stand-growth model. Natural Resource Modeling, 8(2):139–161, Spring 1994. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Derrick:1994:OCP

[245] William R. Derrick. Optimal crop planting strategies. Natural Resource Modeling, 8(2):163–176, Spring 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:1994:MDF

[246] Gordon Swartzman, Werner Stuetzle, Kristin Kulman, and Nuan Wen. Modeling the distribution of fish schools in the Bering Sea: morphological school identification. Natural Resource Modeling, 8(2):177–194, Spring 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Getz:1994:BR

[247] Wayne M. Getz. Book review. Natural Resource Modeling, 8(2):195–197, Spring 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wear:1994:ETS

[248] David N. Wear and Peter J. Parks. The economics of timber supply: an analytical synthesis of modeling approaches. Natural Resource Modeling, 8(3):199–223, Summmer 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Caputo:1994:MCP

[249] Michael R. Caputo and Dean Lueck. Modeling common property ownership as a dynamic contract. Natural Resource Modeling, 8(3):225–245, Summmer 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rowse:1994:DCE

[250] John Rowse. Depletion cost effects, user costs and intertemporal supplies of a nonrenewable resource. Natural Resource Modeling, 8(3):247–271, Summmer 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Just:1994:PVG

[251] Richard E. Just, Erik Lichtenberg, and David Zilberman. Partial versus general storage policy: commodities and resources. Natural Resource Modeling, 8(3):273–292, Summmer 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hilborn:1994:BR

[252] Ray Hilborn. Book review. Natural Resource Modeling, 8(3):293–294, Summmer 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1994:E

[253] Anonymous. Errata. Natural Resource Modeling, 8(3):295, Summmer 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cushing:1994:NRV

[254] J. M. Cushing and Zhou Yicang. The net reproductive value and stability in matrix population models. Natural Resource Modeling, 8(4):297–333, Fall 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Reed:1994:CBA

[255] William J. Reed and Jane J. Ye. Cost-benefit analysis applied to wilderness preservation — option value uncertainty and ditonicity. Natural Resource Modeling, 8(4):335–372, Fall 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Plantinga:1994:OFS

[256] Andrew J. Plantinga and Richard A. Birdsey. Optimal forest stand management when benefits are derived from carbon. Natural Resource Modeling, 8(4):373–387, Fall 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tsur:1994:ESN

[257] Yacov Tsur and Amos Zemel. Endangered species and natural resource exploitation: extinction vs. coexistence. Natural Resource Modeling, 8 (4):389–413, Fall 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Shelton:1994:BR

[258] Peter A. Shelton. Book review. Natural Resource Modeling, 8(4):415– 417, Fall 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1994:VI

[259] Anonymous. Volume index. Natural Resource Modeling, 8(4):419–420, Fall 1994. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bevers:1995:SFM

[260] Michael Bevers, John Hof, Brian Kent, and Martin G. Raphael. Sustainable forest management for optimizing multispecies wildlife habitat: a coastal douglas-fir example. Natural Resource Modeling, 9(1):1–23, Winter 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gong:1995:OAI

[261] Peichen Gong. An optimization approach to identifying timber supply function coefficients. Natural Resource Modeling, 9(1):25–50, Winter 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hannesson:1995:SFC

[262] Röugnvaldur Hannesson. Sequential fishing: cooperative and noncooperative equilibria. Natural Resource Modeling, 9(1):51–59, Winter 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Heinricher:1995:AOC

[263] A. Heinricher, S. Lenhart, and A. Solomon. The application of optimal control methodology to a well-stirred bioreactor. Natural Resource Modeling, 9(1):61–80, Winter 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Parks:1995:SCE

[264] Peter J. Parks, Randall A. Kramer, and Ralph E. Heimlich. Simulating cost-effective wetlands reserves: a comparison of positive and normative approaches. Natural Resource Modeling, 9(1):81–96, Winter 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ludwig:1995:BR

[265] Donald Ludwig. Book review. Natural Resource Modeling, 9(1):97–98, Winter 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Clarke:1995:ODW

[266] Harry Clarke. Optimal depletion when development makes an unused resource stock more valuable. Natural Resource Modeling, 9(2):99–119, Spring 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McGarvey:1995:EDS

[267] Richard McGarvey. The effect of decreasing system size on birth and death models of open-access fisheries and predator-prey ecosystems. Natural Resource Modeling, 9(2):121–146, Spring 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rowse:1995:COA

[268] John Rowse. Computing optimal allocations for discrete-time nonlinear natural resource models. Natural Resource Modeling, 9(2):147–175, Spring 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sunding:1995:FTC

[269] David Sunding, David Zilberman, Gordon Rausser, and Alan Marco. Flexible technology and the cost of improving groundwater quality. Natural Resource Modeling, 9(2):177–192, Spring 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Scheel:1995:SR

[270] David Scheel. Software reviews. Natural Resource Modeling, 9(2):193– 195, Spring 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Benford:1995:MNC

[271] Frank A. Benford. A model of the North Carolina bay scallop fishery with endogenous fishing effort and entry. Natural Resource Modeling, 9(3):197–228, Summmer 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Love:1995:RCM

[272] H. Alan Love, Diana M. Burton, Gilbert Sylvia, and Shangli Lei. Regulatory controls and market power exertion: a study of the Pacific halibut industry. Natural Resource Modeling, 9(3):229–253, Summmer 1995. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Plessner:1995:EIS

[273] Yakir Plessner and Eli Feinerman. On the economics of irrigation with saline water: a dynamic analysis. Natural Resource Modeling, 9(3):255– 276, Summmer 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Williams:1995:CPN

[274] Nancy A. Williams and Larry L. Dale. Capacity and price of nonrenewable resource deposits. Natural Resource Modeling, 9(3):277–297, Summmer 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gomez:1995:IMN

[275] Irma A. Gomez, Diana M. Burton, and H. Alan Love. Imputing missing natural resource inventory data and the bootstrap. Natural Resource Modeling, 9(4):299–328, Fall 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Heaps:1995:MUV

[276] Terry Heaps. Multiple use values and optimal steady state age distributions. Natural Resource Modeling, 9(4):329–339, Fall 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Onal:1995:NRV

[277] Hayri Önal and Vijay Muralidaran. Nonmarket resource valuation: a bilevel optimization approach. Natural Resource Modeling, 9(4):341–357, Fall 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1995:VI

[278] Anonymous. Volume index. Natural Resource Modeling, 9(4):359–360, Fall 1995. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sandal:1997:EAF

[279] Leif K. Sandal and Stein I. Steinshamn. Efficient allocation in fisheries: domestic issues. Natural Resource Modeling, 10(1):1–2, Winter 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McDonald:1997:EEE

[280] A. David McDonald, Anthony D. M. Smith, André E. Punt, Geoffrey N. Tuck, and Adam J. Davidson. Empirical evaluation of expected returns from research on stock structure for determination of total allowable catch. Natural Resource Modeling, 10(1):3–29, Winter 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sandal:1997:SFM

[281] Leif K. Sandal and Stein Ivar Steinshamn. A stochastic feedback model for optimal management of renewable resources. Natural Resource Modeling, 10(1):31–52, Winter 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Weninger:1997:ATL

[282] Quinn R. Weninger and Richard E. Just. An analysis of transition from limited entry to transferable quota: non-marshallian principles for fisheries management. Natural Resource Modeling, 10(1):53–83, Winter 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sandal:1997:I

[283] Leif K. Sandal and Stein I. Steinshamn. Introduction. Natural Resource Modeling, 10(2):85–86, Spring 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kaitala:1997:CMH

[284] Veijo Kaitala and Gordon Munro. The conservation and management of high seas fishery resources under the new law of the sea. Natural Resource Modeling, 10(2):87–108, Spring 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lee:1997:HSM

[285] C. S. Lee and B. S. Goh. Harvesting strategies for a migratory fish stock based on a model with uncertainty. Natural Resource Modeling, 10(2): 109–128, Spring 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McKelvey:1997:GTI

[286] Robert McKelvey. Game-theoretic insights into the international management of fisheries. Natural Resource Modeling, 10(2):129–171, Spring 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kim:1997:ASM

[287] C. S. Kim, C. Sandretto, R. A. Fleming, and R. A. Adams. An alternative specification for modeling groundwater dynamics. Natural Resource Modeling, 10(3):173–183, Summmer 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McDonald:1997:TEE

[288] A. David McDonald and Anthony D. M. Smith. A tutorial on evaluating expected returns from research for fishery management using Bayes’ Theorem. Natural Resource Modeling, 10(3):185–216, Summmer 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Quinn:1997:ESE

[289] Terrance J. Quinn II and Richard Gates. Estimation of salmon escapement: models with entry, mortality and stochasticity. Natural Resource Modeling, 10(3):217–250, Summmer 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yin:1997:WCS

[290] Runsheng Yin and David Newman. When to cut a stand of trees? Natural Resource Modeling, 10(3):251–261, Summmer 1997. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Alavalapati:1997:MEI

[291] Janaki R. R. Alavalapati and Martin K. Luckert. Modeling the effect of institutional constraints on short-run timber supply on public land: a case study of quota holders in Alberta. Natural Resource Modeling, 10(4):263–282, Fall 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Armstrong:1997:EFS

[292] Glen W. Armstrong, Frank S. Novak, and Wiktor L. Adamowicz. Environmental and financial sustainability of forest management practices. Natural Resource Modeling, 10(4):283–301, Fall 1997. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hertzler:1997:ESM

[293] Greg Hertzler, Julie Harman, and Robert K. Lindner. Estimating a stochastic model of population dynamics with an application to kangaroos. Natural Resource Modeling, 10(4):303–343, Fall 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ligon:1997:CED

[294] Ethan Ligon and Urvashi Narain. Computing the equilibria of dynamic common property games. Natural Resource Modeling, 10(4):345–369, Fall 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1997:VI

[295] Anonymous. Volume index. Natural Resource Modeling, 10(4):371–372, Fall 1997. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Erbe:1998:ISIa

[296] L. H. Erbe and H. I. Freedman. Introduction to special issue. Natural Resource Modeling, 11(1):1–4, Spring 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Buchanan:1998:DFP

[297] J. Robert Buchanan. Discontinuous forcing of periodic solutions in C^{1} vector fields with applications to population models. Natural Resource Modeling, 11(1):5–19, Spring 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cantrell:1998:PPD

[298] Robert Stephen Cantrell and Chris Cosner. Practical persistence in diffusive food chain models. Natural Resource Modeling, 11(1):21–34, Spring 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Solomonovich:1998:SBE

[299] M. Solomonovich, H. I. Freedman, L. P. Apedaile, S. G. M. Schilizzi, and L. Belostotski. Stability and bifurcations in an environmental recovery model of economic agriculture-industry interactions. Natural Resource Modeling, 11(1):35–79, Spring 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Erbe:1998:ISIb

[300] L. H. Erbe and H. I. Freedman. Introduction to special issue. Natural Resource Modeling, 11(2):81, Summmer 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hastings:1998:NOF

[301] Harold M. Hastings and Robert Kissell. Is the Nile outflow fractal? Hurst’s analysis revisited. Natural Resource Modeling, 11(2):83–93, Summmer 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tam:1998:ALC

[302] Judy Hong Tam. Application of Lanchester combat model in the Ardennes Campaign. Natural Resource Modeling, 11(2):95–116, Summmer 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mangum:1998:APB

[303] W. A. Mangum. Analysis of a population biology model describing the interaction between the parasitic mite, Varroa jacobsoni oudemans and the honey bee, Apis mellifera L., using nonlinear difference equations. Natural Resource Modeling, 11(2):117–130, Summmer 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ruan:1998:DDI

[304] Shigui Ruan. Diffusion-driven instability in the Gierer–Meinhardt model of morphogenesis. Natural Resource Modeling, 11(2):131–141, Summmer 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Singh:1998:NST

[305] Manmohan Singh, Alan Easton, Gurong Cui, and Irina Kozlova. Numerical study of the two-dimensional spruce budworm reaction-diffusion equation with density dependent diffusion. Natural Resource Modeling, 11(2):143–154, Summmer 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sumner:1998:SPB

[306] Suzanne Sumner. Stable periodic behavior in pioneer-climax competing species models with constant rate forcing. Natural Resource Modeling, 11(2):155–171, Summmer 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Becker:1998:CCU

[307] Nir Becker and K. William Easter. Conflict and cooperation in utilizing a common property resource. Natural Resource Modeling, 11(3):173– 196, Fall 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hof:1998:OSD

[308] John Hof. Optimizing spatial and dynamic population-based control strategies for invading forest pests. Natural Resource Modeling, 11(3): 197–216, Fall 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sandal:1998:DCT

[309] Leif K. Sandal and Stein I. Steinshamn. Dynamic corrective taxes with flow and stock externalities: a feedback approach. Natural Resource Modeling, 11(3):217–239, Fall 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Schuhmann:1998:MDF

[310] Peter W. Schuhmann. Modeling dynamics of fishery harvest reallocations: an analysis of the North Carolina red drum fishery. Natural Resource Modeling, 11(3):241–271, Fall 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mesterton-Gibbons:1998:ISI

[311] Michael Mesterton-Gibbons. Introduction to special issue. Natural Resource Modeling, 11(4):273–274, Winter 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kaitala:1998:SBC

[312] Veijo Kaitala and Marko Lindroos. Sharing the benefits of cooperation in high seas fisheries: a characteristic function game approach. Natural Resource Modeling, 11(4):275–299, Winter 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Krawczyk:1998:SSS

[313] Jacek B. Krawczyk, Odile Pourtallier, and Mabel Tidball. A steady-state satisfactory solution to an environmental game with piece-wise defined payoffs. Natural Resource Modeling, 11(4):301–329, Winter 1998. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Schimmelpfennig:1998:LRE

[314] David Schimmelpfennig. Long-run equilibria in an economy with a greenhouse effect. Natural Resource Modeling, 11(4):331–355, Winter 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xepapadeas:1998:ORD

[315] Anastasios Xepapadeas. Optimal resource development and irreversibilities: cooperative and noncooperative solutions. Natural Resource Modeling, 11(4):357–378, Winter 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bhat:1998:TCN

[316] Mahadev G. Bhat, Robert R. Alexander, and Burton C. English. Toward controlling nonpoint source pollution of groundwater: a hierarchical policy formulation game. Natural Resource Modeling, 11(4):379–403, Winter 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:1998:VI

[317] Anonymous. Volume index. Natural Resource Modeling, 11(4):405–406, Winter 1998. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:1999:ISI

[318] Roland H. Lamberson. Introduction to special issue: persistence of structured populations. Natural Resource Modeling, 12(1):1–4, March 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Carroll:1999:SSS

[319] Joseph E. Carroll and Roland H. Lamberson. Sources, sinks and selectivity. Natural Resource Modeling, 12(1):5–36, March 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

VanKirk:1999:EPP

[320] R. W. Van Kirk and M. A. Lewis. Edge permeability and population persistence in isolated habitat patches. Natural Resource Modeling, 12 (1):37–64, March 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Higgins:1999:IAS

[321] Kevin Higgins. The impact of age-, space-and stochastic-structure on the extinction probability of a Chinook salmon metapopulation. Natural Resource Modeling, 12(1):65–108, March 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Conner:1999:EIH

[322] Mary M. Conner and Gary C. White. Effects of individual heterogeneity in estimating the persistence of small populations. Natural Resource Modeling, 12(1):109–127, March 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Delong:1999:HBM

[323] Allison Kimber Delong and Roland H. Lamberson. A habitat based model for the distribution of forest interior nesting birds in a fragmented landscape. Natural Resource Modeling, 12(1):129–146, March 1999. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Keeling:1999:SDT

[324] M. J. Keeling, E. J. Milner-Gulland, and L. M. Clayton. Spatial dynamics of two harvested wild pig populations. Natural Resource Modeling, 12 (1):147–169, March 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Crowe:1999:ISP

[325] Kathleen Crowe. An introduction to structured population dynamics. Natural Resource Modeling, 12(1):171–174, March 1999. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bhat:1999:OCM

[326] Mahadev G. Bhat, K. Renee Fister, and Suzanne Lenhart. An optimal control model for the surface runoff contamination of a large river basin. Natural Resource Modeling, 12(2):175–195, June 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Getz:1999:MPM

[327] Wayne M. Getz and Norman Owen-Smith. A metaphysiological population model of storage in variable environments. Natural Resource Modeling, 12(2):197–230, June 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Helu:1999:IBF

[328] S. Langitoto Helu, James J. Anderson, and David B. Sampson. An individual-based fishery model and assessing fishery stability. Natural Resource Modeling, 12(2):231–247, June 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kim:1999:CGQ

[329] C. S. Kim, Carmen Sandretto, and Donna Lee. Controlling groundwater quality with endogenous regulatory instruments. Natural Resource Modeling, 12(2):249–272, June 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cai:1999:HPU

[330] Deqin Cai. Harvest planning under uncertain extinction time. Natural Resource Modeling, 12(3):273–289, September 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lofgren:1999:WMC

[331] Karl-Gustaf Lofgren. Welfare measurement and cost-benefit analysis in Nash and Stackelberg differential fish games. Natural Resource Modeling, 12(3):291–305, September 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Prato:1999:RBM

[332] Tony Prato. Risk-based multiattribute decision-making in property and watershed management. Natural Resource Modeling, 12(3):307–334, September 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Strange:1999:OCM

[333] Niels Strange, John Douglas Brodie, Henrik Meilby, and Finn Helles. Optimal control of multiple-use products: the case of timber, forage and water production. Natural Resource Modeling, 12(3):335–354, September 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Woodward:1999:OSM

[334] Richard T. Woodward and Richard C. Bishop. Optimal-sustainable management of multi-species fisheries: lessons from a predator-prey model. Natural Resource Modeling, 12(3):355–377, September 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Smith:1999:ISI

[335] Anthony D. M. Smith and A. David McDonald. Introduction to special issue: 1997 World Conference on Natural Resource Modelling. Natural Resource Modeling, 12(4):379–381, December 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Landsberg:1999:MFP

[336] Joe Landsberg and Nicholas C. Coops. Modeling forest productivity across large areas and long periods. Natural Resource Modeling, 12(4): 383–411, December 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lyon:1999:CVN

[337] Kenneth S. Lyon. The costate variable in natural resource optimal control problems. Natural Resource Modeling, 12(4):413–426, December 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mckelvey:1999:CEC

[338] Robert Mckelvey. Coexistence or exclusion in a competitive common-pool fishery: a revisionist view. Natural Resource Modeling, 12(4):427–460, December 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Skonhoft:1999:EUL

[339] Anders Skonhoft. Exploitation of an unmanaged local common. on the problems of overgrazing, regulation and distribution. Natural Resource Modeling, 12(4):461–479, December 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Supriatna:1999:HTP

[340] Asep K. Supriatna and Hugh P. Possingham. Harvesting a two-patch predator-prey metapopulation. Natural Resource Modeling, 12(4):481– 498, December 1999. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:2000:E

[341] Roland H. Lamberson. Editorial. Natural Resource Modeling, 13(1):1–3, March 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mesterton-Gibbons:2000:ISI

[342] Michael P. Mesterton-Gibbons and E. J. Milner-Gulland. Introduction to special issue: conserving African wildlife: modeling issues. Natural Resource Modeling, 13(1):5–11, March 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Norton-Griffiths:2000:WLK

[343] M. Norton-Griffiths. Wildlife losses in Kenya: an analysis of conservation policy. Natural Resource Modeling, 13(1):13–34, March 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hearne:2000:MTE

[344] John W. Hearne and Johan Swart. Modeling: a tool for enhancing the survival prospects of African wildlife. Natural Resource Modeling, 13 (1):35–55, March 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Owen-Smith:2000:MPD

[345] Norman Owen-Smith. Modeling the population dynamics of a subtropical ungulate in a variable environment: rain, cold and predators. Natural Resource Modeling, 13(1):57–87, March 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Walsh:2000:SFM

[346] Peter D. Walsh, Marc Thibault, Yves Mihindou, Daniel Idiata, Christian Mbina, and Lee J. T. White. A statistical framework for monitoring forest elephants. Natural Resource Modeling, 13(1):89–134, March 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Burn:2000:SAS

[347] R. W. Burn and F. M. Underwood. Statistical aspects of sampling populations of forest elephants. Natural Resource Modeling, 13(1):135–150, March 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hofer:2000:MSD

[348] Heribert Hofer, Kenneth L. I. Campbell, Marion L. East, and Sally A. Huish. Modeling the spatial distribution of the economic costs and benefits of illegal game meat hunting in the Serengeti. Natural Resource Modeling, 13(1):151–177, March 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kalachev:2000:ISIa

[349] Leonid V. Kalachev. Introduction to special issues: volume 13, numbers 2 and 3: perturbation methods and their applications. Natural Resource Modeling, 13(2):179–180, June 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Macgillivray:2000:ISP

[350] A. D. Macgillivray. An introduction to singular perturbations. Natural Resource Modeling, 13(2):181–217, June 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wan:2000:LEM

[351] Frederic Y. M. Wan. Locational equilibrium models for publicly owned residential land. Natural Resource Modeling, 13(2):219–246, June 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Butuzov:2000:SPR

[352] V. F. Butuzov, N. N. Nefedov, and K. R. Schneider. Singularly perturbed reaction-diffusion systems in cases of exchange of stabilities. Natural Resource Modeling, 13(2):247–269, June 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ward:2000:DBP

[353] Michael J. Ward. Diffusion and bifurcation problems in singularly perturbed domains. Natural Resource Modeling, 13(2):271–302, June 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kalachev:2000:ISIb

[354] Leonid V. Kalachev. Introduction to special issues: volume 13, numbers 2 and 3: perturbation methods and their applications. Natural Resource Modeling, 13(3):303–304, September 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kalachev:2000:AMA

[355] Leonid V. Kalachev. Asymptotic methods: application to reduction of models. Natural Resource Modeling, 13(3):305–338, September 2000. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Witelski:2000:ATW

[356] Thomas P. Witelski, Kinya Ono, and Tasso J. Kaper. On axisymmetric traveling waves and radial solutions of semi-linear elliptic equations.

Natural Resource Modeling, 13(3):339–388, September 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Shih:2000:CRO

[357] Shagi-Di Shih and Tarek P. Mathew. Computation of the relaxation oscillation period in Lotka–Volterra systems. Natural Resource Modeling, 13(3):389–418, September 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cairns:2000:CHR

[358] Robert D. Cairns and Zhao Yang. The converse of Hartwick’s rule and uniqueness of the sustainable path. Natural Resource Modeling, 13(4): 1–10, December 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Agee:2000:MPC

[359] Mark D. Agee. Monopoly, potential competition and private stock information in exhaustible resource markets. Natural Resource Modeling, 13(4):419–433, December 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anderson:2000:EII

[360] Lee G. Anderson. The effects of ITQ implementation: a dynamic approach. Natural Resource Modeling, 13(4):435–470, December 2000. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bar-Shira:2000:LRI

[361] Ziv Bar-Shira and Israel Finkelshtain. The long-run inefficiency of blockrate pricing. Natural Resource Modeling, 13(4):471–492, December 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rowse:2000:DRN

[362] John Rowse. Does a renewable natural resource usually have many nearoptimal allocation paths? Natural Resource Modeling, 13(4):503–533, December 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Singh:2000:NSS

[363] Manmohan Singh, Alan Easton, and Irina Kozlova. A numerical study of the spruce budworm reaction-diffusion equation with hostile boundaries. Natural Resource Modeling, 13(4):535–549, December 2000. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Arino:2001:ISIa

[364] Ovide Arino and Rafael Bravo de la Parra. Introduction to special issue: Alcalá First International Conference on Mathematical Ecology (AICME). Natural Resource Modeling, 14(1):1–3, March 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

GomezGomez:2001:OET

[365] Carlos Mario Gómez Gómez. On optimal environmental taxation and enforcement: information, monitoring and efficiency. Natural Resource Modeling, 14(1):5–30, March 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Karev:2001:SMP

[366] Georgy P. Karev. Structured models of pattern formation of tree populations. Natural Resource Modeling, 14(1):31–43, March 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kraev:2001:EUH

[367] Egor A. Kraev. Existence and uniqueness for height structured hierarchical population models. Natural Resource Modeling, 14(1):45–70, March 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lischke:2001:NDF

[368] H. Lischke. New developments in forest modeling: convergence between applied and theoretical approaches. Natural Resource Modeling, 14(1): 71–102, March 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Loffler:2001:IIV

[369] T. J. Löffler and H. Lischke. Incorporation and influence of variability in an aggregated forest model. Natural Resource Modeling, 14(1):103–137, March 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Montero:2001:SPO

[370] J. A. Montero. A study of the profitability for an optimal control problem when the size of the domain changes. Natural Resource Modeling, 14(1): 139–146, March 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Picard:2001:SPI

[371] Nicolas Picard, Avner Bar-Hen, and Alain Franc Engref. Spatial pattern induced by asymmetric competition: a modeling approach. Natural Resource Modeling, 14(1):147–175, March 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cirita:2001:POC

[372] Sergio Rinaldi Cirita, Carlo Piccardi, and Alessandra Gragnani. Pest outbreaks control: the approach of peak-to-peak dynamics. Natural Resource Modeling, 14(1):177–195, March 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Arino:2001:ISIb

[373] Ovide Arino and Rafael Bravo de la Parra. Introduction to special issue: Alcalæ First International Conference on Mathematical Ecology (AICME). Natural Resource Modeling, 14(2):197–198, June 2001. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Braumann:2001:CEC

[374] Carlos A. Braumann. Constant effort and constant quotafishing policies with cut-offs in a random environment. Natural Resource Modeling, 14 (2):199–232, June 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Horwood:2001:PBE

[375] Joseph Horwood. Population biology and ecology of the sole. Natural Resource Modeling, 14(2):233–256, June 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Pardo:2001:WCR

[376] Olivier Pardo and Ovide Arino. Weight-controlled recruitment of the anchovy in the late larval stage. Natural Resource Modeling, 14(2):257– 286, June 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

RaISsi:2001:FBM

[377] N. Raïssi. Features of bioeconomics models for the optimal management of a fishery exploited by two different fleets. Natural Resource Modeling, 14(2):287–310, June 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Touzeau:2001:RFL

[378] Suzanne Touzeau and Jean-Luc GouzË. Regulation of a fishery: from a local optimal control problem to an “invariant domain” approach. Natural Resource Modeling, 14(2):311–333, June 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cantrell:2001:EDS

[379] Robert Stephen Cantrell and Chris Cosner. Effects of domain size on the persistence of populations in a diffusive food-chain model with Beddington–DeAngelis functional response. Natural Resource Modeling, 14(3):335–367, September 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Quinn:2001:ESE

[380] John Quinn and Gorazd Ruseski. Effort subsidies and entry deterrence in transboundary fisheries. Natural Resource Modeling, 14(3):369–389, September 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sanchirico:2001:DSE

[381] James N. Sanchirico and James E. Wilen. Dynamics of spatial exploitation: a metapopulation approach. Natural Resource Modeling, 14(3): 391–418, September 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sandal:2001:SFA

[382] Leif K. Sandal and Stein I. Steinshamn. A simplified feedback approach to optimal resource management. Natural Resource Modeling, 14(3):419– 432, September 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yuancai:2001:FAB

[383] Lei Yuancai, Carlos Pacheco Marques, and João Manuel Bento. Features and applications of [von] Bertalanffy–Richards’ and Schnute’s growth equations. Natural Resource Modeling, 14(3):433–451, September 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:2001:SRJ

[384] Gordon L. Swartzman. Software review — Java. Natural Resource Modeling, 14(3):453–456, September 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2001:TM

[385] Anonymous. Truth in modelling. Natural Resource Modeling, 14(3): 457–463, September 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Railsback:2001:GRP

[386] Steven F. Railsback. Getting “results”: the pattern-oriented approach to analyzing natural systems with individual-based models. Natural Resource Modeling, 14(3):465–475, September 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dinar:2001:SEW

[387] Ariel Dinar. Scale and equity in water resource development: a Nash bargaining model. Natural Resource Modeling, 14(4):477–494, December 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Li:2001:FDP

[388] Chao Li and Hugh J. Barclay. Fire disturbance patterns and forest age structure. Natural Resource Modeling, 14(4):495–521, December 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Routledge:2001:MSV

[389] Rick Routledge. Mixed-stock vs. terminal fisheries: a bioeconomic model. Natural Resource Modeling, 14(4):523–539, December 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Strange:2001:LUO

[390] Niels Strange, Henrik Meilby, and Peter Bogetoft. Land use optimization using self-organizing algorithms. Natural Resource Modeling, 14(4):541– 574, December 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yin:2001:SOP

[391] Runsheng Yin. Spotted owl protection, booming housing market, and log price changes in the Pacific Northwest. Natural Resource Modeling, 14(4):575–592, December 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Weckerly:2001:MPM

[392] Floyd W. Weckerly. Matrix population models: construction analysis and interpretation. Natural Resource Modeling, 14(4):593–595, December 2001. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:2002:ISI

[393] Roland H. Lamberson. Introduction to special issue: what does it take to make individual-based models realize their potential? Natural Resource Modeling, 15(1):1–4, March 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ropella:2002:SEC

[394] Glen E. Ropella, Steven F. Railsback, and Stephen K. Jackson. Software engineering considerations for individual-based models. Natural Resource Modeling, 15(1):5–22, March 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Grimm:2002:VDW

[395] Volker Grimm. Visual debugging: a way of analyzing, understanding and communicating bottom-up simulation models in ecology. Natural Resource Modeling, 15(1):23–38, March 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Berger:2002:TSI

[396] Uta Berger, Hanno Hildenbrandt, and Volker Grimm. Towards a standard for the individual-based modeling of plant populations: selfthinning and the field-of-neighborhood approach. Natural Resource Modeling, 15(1):39–54, March 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anderson:2002:ABE

[397] James J. Anderson. An agent-based event driven foraging model. Natural Resource Modeling, 15(1):55–82, March 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Railsback:2002:PLA

[398] Steven F. Railsback, Bret C. Harvey, Roland H. Lamberson, Derek E. Lee, Nathan J. Claasen, and Shuzo Yoshihara. Population-level analysis and validation of an individual-based cutthroat trout model. Natural Resource Modeling, 15(1):83–110, March 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Harper:2002:MMC

[399] Steven J. Harper, James D. Westervelt, and Ann-Marie Shapiro. Modeling the movements of cowbirds: application towards management at the landscape scale. Natural Resource Modeling, 15(1):111–131, March 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Adami:2002:IME

[400] C. Adami. Ab initio modeling of ecosystems with artificial life. Natural Resource Modeling, 15(1):133–145, March 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:2002:IBM

[401] Roland H. Lamberson. Individual-based models on the web. Natural Resource Modeling, 15(1):147–148, March 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kraemer:2002:CMD

[402] Michael A. Kraemer, Leonid V. Kalachev, and Dean W. Coble. A class of models describing age structure dynamics in a natural forest. Natural Resource Modeling, 15(2):149–200, June 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sunding:2002:MCR

[403] David Sunding, David Zilberman, Richard Howitt, Ariel Dinar, and Neal Macdougall. Measuring the costs of reallocating water from agriculture: a multi-model approach. Natural Resource Modeling, 15(2):201–225, June 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Thompson:2002:TSI

[404] William L. Thompson and Danny C. Lee. A two-stage informationtheoretic approach to modeling landscape-level attributes and maximum recruitment of Chinook salmon in the Columbia River Basin. Natural Resource Modeling, 15(2):227–257, June 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Swartzman:2002:NST

[405] Gordon L. Swartzman. Any new software tools? Natural Resource Modeling, 15(2):259–260, June 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sumaila:2002:EMM

[406] Ussif Rashid Sumaila and Anthony T. Charles. Economic models of marine protected areas: an introduction. Natural Resource Modeling, 15(3):261–272, September 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hannesson:2002:EMR

[407] RÖGnvaldur Hannesson. The economics of marine reserves. Natural Resource Modeling, 15(3):273–290, September 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sanchirico:2002:IMR

[408] James N. Sanchirico and James E. Wilen. The impacts of marine reserves on limited-entry fisheries. Natural Resource Modeling, 15(3):291–310, September 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anderson:2002:BAM

[409] Lee G. Anderson. A bioeconomic analysis of marine reserves. Natural Resource Modeling, 15(3):311–334, September 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hall:2002:CVM

[410] Darwin C. Hall, Jane V. Hall, and Steven N. Murray. Contingent valuation of marine protected areas: Southern California rocky intertidal ecosystems. Natural Resource Modeling, 15(3):335–368, September 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Holland:2002:IMP

[411] Daniel S. Holland. Integrating marine protected areas into models for fishery assessment and management. Natural Resource Modeling, 15(3): 369–386, September 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Boncoeur:2002:FFS

[412] Jean Boncoeur, Frédérique Alban, Olivier Guyader Ifremer, and Olivier Thébaud Ifremer. Fish, fishers, seals and tourists: economic consequences of creating a marine reserve in a multi-species, multi-activity context. Natural Resource Modeling, 15(4):387–411, December 2002. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Beattie:2002:MBE

[413] Alasdair Beattie, Ussif Rashid Sumaila, Villy Christensen, and Daniel Pauly. A model for the bioeconomic evaluation of marine protected area size and placement in the North Sea. Natural Resource Modeling, 15(4): 413–437, December 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sumaila:2002:MPA

[414] Ussif Rashid Sumaila. Marine protected area performance in a model of the fishery. Natural Resource Modeling, 15(4):439–451, December 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rodwell:2002:MTM

[415] Lynda D. Rodwell, Edward B. Barbier, Callum M. Roberts, and Tim R. Mcclanahan. A model of tropical marine reserve-fishery linkages. Natural Resource Modeling, 15(4):453–486, December 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Roberts:2002:FBF

[416] Callum M. Roberts and Helen Sargant. Fishery benefits of fully protected marine reserves: why habitat and behavior are important. Natural Resource Modeling, 15(4):487–507, December 2002. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Badger:2003:GMP

[417] Lee Badger. A global model of population-resource interaction. Natural Resource Modeling, 16(1):1–19, March 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bulte:2003:MEI

[418] Erwin H. Bulte and Richard Damania. Managing ecologically interdependent species. Natural Resource Modeling, 16(1):21–38, March 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Caputo:2003:NRE

[419] Michael R. Caputo and Dean Lueck. Natural resource exploitation under common property rights. Natural Resource Modeling, 16(1):39–67, March 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gong:2003:TSU

[420] Peichen Gong and Karl-Gustaf Löfgren. Timber supply under demand uncertainty: welfare gains from perfect competition with rational expectations. Natural Resource Modeling, 16(1):69–97, March 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Shukla:2003:MEP

[421] J. B. Shukla, A. K. Agrawal, P. Sinha, and B. Dubey. Modeling effects of primary and secondary toxicants on renewable resources. Natural Resource Modeling, 16(1):99–120, March 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cairns:2003:MD

[422] Robert D. Cairns. Maximin without discounting. Natural Resource Modeling, 16(2):121–126, June 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cripe:2003:EIS

[423] Greg Cripe. The effect of information on a stochastic, spatially distributed fishery. Natural Resource Modeling, 16(2):127–143, June 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hartwick:2003:NIS

[424] John M. Hartwick. “net investment” and sustainability. Natural Resource Modeling, 16(2):145–160, June 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Petith:2003:PCG

[425] Howard Petith. The possibility of continuous growth with exhaustible resources: unknowingly an agreement has been reached, but it may not be correct. Natural Resource Modeling, 16(2):161–173, June 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Pintassilgo:2003:CAM

[426] Pedro Pintassilgo. A coalition approach to the management of high seas fisheries in the presence of externalities. Natural Resource Modeling, 16 (2):175–197, June 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wichmann:2003:FTS

[427] Lars Wichmann. Forecasting tree size distribution by numerical approximation of the forward Kolmogorov equation. Natural Resource Modeling, 16(2):199–212, June 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Withagen:2003:MDS

[428] Cees Withagen, Geir B. Asheim, and Wolfgang Buchholz. Maximin, discounting, and separating hyperplanes. Natural Resource Modeling, 16 (2):213–217, June 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Withagen:2003:SPS

[429] Cees Withagen, Geir B. Asheim, and Wolfgang Buchholz. On the sustainable program in Solow’s model. Natural Resource Modeling, 16(2): 219–231, June 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic). See erratum [534].

Brauer:2003:PEP

[430] Fred Brauer and David A. Sánchez. Periodic environments and periodic harvesting. Natural Resource Modeling, 16(3):233–244, September 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Clarke:2003:IBC

[431] Harry Clarke. International biodiversity conservation agreements. Natural Resource Modeling, 16(3):245–257, September 2003. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Eichner:2003:ISA

[432] Thomas Eichner and Rüdiger Pethig. The impact of scarcity and abundance in food chains on species population dynamics. Natural Resource Modeling, 16(3):259–303, September 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

NAeVdal:2003:ORN

[433] Eric Nævdal. Optimal regulation of natural resources in the presence of irreversible threshold effects. Natural Resource Modeling, 16(3):305–333, September 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:2003:E

[434] Rollie Lamberson. Editorial. Natural Resource Modeling, 16(4):335–336, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ianelli:2003:ISI

[435] James Ianelli and Roland H. Lamberson. Introduction to special issue: modeling in fisheries science, past, present and future. Natural Resource Modeling, 16(4):337–340, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Quinn:2003:RDF

[436] Terrance J. Quinn II. Ruminations on the development and future of population dynamics models in fisheries. Natural Resource Modeling, 16(4):341–392, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Schnute:2003:DFM

[437] Jon T. Schnute. Designing fishery models: a personal adventure. Natural Resource Modeling, 16(4):393–413, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Stefansson:2003:IMM

[438] Gunnar Stefansson. Issues in multispecies models. Natural Resource Modeling, 16(4):415–437, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

PopeVi:2003:GAM

[439] John Pope Vi. Golden ages or magic moments? Natural Resource Modeling, 16(4):439–464, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Maunder:2003:PSF

[440] Mark N. Maunder. Paradigm shifts in fisheries stock assessment: from integrated analysis to Bayesian analysis and back again. Natural Resource Modeling, 16(4):465–475, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Prager:2003:GAN

[441] Michael H. Prager and Erik H. Williams. From the golden age to the new industrial age: fishery modeling in the early 21st century. Natural Resource Modeling, 16(4):477–489, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Clark:2003:MWY

[442] William G. Clark. A model for the world: 80 years of model development and application at the International Pacific Halibut Commission. Natural Resource Modeling, 16(4):491–503, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2003:E

[443] Anonymous. Errata. Natural Resource Modeling, 16(4):505, December 2003. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Amacher:2004:FRS

[444] Gregory S. Amacher, Erkki Koskela, and Markku Ollikainen. Forest rotations and stand interdependency: ownership structure and timing of decisions. Natural Resource Modeling, 17(1):1–43, March 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lindroos:2004:RCM

[445] Marko Lindroos. Restricted coalitions in the management of regional fisheries organizations. Natural Resource Modeling, 17(1):45–69, March 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mullon:2004:VMT

[446] Christian Mullon, Philippe Cury, and Lynne Shannon. Viability model of trophic interactions in marine ecosystems. Natural Resource Modeling, 17(1):71–102, March 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Murray:2004:MHM

[447] Alexander G. Murray. Managing harvesting to minimize the impact of epidemics on wild fish stocks. Natural Resource Modeling, 17(2):103– 121, June 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Prestemon:2004:TAT

[448] Jeffrey P. Prestemon, John M. Pye, and Thomas P. Holmes. Temporal aggregation and testing for timber price behavior. Natural Resource Modeling, 17(2):123–162, June 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Saphores:2004:EUT

[449] Jean-Daniel M. Saphores. Environmental uncertainty and the timing of environmental policy. Natural Resource Modeling, 17(2):163–190, June 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Schuhmann:2004:MDF

[450] Peter W. Schuhmann and Kurt A. Schwabe. Modeling the dynamics of fishery stock recovery: implications for the assessment of natural resource damages. Natural Resource Modeling, 17(2):191–212, June 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ackleh:2004:SFQ

[451] Azmy S. Ackleh and Keng Deng. Survival of the fittest in a quasilinear size-structured population model. Natural Resource Modeling, 17(3):213– 228, September 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dubey:2004:MEH

[452] B. Dubey. Models for the effect of high speed wind on the depletion of fertile topsoil. Natural Resource Modeling, 17(3):229–249, September 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kar:2004:SHT

[453] Tapan Kumar Kar and K. S. Chaudhuri. On selective harvesting of two competing fish species in the presence of environmental fluctuation. Natural Resource Modeling, 17(3):251–272, September 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Smith:2004:FYC

[454] Martin D. Smith. Fishing yield, curvature and spatial behavior: implications for modeling marine reserves. Natural Resource Modeling, 17(3): 273–298, September 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Stainback:2004:MCR

[455] G. Andrew Stainback and Janaki R. R. A. Lavalapati. Modeling catastrophic risk in economic analysis of forest carbon sequestration. Natural Resource Modeling, 17(3):299–317, September 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hearne:2004:ENS

[456] John Hearne and Yiannis Matsinos. Editorial note on special issue. Natural Resource Modeling, 17(4):319–320, December 2004. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kizos:2004:TLM

[457] Thanasis Kizos and Ioannis Spilanis. The transformation of landscape: modeling policy and social impacts on the agricultural landscape of Lesvos. Natural Resource Modeling, 17(4):321–358, December 2004. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Koutsias:2004:FOP

[458] Nikos Koutsias, Kostas D. Kalabokidis, and Britta Allgöwer. Fire occurrence patterns at landscape level: beyond positional accuracy of ignition points with kernel density estimation methods. Natural Resource Modeling, 17(4):359–375, December 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lyon:2004:NRE

[459] Kenneth S. Lyon and Dug Man Lee. Nonrenewable resource extractions with a pollution side effect: a comparative dynamic analysis. Natural Resource Modeling, 17(4):377–392, December 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mardle:2004:BIE

[460] S. Mardle, C. James, C. Pipitone, and M. Kienzle. Bioeconomic interactions in an established fishing exclusion zone: the Gulf of Castellammare, NW Sicily. Natural Resource Modeling, 17(4):393–422, December 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Matsinos:2004:IVS

[461] Yiannis G. Matsinos and Eleni Papadopoulou. Investigating the viability of squirrel populations; a modeling approach for the Island of Lesvos, Greece. Natural Resource Modeling, 17(4):423–444, December 2004. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mazaris:2004:MES

[462] Antonios D. Mazaris, Eyagelia Kornaraki, Yiannis G. Matsinos, and Dimitrios Margaritoulis. Modeling the effect of sea surface temperature on sea turtle nesting activities by investigating seasonal trends. Natural Resource Modeling, 17(4):445–465, December 2004. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mcdonald:2004:IBC

[463] A. David Mcdonald, Leif K. Sandal, L. Richard Little, Arne-Christian Lund, and Stein-Ivar Steinshamn. Inferring a biopolitical consensus view of stochastic dynamics for management of a transboundary fishery. Natural Resource Modeling, 17(4):467–487, December 2004. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fonnesbeck:2005:SDW

[464] Christopher J. Fonnesbeck. Solving dynamic wildlife resource optimization problems using reinforcement learning. Natural Resource Modeling, 18(1):1–40, March 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mcausland:2005:LDP

[465] Carol Mcausland. Learning by doing in the presence of an open access renewable resource: is growth sustainable? Natural Resource Modeling, 18(1):41–68, March 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Skonhoft:2005:CWB

[466] Anders Skonhoft and Claire W. Armstrong. Conservation of wildlife. a bio-economic model of a wildlife reserve under the pressure of habitat destruction and harvesting outside the reserve. Natural Resource Modeling, 18(1):69–90, March 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Stollery:2005:CCO

[467] Kenneth R. Stollery. Climate change and optimal rotation in a flammable forest. Natural Resource Modeling, 18(1):91–112, March 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Beladi:2005:FIP

[468] Hamid Beladi and Reza Oladi. Foreign investment policies and environment. Natural Resource Modeling, 18(2):113–126, June 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Danielsson:2005:EEL

[469] Asgeir Danielsson. Entry and exit of labor and capital in a fishery. Natural Resource Modeling, 18(2):127–156, June 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Flaaten:2005:URP

[470] Ola Flaaten and Einar Mjølhus. Using reserves to protect fish and wildlife simplified modeling approaches. Natural Resource Modeling, 18(2):157– 182, June 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Orea:2005:MMP

[471] Luis Orea, Antonio Alvarez, and Catherine J. Morrison Paul. Modeling and measuring production processes for a multi-species fishery: alternative technical efficiency estimates for the Northern Spain hake fishery. Natural Resource Modeling, 18(2):183–213, June 2005. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Simoy:2005:SSM

[472] Maria Verónica Simoy and Graciela Ana Canziani. A stage-structured model with two time-steps for analyzing the population dynamics of Rhea americana under variable environmental conditions. Natural Resource Modeling, 18(2):215–233, June 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lenhart:2005:ENSa

[473] Suzanne Lenhart and Horst Thieme. Editorial note on special issue. Natural Resource Modeling, 18(3):235–236, September 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Caliendo:2005:ODC

[474] Frank Caliendo and Kenneth S. Lyon. Optimal discounting in control problems that span multiple generations. Natural Resource Modeling, 18(3):237–259, September 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fitzpatrick:2005:NAS

[475] Ben G. Fitzpatrick. Numerical analysis and simulation of resourceexploration models. Natural Resource Modeling, 18(3):261–279, September 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Paudel:2005:DOW

[476] Krishna P. Paudel, Ashutosh S. Limaye, L. Upton Hatch, James F. Cruise, and Fuad Musleh. Development of an optimal water allocation decision tool for the major crops during the water deficit period in the Southeast United States. Natural Resource Modeling, 18(3):281–306, September 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Salinas:2005:CMH

[477] RenÉ A. Salinas, Suzanne Lenhart, and Louis J. Gross. Control of a metapopulation harvesting model for black bears. Natural Resource Modeling, 18(3):307–321, September 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Joshi:2005:SCP

[478] Hem Raj Joshi, Suzanne Lenhart, and Maïtine Bergounioux. Solving a crop problem by an optimal control method. Natural Resource Modeling, 18(3):323–346, September 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Medvinsky:2005:MIP

[479] Alexander B. Medvinsky, Maria M. Gonik, Vassili V. Velkov, Bai-Lian Li, and Horst Malchow. Modeling invasion of pests resistant to Bt toxins produced by genetically modified plants: recessive vs. dominant invaders. Natural Resource Modeling, 18(3):347–362, September 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Beres:2005:RHU

[480] Karl A. Beres, Robert L. Wallace, and Hendrik H. Segers. Rotifers and Hubbell’s unified neutral theory of biodiversity and biogeography. Natural Resource Modeling, 18(3):363–376, September 2005. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lenhart:2005:ENSb

[481] Suzanne Lenhart and Horst R. Thieme. Editorial note on special issue. Natural Resource Modeling, 18(4):377–378, December 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Martcheva:2005:MMD

[482] Maia Martcheva and Horst R. Thieme. A metapopulation model with discrete size structure. Natural Resource Modeling, 18(4):379–413, December 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Feng:2005:DAS

[483] Zhilan Feng, Libin Rong, and Robert K. Swihart. Dynamics of an agestructured metapopulation model. Natural Resource Modeling, 18(4): 415–440, December 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Phillips:2005:HPO

[484] Karl W. Phillips, Smruti P. Damania, James L. Hayward, Shandelle M. Henson, and Clara J. Logan. Habitat patch occupancy dynamics of glaucous-winged gulls (larus glaucescens) i: a discrete-time model. Natural Resource Modeling, 18(4):441–468, December 2005. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Damania:2005:HPO

[485] Smruti P. Damania, Karl W. Phillips, Shandelle M. Henson, and James L. Hayward. Habitat patch occupancy dynamics of glaucouswinged gulls (larus glaucescens) ii: a continuous-time model. Natural Resource Modeling, 18(4):469–499, December 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Conrad:2005:OAE

[486] Jon M. Conrad. Open access and extinction of the passenger pigeon in North America. Natural Resource Modeling, 18(4):501–519, December 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Apaloo:2005:ICS

[487] J. Apaloo. Inaccessible continuously stable strategies. Natural Resource Modeling, 18(4):521–535, December 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Koonce:2005:RJD

[488] Alexander E. Koonce. The role of juvenile dispersal in the replacement of golden-winged warblers by blue-winged warblers. Natural Resource Modeling, 18(4):537–547, December 2005. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fray:2006:ENS

[489] Bob Fray. Editorial note on special issue. Natural Resource Modeling, 19(1):1–2, March 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Reed:2006:BIS

[490] William J. Reed. A brief introduction to scale-free networks. Natural Resource Modeling, 19(1):3–14, March 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Abt:2006:TMF

[491] Karen L. Abt and Jeffrey P. Prestemon. Timber markets and fuel treatments in the Western U.S. Natural Resource Modeling, 19(1):15–43, March 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lyon:2006:CVS

[492] Kenneth S. Lyon and Saket Pande. The costate variable in a stochastic renewable resource model. Natural Resource Modeling, 19(1):45–66, March 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xiang:2006:MEN

[493] Xiaorong Xiang, Yingping Huang, Gregory Madey, Steve Cabaniss, Leilani Arthurs, and Patricia Maurice. Modeling the evolution of natural organic matter in the environment with an agent-based stochastic approach. Natural Resource Modeling, 19(1):67–90, March 2006. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Chen:2006:CFC

[494] Wenjun Chen, Richard H. Moll, Brian D. Haddon, Sylvain Leblanc, Goran Pavlic, Robert Fraser, Richard Fernandes, Rasim Latfovic, Josef Cihlar, and Simon R. J. Bridge. Canada’s forest cover indicator: definition, methodology and results. Natural Resource Modeling, 19(1):91–116, March 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Emmert:2006:PED

[495] Keith E. Emmert and Linda J. S. Allen. Population extinction in deterministicand stochastic discrete-time epidemic models with periodic coefficients with applications to amphibian populations. Natural Resource Modeling, 19(2):117–164, June 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Finus:2006:PIE

[496] Michael Finus and Bianca Rundshagen. Participation in international environmental agreements: the role of timing and regulation. Natural Resource Modeling, 19(2):165–200, June 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jin:2006:EVE

[497] Di Jin, Porter Hoagland, and Andrew Solow. The economic value of environmental research in understanding the relative contributions of sources of nutrients to coastal waters. Natural Resource Modeling, 19 (2):201–219, June 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Reithe:2006:MRM

[498] Siv Reithe. Marine reserves as a measure to control bycatch problems: the importance of multispecies interactions. Natural Resource Modeling, 19(2):221–242, June 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rowse:2006:HTD

[499] John Rowse. On hyperbolic time discounting in exhaustible resource models: an application to world oil resources. Natural Resource Modeling, 19(2):243–277, June 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Eichner:2006:MPP

[500] Thomas Eichner and Rüdiger Pethig. A microfoundation of predatorprey dynamics. Natural Resource Modeling, 19(3):279–321, September 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Eyckmans:2006:CFG

[501] Johan Eyckmans and Michael Finus. Coalition formation in a global warming game: how the design of protocols affects the success of environmental treaty-making. Natural Resource Modeling, 19(3):323–358, September 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hediger:2006:WSS

[502] Werner Hediger. Weak and strong sustainability, environmental conservation and economic growth. Natural Resource Modeling, 19(3):359–394, September 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hellwig:2006:AAS

[503] Klaus Hellwig. An axiomatic approach to sustainability. Natural Resource Modeling, 19(3):395–403, September 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Holland:2006:MRB

[504] Daniel S. Holland and Kurt E. Schnier. Modeling a rights-based approach for managing habitat impacts of fisheries. Natural Resource Modeling, 19(3):405–435, September 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hearne:2006:ENS

[505] John Hearne. Editorial note on special issue: managing and understanding complex systems. Natural Resource Modeling, 19(4):437–440, December 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Punt:2006:FPA

[506] AndrÉ E. Punt. The FAO precautionary approach after almost 10 years: have we progressed towards implementing simulation-tested feedbackcontrol management systems for fisheries management? Natural Resource Modeling, 19(4):441–464, December 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Elmahdi:2006:USD

[507] Amgad Elmahdi, Hector Malano, and Shahbaz Khan. Using a system dynamics approach to model sustainability indicators for irrigation systems in Australia. Natural Resource Modeling, 19(4):465–481, December 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kirby:2006:CBC

[508] Mac Kirby, M. Ejaz Qureshi, Mohammed Mainuddin, and Brenda Dyack. Catchment behavior and counter-cyclical water trade: an integrated model. Natural Resource Modeling, 19(4):483–510, December 2006. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Smajgl:2006:QEW

[509] Alexander Smajgl. Quantitative evaluation of water use benefits — an integrative modeling approach for the Great Barrier Reef Region. Natural Resource Modeling, 19(4):511–538, December 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Roebeling:2006:EGB

[510] Peter C. Roebeling. Efficiency in Great Barrier Reef water pollution control: a case study for the Douglas Shire. Natural Resource Modeling, 19(4):539–556, December 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tahvonen:2006:OHF

[511] Olli Tahvonen and Markku Kallio. Optimal harvesting of forest age classes under price uncertainty and risk aversion. Natural Resource Modeling, 19(4):557–585, December 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mcarthur:2006:MDG

[512] Lynne Mcarthur, John Boland, and Fleur Tiver. A model to detect grazing sensitivity of myoporum platycarpum in the arid rangelands of South Australia. Natural Resource Modeling, 19(4):587–607, December 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Stage:2006:OHA

[513] Jesper Stage. Optimal harvesting in an age-class model with age-specific mortalities: an example from Namibian linefishing. Natural Resource Modeling, 19(4):609–631, December 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hannesson:2006:SNA

[514] RÖGnvaldur Hannesson. Sharing the Northeast Arctic cod: possible effects of climate change. Natural Resource Modeling, 19(4):633–654, December 2006. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Battaile:2006:DDE

[515] Brian C. Battaile and Terrance J. Quinn II. A DeLury depletion estimator for walleye pollock (Theragra chalco-gramma) in the Eastern Bering Sea. Natural Resource Modeling, 19(4):655–674, December 2006. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tschirhart:2007:ISI

[516] John Tschirhart. Introduction to special issue: integrated modeling of economies and ecosystems. Natural Resource Modeling, 20(1):1–6, March 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Christiaans:2007:MLC

[517] Thomas Christiaans, Thomas Eichner, and Rüdiger Pethig. A micro-level ‘consumer approach’ to species population dynamics. Natural Resource Modeling, 20(1):7–59, March 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Finnoff:2007:UOT

[518] David Finnoff and John Tschirhart. Using oligopoly theory to examine individual plant versus community optimization and evolutionary stable objectives. Natural Resource Modeling, 20(1):61–85, March 2007. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Baumgartner:2007:IVB

[519] Stefan Baumgärtner. The insurance value of biodiversity in the provision of ecosystem services. Natural Resource Modeling, 20(1):87–127, March 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bradshaw:2007:EEM

[520] Corey J. A. Bradshaw and Barry W. Brook. Ecological-economic models of sustainable harvest for an endangered but exotic megaherbivore in Northern Australia. Natural Resource Modeling, 20(1):129–156, March 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hannesson:2007:ISI

[521] Rögnvaldur Hannesson. Introduction to special issue: economic effects of climate change on fisheries. Natural Resource Modeling, 20(2):157– 162, June 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Arnason:2007:CCF

[522] Ragnar Arnason. Climate change and fisheries: assessing the economic impact in Iceland and Greenland. Natural Resource Modeling, 20(2):163– 197, June 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Eide:2007:EIG

[523] Arne Eide. Economic impacts of global warming: the case of the Barents Sea fisheries. Natural Resource Modeling, 20(2):199–221, June 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rockmann:2007:REB

[524] Christine Röckmann, Uwe A. Schneider, Michael A. St. John, and Richard S. J. Tol. Rebuilding the Eastern Baltic cod stock under environmental change — a preliminary approach using stock, environmental, and management constraints. Natural Resource Modeling, 20(2):223–262, June 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Golubtsov:2007:IIS

[525] Peter V. Golubtsov and Robert Mckelvey. The incomplete-information split-stream fish war: examining the implications of competing risks. Natural Resource Modeling, 20(2):263–300, June 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hannesson:2007:GWF

[526] RÖGnvaldur Hannesson. Global warming and fish migrations. Natural Resource Modeling, 20(2):301–319, June 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kaje:2007:VSR

[527] Janne H. Kaje and Daniel D. Huppert. The value of short-run climate forecasts in managing the coastal Coho salmon (Oncorhynchus kisutch) fishery in Washington State. Natural Resource Modeling, 20(2):321–349, June 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Potapov:2007:OCB

[528] Alexei B. Potapov, Mark A. Lewis, and David C. Finnoff. Optimal control of biological invasions in lake networks. Natural Resource Modeling, 20(3):351–379, September 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kunce:2007:RIG

[529] Mitch Kunce. Royalty incentives and Gulf of Mexico oil production. Natural Resource Modeling, 20(3):381–404, September 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Chivers:2007:WGD

[530] W. J. Chivers, R. D. Herbert, and W. Gladstone. Within-generational and diversity-dependent effects in an individual-based model of predatorprey interaction. Natural Resource Modeling, 20(3):405–413, September 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ruseski:2007:HFD

[531] Gorazd Ruseski and John Quinn. Human fertility decisions and common property resources: a dynamic analysis. Natural Resource Modeling, 20(3):415–433, September 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Skabar:2007:MSD

[532] Andrew Skabar. Modeling the spatial distribution of mineral deposits using neural networks. Natural Resource Modeling, 20(3):435–450, September 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Paudel:2007:AMA

[533] Krishna P. Paudel, Michael A. Dunn, Doleswar Bhandari, Richard P. Vlosky, and Kurt M. Guidry. Alternative methods to analyze the rank ordered data: a case of invasive species control. Natural Resource Modeling, 20(3):451–471, September 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Withagen:2007:ESP

[534] Cees Withagen, Geir B. Asheim, and Wolfgang Buchholz. Erratum for “On the sustainable program in Solow’s model”. Natural Resource Modeling, 20(3):473–475, September 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic). See [429].

Alvarez:2007:FRP

[535] Luis H. R. Alvarez and Erkki Koskela. The forest rotation problem with stochastic harvest and amenity value. Natural Resource Modeling, 20(4): 477–509, December 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fenichel:2007:JDE

[536] Eli P. Fenichel and Richard D. Horan. Jointly-determined ecological thresholds and economic trade-offs in wildlife disease management. Natural Resource Modeling, 20(4):511–547, December 2007. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Logan:2007:ATP

[537] J. David Logan and William Wolesensky. Accounting for temperature in predator functional responses. Natural Resource Modeling, 20(4):549– 574, December 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Maiti:2007:PSA

[538] Alakes Maiti, Bibek Patra, and G. P. Samanta. Persistence and stability in aratio-dependent predator-prey system with delay and harvesting. Natural Resource Modeling, 20(4):575–600, December 2007. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:2008:ISI

[539] Roland H. Lamberson and Sharon L. Brown. Introduction to special issue: Getting the details right, modeling for the management of complex systems. Natural Resource Modeling, 21(1):1–2, Spring 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Goss-Custard:2008:IBM

[540] John D. Goss-Custard and Richard A. Stillman. Individual-based models and the management of shorebird populations. Natural Resource Modeling, 21(1):3–71, Spring 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Chivers:2008:SEI

[541] W. J. Chivers, W. Gladstone, and R. D. Herbert. Spatial effects in an individual-based model of producer–herbivore interaction. Natural Resource Modeling, 21(1):72–92, Spring 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kanarek:2008:IBM

[542] Andrew R. Kanarek, Roland H. Lamberson, and Jeffrey M. Black. An individual-based model for traditional foraging behavior: investigating effects of environmental fluctuation. Natural Resource Modeling, 21(1): 93–116, Spring 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Criddle:2008:SSB

[543] Keith R. Criddle and Mark Herrmann. A state space bioeconomic model of Pacific halibut. Natural Resource Modeling, 21(1):117–147, Spring 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hashim:2008:SSI

[544] N. M. Nik Hashim. Steady state and intertemporal management models for marine fisheries. Natural Resource Modeling, 21(1):148–177, Spring 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hennlock:2008:IMP

[545] Magnus Hennlock. An international marine pollutant sink in an asymmetric environmental technology game. Natural Resource Modeling, 21 (1):178–203, Spring 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Strub:2008:IAR

[546] Mike R. Strub and Ralph L. Amateis. Isometric and allometric relationships between large and small-scale tree spacing studies. Natural Resource Modeling, 21(2):205–224, Summmer 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fuller:2008:UNA

[547] Michael M. Fuller, Andreas Wagner, and Brian J. Enquist. Using network analysis to characterize forest structure. Natural Resource Modeling, 21(2):225–247, Summmer 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lawrie:2008:MSL

[548] Jock Lawrie. A method for simplifying large ecosystem models. Natural Resource Modeling, 21(2):248–263, Summmer 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Stauffer:2008:ABS

[549] Howard B. Stauffer. Application of Bayesian statistical inference and decision theory to a fundamental problem in natural resource science: the adaptive management of an endangered species. Natural Resource Modeling, 21(2):264–284, Summmer 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lyon:2008:ODB

[550] Kenneth S. Lyon and Frank Caliendo. Optimal discounting of benefits from cleanup at waste sites. Natural Resource Modeling, 21(2):285–313, Summmer 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rizzardi:2008:SSM

[551] Mark Rizzardi. A state-space model for univariate ordinal-valued time series. Natural Resource Modeling, 21(2):314–329, Summmer 2008. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Shryock:2008:RDE

[552] Kacee A. Shryock, Sharon L. Brown, Nathan J. Sanders, and Elizabeth Burroughs. A reaction–diffusion equation modeling the invasion of the Argentine ant population, Linepithema humile, at Jasper Ridge Biological Preserve. Natural Resource Modeling, 21(2):330–342, Summmer 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ami:2008:FCE

[553] D. Ami, N. Hilgert, S. Pardo, and M. Tidball. Is fishing compatible with environmental conservation: a stochastic model with an element of selfprotection. Natural Resource Modeling, 21(3):343–365, September 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lindroos:2008:CIF

[554] Marko Lindroos. Coalitions in international fisheries management. Natural Resource Modeling, 21(3):366–384, September 2008. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gong:2008:IRA

[555] Peichen Gong and Karl-Gustaf Löfgren. Impact of risk aversion on the optimal rotation with stochastic price. Natural Resource Modeling, 21(3): 385–415, September 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lorentzen:2008:MCC

[556] Torbjørn Lorentzen. Modeling climate change and the effect on the Norwegian salmon farming industry. Natural Resource Modeling, 21(3):416– 435, September 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sandal:2008:OET

[557] Leif K. Sandal and Stein I. Steinshamn. Optimal environmental taxes: effects of pollution decay and consumer awareness. Natural Resource Modeling, 21(3):436–465, September 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wirl:2008:IIS

[558] Franz Wirl. Intertemporal investments into synfuels. Natural Resource Modeling, 21(3):466–488, September 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Heavilin:2008:NMF

[559] Justin Heavilin and James Powell. A novel method of fitting spatiotemporal models to data, with applications to the dynamics of mountain pine beetles. Natural Resource Modeling, 21(4):489–524, Winter 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tahvonen:2008:HAS

[560] Olli Tahvonen. Harvesting an age-structured population as biomass: does it work? Natural Resource Modeling, 21(4):525–550, Winter 2008. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Touza:2008:BAF

[561] Julia Touza, Mette Termansen, and Charles Perrings. A bioeconomic approach to the Faustmann–Hartman model: ecological interactions in managed forest. Natural Resource Modeling, 21(4):551–581, Winter 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ekerhovd:2008:EDS

[562] Nils-Arne Ekerhovd. The effects of different strategic variables in noncooperative fisheries games. Natural Resource Modeling, 21(4):582–606, Winter 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Caputo:2008:IQP

[563] Michael R. Caputo. The intrinsic qualitative properties of the classical optimal stopping problem are invariant to the functional form of the discount function. Natural Resource Modeling, 21(4):607–624, Winter 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Olaussen:2008:EBI

[564] Jon Olaf Olaussen and Anders Skonhoft. On the economics of biological invasion: an application to recreational fishing. Natural Resource Modeling, 21(4):625–653, Winter 2008. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2009:E

[565] Anonymous. Editorial. Natural Resource Modeling, 22(1):iv–v, Spring 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

ROCkmann:2009:REB

[566] Christine Röckmann, Richard S. J. Tol, Uwe A. Schneider, and Michael A. St. John. Rebuilding the Eastern Baltic cod stock under environmental change (Part II): taking into account the costs of a marine protected area. Natural Resource Modeling, 22(1):1–25, Spring 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hunter:2009:NNE

[567] Greg Hunter. Noah’s nonconcavity: on the existence of nontrivial interior solutions to the problem of cost-effective conservation planning. Natural Resource Modeling, 22(1):26–41, Spring 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Aillery:2009:MFS

[568] Marcel Aillery, Noel Gollehon, Vince Breneman, and Shawn Bucholtz. Modeling firm spatial interdependence using national data coverages: a regional application to manure management. Natural Resource Modeling, 22(1):42–66, Spring 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kinzey:2009:MSS

[569] Douglas Kinzey and André E. Punt. Multispecies and single-species models of fish population dynamics: comparing parameter estimates. Natural Resource Modeling, 22(1):67–104, Spring 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Felthoven:2009:MPC

[570] Ronald G. Felthoven, Catherine J. Morrison Paul, and Marcelo Torres. Measuring productivity and its components for fisheries: the case of the Alaskan pollock fishery, 1994–2003. Natural Resource Modeling, 22(1): 105–136, Spring 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fishman:2009:TVP

[571] Yaron Fishman, Nir Becker, and Mordechai Shechter. Treatment versus prevention of nitrogen fertilizer pollution: an inter-sectoral externality policy model. Natural Resource Modeling, 22(1):137–171, Spring 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ding:2009:OHS

[572] Wandi Ding and Suzanne Lenhart. Optimal harvesting of a spatially explicit fishery model. Natural Resource Modeling, 22(2):173–211, May 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yamazaki:2009:OVI

[573] Satoshi Yamazaki, Tom Kompas, and R. Quentin Grafton. Output versus input controls under uncertainty: the case of a fishery. Natural Resource Modeling, 22(2):212–236, May 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tchuenche:2009:SAT

[574] Jean M. Tchuenche and Christinah Chiyaka. Stability analysis of a tritrophic food chain model with an adaptive parameter for the predator. Natural Resource Modeling, 22(2):237–256, May 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Potapov:2009:SML

[575] Alex Potapov. Stochastic model of lake system invasion and its optimal control: neurodynamic programming as a solution method. Natural Resource Modeling, 22(2):257–288, May 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tahvonen:2009:OCB

[576] Olli Tahvonen. Optimal choice between even- and uneven-aged forestry. Natural Resource Modeling, 22(2):289–321, May 2009. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Joshi:2009:ODH

[577] Hem Raj Joshi, Guillermo E. Herrera, Suzanne Lenhart, and Michael G. Neubert. Optimal dynamic harvest of a mobile renewable resource. Natural Resource Modeling, 22(2):322–343, May 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Susaeta:2009:MIB

[578] Andres Susaeta, Janaki R. R. Alavalapati, and Douglas R. Carter. Modeling impacts of bioenergy markets on nonindustrial private forest management in the Southeastern United States. Natural Resource Modeling, 22(3):345–369, August 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Johannesen:2009:GMU

[579] Anne B. Johannesen and Anders Skonhoft. Growth and measurement uncertainty in an unregulated fishery. Natural Resource Modeling, 22(3): 370–392, August 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jensen:2009:RPR

[580] Frank Jensen and Lone G. Kronbak. Random penalties and renewable resources: a mechanism to reach optimal landings in fisheries. Natural Resource Modeling, 22(3):393–414, August 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mukhopadhyay:2009:DIS

[581] B. Mukhopadhyay and R. Bhattacharyya. Diffusion induced shift of bifurcation point in a mangrove ecosystem food-chain model with harvesting. Natural Resource Modeling, 22(3):415–436, August 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ranjan:2009:DRA

[582] Ram Ranjan and Sanhita Athalye. Drought resilience in agriculture: the role of technological options, land use dynamics, and risk perception. Natural Resource Modeling, 22(3):437–462, August 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tausch:2009:SBA

[583] Robin J. Tausch. A structurally based analytic model for estimation of biomass and fuel loads of woodland trees. Natural Resource Modeling, 22(4):463–488, November 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Agarwal:2009:USD

[584] Manju Agarwal and Abhinav Tandon. Unsteady state dispersion of air pollutants under the effects of delayed and nondelayed removal mechanisms. Natural Resource Modeling, 22(4):489–510, November 2009. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Diamantopoulou:2009:ANN

[585] Maria J. Diamantopoulou, Elias Milios, Dimitrios Doganos, and Ioannis Bistinas. Artificial neural network modeling for reforestation design through the dominant trees bole-volume estimation. Natural Resource Modeling, 22(4):511–543, November 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hayward:2009:PGH

[586] James L. Hayward, Shandelle M. Henson, Richard D. Tkachuck, Cynthia M. Tkachuck, Brianna G. Payne, and Cassaundra K. Boothby. Predicting gull/human conflicts with mathematical models: a tool for management. Natural Resource Modeling, 22(4):544–563, November 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mullon:2009:MGF

[587] C. Mullon, J-F. Mittaine, O. Thébaud, G. Péron, G. Merino, and M. Barange. Modeling the global fishmeal and fish oil markets. Natural Resource Modeling, 22(4):564–609, November 2009. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jones:2010:TUR

[588] Nina S. Jones. Is there unit root in the nitrogen oxides emissions: a Monte Carlo investigation? Natural Resource Modeling, 23(1):1–21, February 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Masuda:2010:AMC

[589] Tadayoshi Masuda, John F. Yanagida, James E. T. Moncur, and Samir A. El-Swaify. An application of multi-criteria decision making incorporating stochastic production frontiers: a case study of organic coffee production in Kona, Hawai’i. Natural Resource Modeling, 23(1): 22–47, February 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Agarwal:2010:MEU

[590] Manju Agarwal and Abhinav Tandon. Modeling the effects of urban heat islands generated mesoscale wind on air pollution dispersion in a patchy atmosphere. Natural Resource Modeling, 23(1):48–78, February 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Biancardi:2010:CCD

[591] Marta Biancardi. Competition and cooperation in a dynamical model of natural resources. Natural Resource Modeling, 23(1):79–106, February 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hearne:2010:AME

[592] John W. Hearne. An automated method for extending sensitivity analysis to model functions. Natural Resource Modeling, 23(2):107–120, May 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dupac:2010:SED

[593] Mihai Dupac. Soil erosion and degradation based on sand particles transport caused by wind blowing. Natural Resource Modeling, 23(2):121–137, May 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ngoc:2010:CMR

[594] Quach Thi Khanh Ngoc. Creation of marine reserves and incentives for biodiversity conservation. Natural Resource Modeling, 23(2):138–175, May 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hennessy:2010:BES

[595] David A. Hennessy and Harvey Lapan. Buying ecological services: fragmented reserves, core and periphery national park structure, and the agricultural extensification debate. Natural Resource Modeling, 23(2): 176–217, May 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Charles:2010:FMP

[596] Anthony Charles. Fisheries and marine protected areas: a spatial bioeconomic analysis of distributional impacts. Natural Resource Modeling, 23(2):218–252, May 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bezabih:2010:ECC

[597] Mintewab Bezabih and Tobias Gebäck. Environmental change and the contribution of biodiversity to ecosystem adaptation. Natural Resource Modeling, 23(2):253–284, May 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Doole:2010:RMP

[598] Graeme Doole and Ross Kingwell. Robust mathematical programming for natural resource modeling under parametric uncertainty. Natural Resource Modeling, 23(3):285–302, August 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ewald:2010:SYF

[599] Christian-Oliver Ewald and Wen-Kai Wang. Sustainable yields in fisheries: uncertainty, risk-aversion, and mean-variance analysis. Natural Resource Modeling, 23(3):303–323, August 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Khan:2010:ALF

[600] Shahbaz Khan, Dharma Dassanayake, and Hamza F. Gabriel. An adaptive learning framework for forecasting seasonal water allocations in irrigated catchments. Natural Resource Modeling, 23(3):324–353, August 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Skonhoft:2010:BSV

[601] Anders Skonhoft, Gunnar Austrheim, and Atle Mysterud. A bioeconomic sheep–vegetation trade-off model: an analysis of the Nordic sheep farming system. Natural Resource Modeling, 23(3):354–380, August 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dumont:2010:CEE

[602] Y. Dumont, J. C. Russell, V. Lecomte, and M. Le Corre. Conservation of endangered endemic seabirds within a multi-predator context: the Barau’s petrel in Réunion Island. Natural Resource Modeling, 23(3):

381–436, August 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rogers:2010:PPN

[603] Bruce Rogers and David Murillo. Pigs, parties, and noise: how stochasticity impacts the robustness of the tsembaga cultural practices. Natural Resource Modeling, 23(4):437–466, November 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Biancardi:2010:SFS

[604] Marta Biancardi and Andrea Di Liddo. The size of farsighted stable coalitions in a game of pollution abatement. Natural Resource Modeling, 23(4):467–493, November 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Holland:2010:BRI

[605] Daniel S. Holland and Guillermo E. Herrera. Benefits and risks of increased spatial resolution in the management of fishery metapopulations under uncertainty. Natural Resource Modeling, 23(4):494–520, November 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yadav:2010:ODT

[606] R. R. Yadav, Dilip Kumar Jaiswal, Hareesh Kumar Yadav, and Gul Rana. One-dimensional temporally dependent advection–dispersion equation in porous media: analytical solution. Natural Resource Modeling, 23(4):521–539, November 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Pavlova:2010:HRS

[607] Yulia Pavlova. How rapid should emission reduction be? a gametheoretic approach. Natural Resource Modeling, 23(4):540–564, November 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gengenbach:2010:CRA

[608] Michael F. Gengenbach, Hans-Peter Weikard, and Erik Ansink. Cleaning a river: an analysis of voluntary joint action. Natural Resource Modeling, 23(4):565–590, November 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fujiwara:2010:EPT

[609] Kenji Fujiwara. Environmental policy and trade liberalization: the case of transboundary pollution from consumption. Natural Resource Modeling, 23(4):591–609, November 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Punt:2010:PMP

[610] Maarten J. Punt, Hans-Peter Weikard, Rolf A. Groeneveld, Ekko C. Van Ierland, and Jan H. Stel. Planning marine protected areas: a multiple use game. Natural Resource Modeling, 23(4):610–646, November 2010. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bromaghin:2011:MCS

[611] Jeffrey F. Bromaghin, Ryan M. Nielson, and Jeffrey J. Hard. A model of Chinook salmon population dynamics incorporating size-selective exploitation and inheritance of polygenic correlated traits. Natural Resource Modeling, 24(1):1–47, February 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Duffy:2011:SIA

[612] Kevin J. Duffy. Simulations to investigate animal movement effects on population dynamics. Natural Resource Modeling, 24(1):48–60, February 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Parresol:2011:DTW

[613] Bernard R. Parresol. Derivation of two well-behaved theoretical contagion indices and their sampling properties and application for assessing forest landscape diversity. Natural Resource Modeling, 24(1):61–101, February 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Moloney:2011:POT

[614] P. D. Moloney, J. W. Hearne, I. J. Gordon, and S. R. Mcleod. Portfolio optimization techniques for a mixed-grazing scenario for Australia’s rangelands. Natural Resource Modeling, 24(1):102–116, February 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mcgowan:2011:MMA

[615] Conor P. Mcgowan, David R. Smith, John A. Sweka, Julien Martin, James D. Nichols, Richard Wong, James E. Lyons, Lawrence J. Niles, Kevin Kalasz, Jeffrey Brust, Michelle Klopfer, and Braddock Spear. Multispecies modeling for adaptive management of horseshoe crabs and red knots in the Delaware Bay. Natural Resource Modeling, 24(1):117–156, February 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Powers:2011:CSD

[616] Joseph E. Powers and Elizabeth N. Brooks. Characterizing species distributions by productivity and mortality rates in multispecies models. Natural Resource Modeling, 24(2):157–182, May 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

DaSilveiraCosta:2011:IOG

[617] Michel Iskin Da Silveira Costa and Lucas Del Bianco Faria. Induced oscillations generated by protective threshold policies in the management of exploited populations. Natural Resource Modeling, 24(2):183–206, May 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gordillo:2011:ERM

[618] Luis F. Gordillo. The effects of reducing mating likelihood on population viability. Natural Resource Modeling, 24(2):207–221, May 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Myrseth:2011:DAS

[619] Johanna Myrseth, Katja Enberg, Mikko Heino, and Øyvind Fiksen. Do accurate stock estimates increase harvest and reduce variability in fisheries yields? Natural Resource Modeling, 24(2):222–241, May 2011. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Shukla:2011:MDR

[620] J. B. Shukla, Kusum Lata, and A. K. Misra. Modeling the depletion of a renewable resource by population and industrialization: effect of technology on its conservation. Natural Resource Modeling, 24(2):242– 267, May 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gajardo:2011:HEM

[621] Pedro Gajardo, Julio Peña-Torres, and Héctor Ramírez C. Harvesting economic models and catch-to-biomass dependence: the case of small pelagic fish. Natural Resource Modeling, 24(2):268–296, May 2011. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

JUDez:2011:MEP

[622] Lucinio Júdez, Rosario De Andrés, Miguel Ibáñez, and Elvira Urzainqui. A model for estimating premiums to reduce irrigation on farms. Natural Resource Modeling, 24(3):297–315, August 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kvamsdal:2011:ESM

[623] Sturla F. Kvamsdal. Exogenous shocks and marine reserves. Natural Resource Modeling, 24(3):316–334, August 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Johnson:2011:UTS

[624] Penelope Johnson, Elizabeth Fulton, David C. Smith, Gregory P. Jenkins, and Neville Barrett. The use of telescoping spatial scales to capture inshore to slope dynamics in marine ecosystem modeling. Natural Resource Modeling, 24(3):335–364, August 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Zimmermann:2011:OHF

[625] Fabian Zimmermann, Stein Ivar Steinshamn, and Mikko Heino. Optimal harvest feedback rule accounting for the fishing-up effect and sizedependent pricing. Natural Resource Modeling, 24(3):365–382, August 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sims:2011:OTS

[626] Charles Sims. Optimal timing of salvage harvest in response to a stochastic infestation. Natural Resource Modeling, 24(3):383–408, August 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bhaduri:2011:CCC

[627] Anik Bhaduri, Utpal Manna, Edward Barbier, and Jens Liebe. Climate change and cooperation in transboundary water sharing: an application of stochastic Stackelberg differential games in Volta River Basin. Natural Resource Modeling, 24(4):409–444, November 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Huang:2011:EAC

[628] Jin Huang, Bob Abt, Georg Kindermann, and Sujit Ghosh. Empirical analysis of climate change impact on loblolly pine plantations in the Southern United States. Natural Resource Modeling, 24(4):445–476, November 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Araneda:2011:OHT

[629] Marcelo E. Araneda, Juan M. Hernández, and Eucario Gasca-Leyva. Optimal harvesting time of farmed aquatic populations with nonlinear size-heterogeneous growth. Natural Resource Modeling, 24(4):477–513, November 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

DeLaSen:2011:IBH

[630] M. De La Sen and S. Alonso-Quesada. On the impulsive Beverton– Holt equation and extinction conditions in population dynamics. Natural Resource Modeling, 24(4):514–534, November 2011. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Roberts:2012:ISI

[631] Catherine A. Roberts. Introduction to the special issue: 25 years of natural resource modeling: perspectives of then and now. Natural Resource Modeling, 25(1):1–4, February 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Allen:2012:MMV

[632] L. J. S. Allen, V. L. Brown, C. B. Jonsson, S. L. Klein, S. M. Laverty, K. Magwedere, J. C. Owen, and P. Van Den Driessche. Mathematical modeling of viral zoonoses in wildlife. Natural Resource Modeling, 25 (1):5–51, February 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Conrad:2012:NSM

[633] Jon M. Conrad and Martin D. Smith. Nonspatial and spatial models in bioeconomics. Natural Resource Modeling, 25(1):52–92, February 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Getz:2012:BFA

[634] Wayne M. Getz. A biomass flow approach to population models and food webs. Natural Resource Modeling, 25(1):93–121, February 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hilborn:2012:EQM

[635] Ray Hilborn. The evolution of quantitative marine fisheries management 1985–2010. Natural Resource Modeling, 25(1):122–144, February 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lamberson:2012:BBL

[636] Roland H. Lamberson. A brief and biased look at spatial structure in ecological models: a route to individual-based models. Natural Resource Modeling, 25(1):145–167, February 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Booker:2012:EMW

[637] James F. Booker, Richard E. Howitt, Ari M. Michelsen, and Robert A. Young. Economics and the modeling of water resources and policies. Natural Resource Modeling, 25(1):168–218, February 2012. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

SchlUTer:2012:NHM

[638] M. Schlüter, R. R. J. Mcallister, R. Arlinghaus, N. Bunnefeld, K. Eisenack, F. Hölker, E. J. Milner-Gulland, B. Müller, E. Nicholson, M. Quaas, and M. Stöven. New horizons for managing the environment: a review of coupled social-ecological systems modeling. Natural Resource Modeling, 25(1):219–272, February 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Li:2012:JAD

[639] Nianpeng Li and Abdul-Aziz Yakubu. A juvenile-adult discrete-time production model of exploited fishery systems. Natural Resource Modeling, 25(2):273–324, May 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dubey:2012:MET

[640] B. Dubey. Modeling effects of two interacting populations on a renewable resource. Natural Resource Modeling, 25(2):325–363, May 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Quaas:2012:OGM

[641] Martin F. Quaas and Stefan Baumgärtner. Optimal grazing management rules in semi-arid rangelands with uncertain rainfall. Natural Resource Modeling, 25(2):364–387, May 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kaier:2012:EMP

[642] Klaus Kaier. Economic modeling of the persistence of antimicrobial resistance. Natural Resource Modeling, 25(2):388–402, May 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Moll:2012:ISI

[643] Richard H. H. Moll and Daniel E. Lane. Introduction to the special issue 2011 World Conference on Natural Resource Modeling, Ottawa, Canada. Natural Resource Modeling, 25(3):403–408, August 2012. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Leites:2012:PLU

[644] Laura P. Leites, Gerald E. Rehfeldt, Andrew P. Robinson, Nicholas L. Crookston, and Barry Jaquish. Possibilities and limitations of using historic provenance tests to infer forest species growth responses to climate change. Natural Resource Modeling, 25(3):409–433, August 2012. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

MunozCaravaca:2012:FTC

[645] Alain Muñoz Caravaca, Pascal Douillet, Ovel Díaz García, Renaud Fichez, Roberto H. Herrera, Javier Alcántara-Carrió, and Agustín García Rodríguez. Flushing time in the Cienfuegos Bay, Cuba. Natural Resource Modeling, 25(3):434–455, August 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Paudel:2012:GWI

[646] Krishna P. Paudel and L. Upton Hatch. Global warming, impact on agriculture and adaptation strategy. Natural Resource Modeling, 25(3): 456–481, August 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hallmann:2012:FBA

[647] Fanfan W. Hallmann and Gregory S. Amacher. Forest bioenergy adoption for a risk-averse landowner under uncertain emerging biomass market. Natural Resource Modeling, 25(3):482–510, August 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Abusin:2012:NRM

[648] Sana Abusin, Rashid Hassan, and Greg Hertzler. Natural resource modeling: allowing for inconstant probability of detection and frequency measures of violation within dynamic deterrence fishery models. Natural Resource Modeling, 25(3):511–528, August 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Watts:2012:MEP

[649] Christina M. Watts, Jing Cao, Christopher Panza, Christopher Dugaw, Mark Colwell, and Elizabeth A. Burroughs. Modeling the effects of predator exclosures on a western snowy plover population. Natural Resource Modeling, 25(3):529–547, August 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cho:2012:EFA

[650] Seong-Hoon Cho, J. M. Bowker, Roland K. Roberts, Seung Gyu Kim, Dayton M. Lambert, and Neelam C. Poudyal. Effect of forestland availability by ownership type on license sales for hunting: a spatial approach. Natural Resource Modeling, 25(4):549–573, November 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hillary:2012:ACU

[651] R. M. Hillary. Assessing components of uncertainty in VPA abundance and mortality estimates using an alternative exploitation rate-based algorithm. Natural Resource Modeling, 25(4):574–598, November 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Elofsson:2012:OMI

[652] Katarina Elofsson, Göran Bengtsson, and Ing-Marie Gren. Optimal management of invasive species with different reproduction and survival strategies. Natural Resource Modeling, 25(4):599–628, November 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mandal:2012:MND

[653] Sudipto Mandal, Santanu Ray, and Phani Bhusan Ghosh. Modeling nutrient (dissolved inorganic nitrogen) and plankton dynamics at Sagar Island of Hooghly–Matla estuarine system, West Bengal, India. Natural Resource Modeling, 25(4):629–652, November 2012. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2013:IIa

[654] Anonymous. Issue information. Natural Resource Modeling, 26(1):i–ii, February 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Chen:2013:MIW

[655] Wenting Chen and Anders Skonhoft. On the management of interconnected wildlife populations. Natural Resource Modeling, 26(1):1–25, February 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gong:2013:EEB

[656] Peichen Gong, Karl-Gustaf Löfgren, and Ola Rosvall. Economic evaluation of biotechnological progress: the effect of changing management behavior. Natural Resource Modeling, 26(1):26–52, February 2013. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Shukla:2013:MAA

[657] J. B. Shukla, Shyam Sundar, Shivangi, and Ram Naresh. Modeling and analysis of the acid rain formation due to precipitation and its effect on plant species. Natural Resource Modeling, 26(1):53–65, February 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Blake:2013:ITD

[658] Andon J. Blake. Investigating tax distortions: an applied model of petroleum exploration and extraction decisions. Natural Resource Modeling, 26(1):66–90, February 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mukhopadhyay:2013:VPD

[659] B. Mukhopadhyay and R. Bhattacharyya. Vole population dynamics under the influence of specialist and generalist predation. Natural Resource Modeling, 26(1):91–110, February 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mei:2013:VTH

[660] Bin Mei and Michael L. Clutter. Valuing a timber harvest contract as a high-dimensional American call option via least-squares Monte Carlo simulation. Natural Resource Modeling, 26(1):111–129, February 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2013:IIb

[661] Anonymous. Issue information. Natural Resource Modeling, 26(2):i– ii, May 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ma:2013:MIT

[662] Zhihai Ma, Changhui Peng, Weizhong Li, Qiuan Zhu, Weifeng Wang, Xinzhang Song, and Jianwei Liu. Modeling individual tree mortality rates using marginal and random effects regression models. Natural Resource Modeling, 26(2):131–153, May 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Conrad:2013:EF

[663] Jon M. Conrad and Benjamin P. Leard. Equivalencies in the fishery. Natural Resource Modeling, 26(2):154–163, May 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Punt:2013:MPA

[664] Maarten J. Punt, Hans-Peter Weikard, and Ekko C. Van Ierland. Marine protected areas in the high seas and their impact on international fishing agreements. Natural Resource Modeling, 26(2):164–193, May 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Marcot:2013:HBH

[665] Bruce G. Marcot, Martin G. Raphael, Nathan H. Schumaker, and Beth Galleher. How big and how close? habitat patch size and spacing to conserve a threatened species. Natural Resource Modeling, 26(2):194– 214, May 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ewald:2013:DNR

[666] Christian-Oliver Ewald. Derivatives on nonstorable renewable resources: fish futures and options, not so fishy after all. Natural Resource Modeling, 26(2):215–236, May 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ranjan:2013:MMD

[667] Ram Ranjan. Mathematical modeling of drought resilience in agriculture. Natural Resource Modeling, 26(2):237–258, May 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Thanh:2013:BME

[668] Nguyen Viet Thanh. Bioeconomic model of Eastern Baltic cod under the influence of nutrient enrichment. Natural Resource Modeling, 26(2):259– 280, May 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2013:IIc

[669] Anonymous. Issue information. Natural Resource Modeling, 26(3):i–ii, August 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lazkano:2013:PFM

[670] Itziar Lazkano, Linda Nøstbakken, and Raúl Prellezo. Past and future management of a collapsed fishery: the Bay Of Biscay anchovy. Natural Resource Modeling, 26(3):281–304, August 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Withey:2013:ECC

[671] Patrick Withey and G. Cornelis Van Kooten. The effect of climate change on wetlands and waterfowl in Western Canada: incorporating cropping decisions into a bioeconomic model. Natural Resource Modeling, 26(3): 305–330, August 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wang:2013:EDI

[672] Chenggang Wang and Shyam Nair. The economics of deficit irrigation. Natural Resource Modeling, 26(3):331–364, August 2013. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

OLafsson:2013:QMD

[673] Ari Ólafsson, Sveinn Margeirsson, Eyjólfur Ingi Ásgeirsson, Hlynur StefáNsson, Páll Jensson, Runólfur Gudmundsson, and Sigurjón Arason. Quantitative methods for decision support in the Icelandic fishing industry. Natural Resource Modeling, 26(3):365–384, August 2013. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Adongo:2013:OCA

[674] Donald Adongo, K. Renee Fister, Holly Gaff, and David Hartley. Optimal control applied to Rift Valley Fever. Natural Resource Modeling, 26(3): 385–402, August 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Halicka:2013:SGE

[675] Margaréta Halická and Pavol Jurča. On the sustainable growth in an economy with perfectly substitutable exhaustible resources. Natural Resource Modeling, 26(3):403–434, August 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Simoy:2013:IBM

[676] María. Verónica Simoy, Gustavo Javier Fernández, and Graciela Ana Canziani. An individual-based model to estimate the daily energetic cost of greater rheas and its contribution on population recruitment. Natural Resource Modeling, 26(3):435–454, August 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2013:IId

[677] Anonymous. Issue information. Natural Resource Modeling, 26(4):i–ii, November 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mukanjari:2013:EPB

[678] Samson Mukanjari, Birgit Bednar-Friedl, Edwin Muchapondwa, and Precious Zikhali. Evaluating the prospects of benefit sharing schemes in protecting mountain gorillas in Central Africa. Natural Resource Modeling, 26(4):455–479, November 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Liu:2013:CPR

[679] Xiaozi Liu and Mikko Heino. Comparing proactive and reactive management: managing a transboundary fish stock under changing environment. Natural Resource Modeling, 26(4):480–504, November 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Morin:2013:SDE

[680] Benjamin R. Morin, Eli P. Fenichel, and Carlos Castillo-Chavez. SIR dynamics with economically driven contact rates. Natural Resource Modeling, 26(4):505–525, November 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gray:2013:IAP

[681] Brian R. Gray, Mark D. Holland, Feng Yi, and Leigh Ann Harrod Starcevich. Influences of availability on parameter estimates from site occupancy models with application to submersed aquatic vegetation. Natural Resource Modeling, 26(4):526–545, November 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bischi:2013:MEE

[682] Gian-Italo Bischi, Fabio Lamantia, and Davide Radi. Multispecies exploitation with evolutionary switching of harvesting strategies. Natural Resource Modeling, 26(4):546–571, November 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jin:2013:ORS

[683] Di Jin, Andrew D. Ashton, and Porter Hoagland. Optimal responses to shoreline changes: an integrated economic and geological model with application to curved coasts. Natural Resource Modeling, 26(4):572–604, November 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cowles:2013:MPH

[684] Jonathan D. Cowles, Shandelle M. Henson, James L. Hayward, and Matthew W. Chacko. A method for predicting harbor seal (Phoca vitulina) haulout and monitoring long-term population trends without telemetry. Natural Resource Modeling, 26(4):605–627, November 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bedford:2013:EAM

[685] R. A. Bedford, J. W. Hearne, Y. Wang, H. K. Gorfine, and B. Taylor. Evaluating alternative management strategies for abalone. Natural Resource Modeling, 26(4):628–647, November 2013. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2014:IIa

[686] Anonymous. Issue information. Natural Resource Modeling, 27(1):i–ii, February 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bhaduri:2014:IWS

[687] Anik Bhaduri and Utpal Manna. Impacts of water supply uncertainty and storage on efficient irrigation technology adoption. Natural Resource Modeling, 27(1):1–24, February 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kalkuhl:2014:MCR

[688] Matthias Kalkuhl and Ottmar Edenhofer. Managing the climate rent: how can regulators implement intertemporally efficient mitigation policies? Natural Resource Modeling, 27(1):25–60, February 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bhattacharyya:2014:MSP

[689] R. Bhattacharyya and B. Mukhopadhyay. Mathematical study of a pest control model incorporating sterile insect technique. Natural Resource Modeling, 27(1):61–79, February 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

DaRocha:2014:DAV

[690] José-María Da Rocha and Anders Skonhoft. The dynamics of an animal — vegetation system: sheep farming. Natural Resource Modeling, 27(1):

80–103, February 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ranjan:2014:CSC

[691] Ram Ranjan. Combining social capital and technology for drought resilience in agriculture. Natural Resource Modeling, 27(1):104–127, February 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Vasileiou:2014:OPI

[692] Konstantinos Vasileiou, Panagiotis Mitropoulos, and Ioannis Mitropoulos. Optimizing the performance of irrigated agriculture in Eastern England under different water pricing and regulation strategies. Natural Resource Modeling, 27(1):128–150, February 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hatcher:2014:FSS

[693] Aaron Hatcher. Fishery share systems, ITQ markets, and the distribution of rents. Natural Resource Modeling, 27(1):151–162, February 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2014:IIb

[694] Anonymous. Issue information. Natural Resource Modeling, 27(2):i– ii, May 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Johnson:2014:TPC

[695] Fred A. Johnson, Robert M. Dorazio, Traci D. Castellón, Julien Martin, Jay O. Garcia, and James D. Nichols. Tailoring point counts for inference about avian density: dealing with nondetection and availability. Natural Resource Modeling, 27(2):163–177, May 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wirl:2014:RUD

[696] Franz Wirl and Sebastian Caban. A rationalization of ups and downs of oil prices by sluggish demand, uncertainty, and nonconcavity. Natural Resource Modeling, 27(2):178–196, May 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Strong:2014:GEE

[697] Aaron Strong and Matthew E. Oliver. General equilibrium ecosystem modeling with alternative preference specifications. Natural Resource Modeling, 27(2):197–215, May 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jensen:2014:AAF

[698] Frank Jensen, Lone Grønbæk. Kronbak, and Jens Abildtrup. On avoidance activities in fishery enforcement models. Natural Resource Modeling, 27(2):216–234, May 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Miao:2014:LCO

[699] Ruiqing Miao and David A. Hennessy. To learn or to change: optimal R&D investments under uncertainty in the case of climate change. Natural Resource Modeling, 27(2):235–257, May 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Misra:2014:MAD

[700] A. K. Misra, P. K. Tiwari, Ashish Goyal, and J. B. Shukla. Modeling and analysis of the depletion of organic pollutants by bacteria with explicit dependence on dissolved oxygen. Natural Resource Modeling, 27(2):258– 273, May 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2014:IIc

[701] Anonymous. Issue information. Natural Resource Modeling, 27(3):i–ii, August 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Holma:2014:ECI

[702] Maija Holma, Marko Lindroos, and Soile Oinonen. The economics of conflicting interests: Northern Baltic salmon fishery adaption to gray seal abundance. Natural Resource Modeling, 27(3):275–299, August 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Benshoof:2014:PEF

[703] Chris Benshoof and Jungho Baek. Politics, environment, and fisheries: empirical evidence from Pacific salmon fisheries. Natural Resource Modeling, 27(3):300–310, August 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sidneva:2014:EIE

[704] Nina Sidneva and Eric Zivot. Evaluating the impact of environmental policy on the trend behavior of us emissions of nitrogen oxides and volatile organic compounds. Natural Resource Modeling, 27(3):311–337, August 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Biancardi:2014:IEA

[705] Marta Biancardi and Giovanni Villani. International environmental agreements with developed and developing countries in a dynamic approach. Natural Resource Modeling, 27(3):338–359, August 2014. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hochard:2014:GWP

[706] J. Hochard and D. Finnoff. Gray wolf population projection with intraspecific competition. Natural Resource Modeling, 27(3):360–375, August 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jensen:2014:WRP

[707] Frank Jensen, Jette Bredahl Jacobsen, Niels Strange, and Bo Jellesmark Thorsen. Wildlife reserves, populations, and hunting outcome with smart wildlife. Natural Resource Modeling, 27(3):376–395, August 2014. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bjorndal:2014:NMN

[708] Trond Bjørndal and Marko Lindroos. Noncooperative management of the northeast Atlantic cod fishery: a first mover advantage. Natural Resource Modeling, 27(3):396–410, August 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sigurethardottir:2014:NAS

[709] Sigrídur Sigurdardóttir and Lee Schruben. A new approach to simulating fisheries data for policy making. Natural Resource Modeling, 27(3):411– 428, August 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Boyce:2014:EMP

[710] John R. Boyce. The effect of monopsony power on prorationing and unitization regulation of the common pool. Natural Resource Modeling, 27 (3):429–465, August 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2014:IId

[711] Anonymous. Issue information. Natural Resource Modeling, 27(4):i–ii, November 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hritonenko:2014:MLT

[712] Natali Hritonenko, Yuri Yatsenko, Renan-Ulrich Goetz, and Angels Xabadia. Modeling of a landlord — tenant agricultural system in the environmental context. Natural Resource Modeling, 27(4):467–491, November 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Do:2014:RIL

[713] Kim Hang Pham Do and Ariel Dinar. The role of issue linkage in managing noncooperating basins: the case of the Mekong. Natural Resource Modeling, 27(4):492–518, November 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kitti:2014:DEU

[714] Mitri Kitti. Degree of economic unsustainability in the management of renewable resources. Natural Resource Modeling, 27(4):519–547, November 2014. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2015:IIa

[715] Anonymous. Issue information. Natural Resource Modeling, 28(1):i–ii, February 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fan:2015:EUC

[716] Jin Fan, Shanyong Wang, Yanrui Wu, Jun Li, and Dingtao Zhao. Energyuse choices and allowance trading under the personal carbon trading scheme. Natural Resource Modeling, 28(1):1–17, February 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Salinas:2015:IBM

[717] RenÉ A. Salinas, William H. Stiver, Joseph L. Corn, Suzanne Lenhart, Charles Collins, Marguerite Madden, Kurt C. Vercauteren, Brandon B. Schmit, Ellen Kasari, Agricola Odoi, Graham Hickling, and Hamish Mccallum. An individual-based model for feral hogs in great smoky mountains national park. Natural Resource Modeling, 28(1):18–36, February 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Marcot:2015:ASU

[718] Bruce G. Marcot, Peter H. Singleton, and Nathan H. Schumaker. Analysis of sensitivity and uncertainty in an individual-based model of a threatened wildlife species. Natural Resource Modeling, 28(1):37–58, February 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Zengin:2015:MHS

[719] Hayati Zengin, Ünal Asan, Sinan Destan, Murat Engin Ünal, Ahmet Yeşil, Pete Bettinger, and Ahmet Salih Değermenci. Modeling harvest scheduling in multifunctional planning of forests for longterm water yield optimization. Natural Resource Modeling, 28(1):59–85, February 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lopes:2015:OCA

[720] Adrian A. Lopes. Organized crimes against nature: elephants in Southern Africa. Natural Resource Modeling, 28(1):86–107, February 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2015:IIb

[721] Anonymous. Issue information. Natural Resource Modeling, 28(2):i– ii, May 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yang:2015:MFM

[722] Zong-Chang Yang. Modeling and forecasting monthly average water levels based on the elliptic orbit model. Natural Resource Modeling, 28 (2):109–132, May 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Vasilieva:2015:HNU

[723] Olga Vasilieva. From harvesting to nonharvesting utility: an optimal control approach to species conservation. Natural Resource Modeling, 28 (2):133–151, May 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Laura-Guarachi:2015:MWF

[724] Leonardo R. Laura-Guarachi and Onésimo Hernández-Lerma. The mitra-wan forestry model: a discrete-time optimal control problem. Natural Resource Modeling, 28(2):152–168, May 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

DaSCosta:2015:CAE

[725] Michel Iskin Da S.Costa and Lucas Dos Anjos. Component Allee effects and stage-specific predation: a brief insight into conservation and biological control issues. Natural Resource Modeling, 28(2):169–183, May 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

VanKirk:2015:AUM

[726] Kray F. Van Kirk, Terrance J. Quinn II, Jeremy S. Collie, and Z. T. A’mar. Assessing uncertainty in a multispecies age-structured assessment framework: the effects of data limitations and model assumptions. Natural Resource Modeling, 28(2):184–205, May 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2015:IIc

[727] Anonymous. Issue information. Natural Resource Modeling, 28(3):i–ii, August 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jaiyeola:2015:DSU

[728] Adesoji T. Jaiyeola and Joseph K. Bwapwa. Dynamics of sedimentation and use of genetic algorithms for estimating sediment yields in a river: a critical review. Natural Resource Modeling, 28(3):207–218, August 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Poudel:2015:SOM

[729] Diwakar Poudel and Leif K. Sandal. Stochastic optimization for multispecies fisheries in the Barents Sea. Natural Resource Modeling, 28(3): 219–243, August 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Moeller:2015:EOM

[730] Holly V. Moeller and Michael G. Neubert. Economically optimal marine reserves without spatial heterogeneity in a simple two-patch model. Natural Resource Modeling, 28(3):244–255, August 2015. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lloyd:2015:MMC

[731] J. M. Lloyd and G. G. L. Meyer. Methods of modeling consumption saturation with uncertainty in optimal stockpile problems. Natural Resource Modeling, 28(3):256–288, August 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tausch:2015:SBA

[732] Robin J. Tausch. A structurally based analytic model of growth and biomass dynamics in single species stands of conifers. Natural Resource Modeling, 28(3):289–320, August 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kvamsdal:2015:EKF

[733] Sturla F. Kvamsdal and Leif K. Sandal. The ensemble Kalman filter for multidimensional bioeconomic models. Natural Resource Modeling, 28 (3):321–347, August 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tsehaye:2015:PMI

[734] Iyob Tsehaye, Michael L. Jones, Brian J. Irwin, David G. Fielder, James E. Breck, and David R. Luukkonen. A predictive model to inform adaptive management of double-crested cormorants and fisheries in Michigan. Natural Resource Modeling, 28(3):348–376, August 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2015:IId

[735] Anonymous. Issue information. Natural Resource Modeling, 28(4):i–ii, November 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Henson:2015:ISI

[736] Shandelle M. Henson, J. M. Cushing, and James L. Hayward. Introduction to special issue on eco-evolutionary dynamics. Natural Resource Modeling, 28(4):377–379, November 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ackleh:2015:DEC

[737] Azmy S. Ackleh, J. M. Cushing, and Paul L. Salceanu. On the dynamics of evolutionary competition models. Natural Resource Modeling, 28(4): 380–397, November 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kang:2015:CED

[738] Yun Kang and Jennifer Harrison Fewell. Co-evolutionary dynamics of a social parasite–host interaction model: obligate versus facultative social parasitism. Natural Resource Modeling, 28(4):398–455, November 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Moberg:2015:BMR

[739] Emily A. Moberg, Esther Shyu, Guillermo E. Herrera, Suzanne Lenhart, Yuan Lou, and Michael G. Neubert. On the bioeconomics of marine reserves when dispersal evolves. Natural Resource Modeling, 28(4):456– 474, November 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Schreiber:2015:EIA

[740] Sebastian J. Schreiber and Swati Patel. Evolutionarily induced alternative states and coexistence in systems with apparent competition. Natural Resource Modeling, 28(4):475–496, November 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cushing:2015:EGT

[741] J. M. Cushing, Shandelle M. Henson, and James L. Hayward. An evolutionary game-theoretic model of cannibalism. Natural Resource Modeling, 28(4):497–521, November 2015. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2016:IIa

[742] Anonymous. Issue information. Natural Resource Modeling, 29(1):1–2, February 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Vardas:2016:MIP

[743] Giannis Vardas and Anastasios Xepapadeas. Managing interacting populations under time scale separation. Natural Resource Modeling, 29(1): 5–35, February 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kelly:2016:OFH

[744] Michael R. Kelly, Jr., Yulong Xing, and Suzanne Lenhart. Optimal fish harvesting for a population modeled by a nonlinear parabolic partial differential equation. Natural Resource Modeling, 29(1):36–70, February 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Levy:2016:MMF

[745] Benjamin Levy, Charles Collins, Suzanne Lenhart, Marguerite Madden, Joseph Corn, RenÉ A. Salinas, and William Stiver. A metapopulation model for feral hogs in great smoky mountains national park. Natural Resource Modeling, 29(1):71–97, February 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Esteban:2016:RGD

[746] Encarna Esteban and Ariel Dinar. The role of groundwater-dependent ecosystems in groundwater management. Natural Resource Modeling, 29 (1):98–129, February 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Livadiotis:2016:KFU

[747] G. Livadiotis, L. Assas, B. Dennis, S. Elaydi, and E. Kwessi. Kappa function as a unifying framework for discrete population modeling. Natural Resource Modeling, 29(1):130–144, February 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Collins:2016:UMS

[748] Obiora C. Collins and Kevin J. Duffy. Understanding multiple species ecosystem dynamics using a consumer resource model. Natural Resource Modeling, 29(1):145–158, February 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Robinson:2016:FAC

[749] Orin J. Robinson, Conor P. Mcgowan, Patrick K. Devers, Rodney W. Brook, Min Huang, Malcom Jones, Daniel G. Mcauley, and Guthrie Zimmerman. A full annual cycle modeling framework for American black ducks. Natural Resource Modeling, 29(1):159–174, February 2016. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2016:IIb

[750] Anonymous. Issue information. Natural Resource Modeling, 29(2):175– 178, May 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Turner:2016:DES

[751] Benjamin L. Turner, Melissa Wuellner, Timothy Nichols, Roger Gates, Luis O. Tedeschi, and Barry H. Dunn. Development and evaluation of a system dynamics model for investigating agriculturally driven land transformation in the North Central United States. Natural Resource Modeling, 29(2):179–228, May 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Uecker:2016:OHS

[752] Hannes Uecker. Optimal harvesting and spatial patterns in a semiarid vegetation system. Natural Resource Modeling, 29(2):229–258, May 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kumari:2016:DWD

[753] Nitu Kumari and Sandeep Sharma. Does water disinfectant play a supportive role in the spread of infectious disease? a mathematical study. Natural Resource Modeling, 29(2):259–288, May 2016. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Seung:2016:REI

[754] Chang Seung and James Ianelli. Regional economic impacts of climate change: a computable general equilibrium analysis for an Alaska fishery. Natural Resource Modeling, 29(2):289–333, May 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2016:IIc

[755] Anonymous. Issue information. Natural Resource Modeling, 29(3):335– 338, August 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Meyer:2016:MRR

[756] Katherine Meyer. A mathematical review of resilience in ecology. Natural Resource Modeling, 29(3):339–352, August 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Alvarez:2016:MCP

[757] María Evangelina Alvarez, Silvia Di Marco, Katrin Erdlenbruch, and Mabel Tidball. A mixed control problem of the management of natural resources. Natural Resource Modeling, 29(3):353–373, August 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

VanDerLaan:2016:CDM

[758] Gerard Van Der Laan and Nigel Moes. Collective decision making in an international river pollution model. Natural Resource Modeling, 29 (3):374–399, August 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ravn-Jonsen:2016:IEE

[759] Lars Ravn-Jonsen, Ken H. Andersen, and Niels Vestergaard. An indicator for ecosystem externalities in fishing. Natural Resource Modeling, 29 (3):400–425, August 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Collins:2016:SAO

[760] O. C. Collins and K. S. Govinder. Stability analysis and optimal vaccination of a waterborne disease model with multiple water sources. Natural Resource Modeling, 29(3):426–447, August 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dawson:2016:MSE

[761] Heather A. Dawson, Michael L. Jones, Brian J. Irwin, Nicholas S. Johnson, Michael C. Wagner, and Melissa D. Szymanski. Management strategy evaluation of pheromone-baited trapping techniques to improve management of invasive sea lamprey. Natural Resource Modeling, 29(3):448– 469, August 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Verma:2016:MER

[762] Maitri Verma. Modeling the effect of rarity value on the exploitation of a wildlife species subjected to the Allee effect. Natural Resource Modeling, 29(3):470–494, August 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2016:IId

[763] Anonymous. Issue information. Natural Resource Modeling, 29(4):495– 498, November 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Collins:2016:OCI

[764] O. C. Collins and K. J. Duffy. Optimal control intervention strategies using an N-patch waterborne disease model. Natural Resource Modeling, 29(4):499–519, November 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Liao:2016:MBT

[765] Chao-Ning Liao and Hsiang-Wei Tsai. Modeling of a block tariff system for groundwater regulation. Natural Resource Modeling, 29(4):520–537, November 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Davison:2016:UAS

[766] Raziel J. Davison and William H. Satterthwaite. Use of age- and stagestructured matrix models to predict life history schedules for semelparous populations. Natural Resource Modeling, 29(4):538–558, November 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lata:2016:MEW

[767] Kusum Lata, B. Dubey, and A. K. Misra. Modeling the effects of wood and non-wood based industries on forestry resources. Natural Resource Modeling, 29(4):559–580, November 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lund:2016:WIU

[768] Arne-Christian Lund and Stein Ivar Steinshamn. Why increased uncertainty may lead to more risky behavior. Natural Resource Modeling, 29(4):581–599, November 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rossi:2016:SLP

[769] Marcelo Margon Rossi and Sonia Ternes. A simple linearization procedure for obtaining R_{0} from invasive epidemic models. Natural Resource Modeling, 29(4):600–609, November 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Skonhoft:2016:MSY

[770] Anders Skonhoft and Peichen Gong. Maximum sustainable yield harvesting in an age-structured fishery population model. Natural Resource Modeling, 29(4):610–632, November 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rodriguez-Rodriguez:2016:CED

[771] Marisabel Rodriguez-Rodriguez and Yun Kang. Colony and evolutionary dynamics of a two-stage model with brood cannibalism and division of labor in social insects. Natural Resource Modeling, 29(4):633–662, November 2016. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2017:IIa

[772] Anonymous. Issue information. Natural Resource Modeling, 30(1):1–4, February 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Doyen:2017:ISI

[773] Luc Doyen. Introduction to special issue on modeling and sustainability of biodiversity and ecosystem services. Natural Resource Modeling, 30 (1):5–9, February 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Drechsler:2017:CUS

[774] Martin Drechsler and Frank Wätzold. Costs of uncoordinated site selection with multiple ecosystem services. Natural Resource Modeling, 30 (1):10–29, February 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Costello:2017:PCT

[775] Christopher Costello and Daniel Kaffine. Private conservation in turfmanaged fisheries. Natural Resource Modeling, 30(1):30–51, February 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cairns:2017:FFF

[776] Robert D. Cairns. Faustmann’s formulas for forests. Natural Resource Modeling, 30(1):52–73, February 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ryan:2017:IPA

[777] Daniel Ryan, Carl Toews, James N. Sanchirico, and Paul R. Armsworth. Implications of policy adjustment costs for fisheries management. Natural Resource Modeling, 30(1):74–90, February 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Grafton:2017:RRN

[778] R. Quentin Grafton and L. Richard Little. Risks, resilience, and natural resource management: lessons from selected findings. Natural Resource Modeling, 30(1):91–111, February 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2017:IIb

[779] Anonymous. Issue information. Natural Resource Modeling, 30(2): e12115:1–e12115:??, May 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jordan:2017:MSA

[780] S. Jordan, C. Nischwitz, R. Ramirez, and L. F. Gordillo. Managing the spread of alfalfa stem nematodes (Ditylenchus dipsaci): The relationship between crop rotation periods and pest reemergence. Natural Resource Modeling, 30(2):e12083:1–e12083:??, May 2017. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yoshimoto:2017:IPA

[781] Atsushi Yoshimoto, Patrick Asante, Masashi Konoshima, and Peter SurovÝ. Integer programming approach to control invasive species spread based on cellular automaton model. Natural Resource Modeling, 30(2), May 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Getz:2017:WAP

[782] Wayne M. Getz, Oliver C. Muellerklein, Richard M. Salter, Colin J. Carlson, Andrew J. Lyons, and Dana P. Seidel. A web app for population viability and harvesting analyses. Natural Resource Modeling, 30 (2):e12120:1–e12120:??, May 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xuan:2017:MRC

[783] Bui Bich Xuan and Claire W. Armstrong. Marine reserve creation and interactions between fisheries and capture-based aquaculture: a bioeconomic model analysis. Natural Resource Modeling, 30(2):e12122:1– e12122:??, May 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hunt:2017:PMP

[784] E. Hunt and A. Tongen. A periodic matrix population model for monarch butterflies. Natural Resource Modeling, 30(2):e12123:1–e12123:??, May 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Manuel:2017:OTC

[785] Devasia Manuel and Richard B. Sowers. Optimal transport to cold chain in perishable hand-picked agriculture. Natural Resource Modeling, 30(2): e12124:1–e12124:??, May 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rautiainen:2017:CTL

[786] Aapo Rautiainen, Jussi Lintunen, and Jussi Uusivuori. Carbon taxation of the land use sector — the economics of soil carbon. Natural Resource Modeling, 30(2):e12126:1–e12126:??, May 2017. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sweeney:2017:HDF

[787] Jonathan R. Sweeney, Richard E. Howitt, Hing Ling Chan, Minling Pan, and PingSun Leung. How do fishery policies affect Hawai’i’s longline fishing industry? Calibrating a positive mathematical programming model. Natural Resource Modeling, 30(2):e12127:1–e12127:??, May 2017. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kanik:2017:PTF

[788] Zafer Kanik and Serkan Kucuksenel. The promise of transferable fishing concessions on EU fisheries. Natural Resource Modeling, 30(2):e12128:1– e12128:??, May 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2017:IIc

[789] Anonymous. Issue information. Natural Resource Modeling, 30(3): e12116:1–e12116:??, August 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Henson:2017:NOP

[790] Shandelle M. Henson. New opportunities for publishing in Natural Resource Modeling. Natural Resource Modeling, 30(3):e12136:1–e12136:??, August 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anagnostou:2017:EME

[791] Eleni Anagnostou, Areti Gianni, and Ierotheos Zacharias. Ecological modeling and eutrophication — a review. Natural Resource Modeling, 30 (3):e12130:1–e12130:??, August 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lade:2017:GME

[792] Steven J. Lade and Susa Niiranen. Generalized modeling of empirical social-ecological systems. Natural Resource Modeling, 30(3):e12129:1– e12129:??, August 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Haider:2017:IAE

[793] Humza S. Haider, Sarah C. Oldfield, Tiffany Tu, Rosa K. Moreno, Jay E. Diffendorfer, Eric A. Eager, and Richard A. Erickson. Incorporating Allee effects into the potential biological removal level. Natural Resource Modeling, 30(3):e12133:1–e12133:??, August 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Durand:2017:ATV

[794] Marie-Hélène Durand, Anna Désilles, Patrick Saint-Pierre, Valérie Angeon, and Harry Ozier-Lafontaine. Agroecological transition: a viability model to assess soil restoration. Natural Resource Modeling, 30(3): e12134:1–e12134:??, August 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Levy:2017:EWH

[795] Benjamin Levy, Charles Collins, Suzanne Lenhart, and William Stiver. Evaluating wild hog preferences to guide control strategies in the Great Smoky Mountains National Park. Natural Resource Modeling, 30(3): e12132:1–e12132:??, August 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bodine:2017:CCB

[796] Erin N. Bodine and Alex Capaldi. Can culling Barred Owls save a declining Northern Spotted Owl population? Natural Resource Modeling, 30 (3):e12131:1–e12131:??, August 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Olmos-Liceaga:2017:DSS

[797] Daniel Olmos-Liceaga, Geiser Villavicencio-Pulido, and Manuel Adrian Acuña-Zegarra. Diffusion as a strategy for survival in an invasion predator–prey model. Natural Resource Modeling, 30(3):e12135:1– e12135:??, August 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2017:IId

[798] Anonymous. Issue information. Natural Resource Modeling, 30(4): e12117:1–e12117:??, November 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Vandermeer:2017:ERF

[799] John Vandermeer. Ecological resilience in the face of catastrophic damage: The case of Hurricane Maria in Puerto Rico. Natural Resource Modeling, 30(4):e12149:1–e12149:??, November 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Elofsson:2017:OMT

[800] Katarina Elofsson, Justice Tei Mensah, and Petter Kjellander. Optimal management of two ecologically interacting deer species — reality matters, beliefs don’t. Natural Resource Modeling, 30(4):e12137:1–e12137:??, November 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Alrashidi:2017:CCM

[801] Mohammed E. Alrashidi, John W. Hearne, Lynne McArthur, and Claudio Zorzan. Cooperative considerations for a mobile resource that transcends property boundaries. Natural Resource Modeling, 30(4):e12138:1– e12138:??, November 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jyotsna:2017:MMS

[802] Kumari Jyotsna and Abhinav Tandon. A mathematical model studying the survival of forest resource–dependent wildlife population in the presence of population pressure–induced mining activities. Natural Resource Modeling, 30(4):e12139:1–e12139:??, November 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Smith:2017:EDC

[803] Gary Smith. Estimation of the demand for construction aggregate. Natural Resource Modeling, 30(4):e12144:1–e12144:??, November 2017. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ranjan:2017:PES

[804] Ram Ranjan. Preserving ecologically sensitive hotspots under endogenous regulation and valuation constraints. Natural Resource Modeling, 30(4):e12145:1–e12145:??, November 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lin:2017:AVC

[805] Li Lin and Pgrni Pussella. Assessment of vulnerability for coastal erosion with GIS and AHP techniques case study: Southern coastline of Sri Lanka. Natural Resource Modeling, 30(4):e12146:1–e12146:??, November 2017. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2018:IIa

[806] Anonymous. Issue information. Natural Resource Modeling, 31(1): e12140:1–e12140:??, February 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Henson:2018:DCR

[807] Shandelle M. Henson and Suzanne L. Robertson. Dedication to Catherine A. Roberts. Natural Resource Modeling, 31(1):e12161:1–e12161:??, February 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Case:2018:MEF

[808] James Case. Modeling the energy future. Natural Resource Modeling, 31(1):e12147:1–e12147:??, February 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Blackwood:2018:MAS

[809] Julie C. Blackwood, Colin Okasaki, André Archer, Eliza W. Matt, Elizabeth Sherman, and Kathryn Montovan. Modeling alternative stable states in Caribbean coral reefs. Natural Resource Modeling, 31(1): e12157:1–e12157:??, February 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Robertson:2018:MMD

[810] Suzanne L. Robertson, Shandelle M. Henson, Timothy Robertson, and J. M. Cushing. A matter of maturity: To delay or not to delay? Continuous-time compartmental models of structured populations in the literature 2000–2016. Natural Resource Modeling, 31(1):e12160:1– e12160:??, February 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bieri:2018:GCH

[811] J. A. Bieri, C. Sample, W. E. Thogmartin, J. E. Diffendorfer, J. E. Earl, R. A. Erickson, P. Federico, D. T. T. Flockhart, S. Nicol, D. Semmens, T. Skraber, R. Wiederholt, and B. J. Mattsson. A guide to calculating habitat-quality metrics to inform conservation of highly mobile species. Natural Resource Modeling, 31(1):e12156:1–e12156:??, February 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Heines:2018:AET

[812] Betsy Heines, Suzanne Lenhart, and Charles Sims. Assessing the economic trade-offs between prevention and suppression of forest fires. Natural Resource Modeling, 31(1):e12159:1–e12159:??, February 2018. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Srivastava:2018:AGH

[813] Nitin Srivastava, Peter Maneykowski, and Richard B. Sowers. Algorithmic geolocation of harvest in hand-picked agriculture. Natural Resource Modeling, 31(1):e12158:1–e12158:??, February 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2018:IIb

[814] Anonymous. Issue information. Natural Resource Modeling, 31(2): e12141:1–e12141:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Osmani:2018:FMS

[815] Dritan Osmani and Richard S. J. Tol. On the farsightedly and myopically stable international environmental agreements. Natural Resource Modeling, 31(2):e12154:1–e12154:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

ZiaeeVafaeyan:2018:LUC

[816] Hilda ZiaeeVafaeyan, Mohammad Hossein Moattar, and Yahya Forghani. Land use change model based on bee colony optimization, Markov chain and a neighborhood decay cellular automata. Natural Resource Modeling, 31(2):e12151:1–e12151:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

vanVuuren:2018:ABS

[817] B. J. van Vuuren, L. Potgieter, and J. H. van Vuuren. An agent-based simulation model of Eldana saccharina Walker. Natural Resource Modeling, 31(2):e12153:1–e12153:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Raoufi:2018:SGY

[818] Roxana Seyed Raoufi, Saeid Soufizadeh, Bahman Amiri Larijani, Majid AghaAlikhani, and Jafar Kambouzia. Simulation of growth and yield of various irrigated rice (Oryza sativa L.) genotypes by AquaCrop under different seedling ages. Natural Resource Modeling, 31(2):e12162:1– e12162:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

He:2018:OAF

[819] Guo-Feng He, Zhi-Guang Li, Ying Yuan, Xue-Hua Li, Long-Hu Hu, and Yun Zhang. Optimization analysis of the factors affecting the soil arching effect between landslide stabilizing piles. Natural Resource Modeling, 31 (2):e12148:1–e12148:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jordan:2018:YRI

[820] Scott Jordan. Yield to the resistance: The impact of nematode resistant varieties on alfalfa yield. Natural Resource Modeling, 31(2):e12150:1– e12150:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Guo:2018:LCC

[821] Shili Guo, Chunjie Li, Shaoquan Liu, and Kui Zhou. Land carrying capacity in rural settlements of three gorges reservoir based on the system dynamic model. Natural Resource Modeling, 31(2):e12152:1–e12152:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kahui:2018:ACH

[822] V. Kahui, B. Moyle, and A. M. Brunell. Alligator conservation and hunting efficiency. Natural Resource Modeling, 31(2):e12155:1–e12155:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jyotsna:2018:NMM

[823] Kumari Jyotsna and Abhinav Tandon. A nonlinear mathematical model investigating the sustainability of an urban system in the presence of haphazard urban development and excessive pollution. Natural Resource Modeling, 31(2):e12163:1–e12163:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Holzer:2018:CTH

[824] Jorge Holzer and Qian Qu. Confidence of the trembling hand: Bayesian learning with data-limited stocks. Natural Resource Modeling, 31(2): e12164:1–e12164:??, May 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2018:IIc

[825] Anonymous. Issue information. Natural Resource Modeling, 31(3): e12142:1–e12142:??, August 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Pastor:2018:ISI

[826] Elsa Pastor. Introduction to the special issue on “vulnerability and resilience of socio-ecological systems”. Natural Resource Modeling, 31(3): e12185:1–e12185:??, August 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bates:2018:BSE

[827] Samuel Bates, Valérie Angeon, Anna Désilles, and Patrick Saint-Pierre. Bounded set of economic uncertainty and robust viability: an illustration from farming systems. Natural Resource Modeling, 31(3):e12186:1– e12186:??, August 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Zeeman:2018:RSV

[828] Mary Lou Zeeman, Katherine Meyer, Erika Bussmann, Alanna Hoyer-Leitzel, Sarah Iams, Ian J. Klasky, Victoria Lee, and Stephen Ligtenberg. Resilience of socially valued properties of natural systems to repeated disturbance: a framework to support value-laden management decisions. Natural Resource Modeling, 31(3):e12170:1–e12170:??, August 2018. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xu:2018:WDF

[829] Qing Xu, Sylvie Huet, Christophe Poix, Isabelle Boisdon, and Guillaume Deffuant. Why do farmers not convert to organic farming? Modeling conversion to organic farming as a major change. Natural Resource Modeling, 31(3):e12171:1–e12171:??, August 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Poudel:2018:IAA

[830] Biswo N. Poudel and Krishna P. Paudel. An integrated approach to analyzing risk in bioeconomic models. Natural Resource Modeling, 31(3): e12172:1–e12172:??, August 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Blanco:2018:MMF

[831] Juan A. Blanco, Deborah S. Page-Dumroese, Martin F. Jurgensen, Michael P. Curran, Joanne M. Tirocke, and Joanna Walitalo. Modelling the management of forest ecosystems: Importance of wood decomposition. Natural Resource Modeling, 31(3):e12173:1–e12173:??, August 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gorfine:2018:MTI

[832] Harry Gorfine, Jim Thomson, Daniel Spring, and Michael Cleland. Modelling trends including effects of natural disturbance in an abalone dive fishery in Australia. Natural Resource Modeling, 31(3):e12175:1– e12175:??, August 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Nichols:2018:AHB

[833] Rachel Nichols, Satoshi Yamazaki, and Sarah Jennings. Allocation of harvest between user groups in a fishery with habitat effect. Natural Resource Modeling, 31(3):e12179:1–e12179:??, August 2018. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Stein:2018:UMS

[834] Alfred Stein and Valentyn A. Tolpekin. Uncertainty in the modeling of spatial big data on a pattern of bushfires holes. Natural Resource Modeling, 31(3):e12180:1–e12180:??, August 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2018:IId

[835] Anonymous. Issue information. Natural Resource Modeling, 31(4): e12143:1–e12143:??, November 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Leander:2018:DSL

[836] Rachel Leander, Wandi Ding, and René A. Salinas. Dedication to Suzanne Lenhart. Natural Resource Modeling, 31(4):e12198:1–e12198:??, November 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Schaefer:2018:MPM

[837] Elsa Schaefer, Kevin A. Caillouët, and Suzanne L. Robertson. Methods for prophylactic management of West Nile virus using a stage-structured avian host-vector model with vaccination, larvicide, and adulticide. Natural Resource Modeling, 31(4):e12165:1–e12165:??, November 2018. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kwessi:2018:NED

[838] Eddy Kwessi, Saber Elaydi, Brian Dennis, and George Livadiotis. Nearly exact discretization of single species population models. Natural Resource Modeling, 31(4):e12167:1–e12167:??, November 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cantrell:2018:EHM

[839] R. S. Cantrell and C. Cosner. Effects of harvesting mediated by dispersal traits. Natural Resource Modeling, 31(4):e12168:1–e12168:??, November 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Baker:2018:OSE

[840] Christopher M. Baker, Fasma Diele, Carmela Marangi, Angela Martiradonna, and Stefania Ragni. Optimal spatiotemporal effort allocation for invasive species removal incorporating a removal handling time and budget. Natural Resource Modeling, 31(4):e12190:1–e12190:??, November 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sims:2018:CFN

[841] Charles Sims, Richard D. Horan, and Benjamin Meadows. Come on feel the noise: Ecological foundations in stochastic bioeconomic models. Natural Resource Modeling, 31(4):e12191:1–e12191:??, November 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yakubu:2018:DTA

[842] Abdul-Aziz Yakubu and Najat Ziyadi. A discrete-time anthrax model in human and herbivore populations. Natural Resource Modeling, 31(4): e12192:1–e12192:??, November 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jang:2018:CHP

[843] Sophia R.-J. Jang, Wenjing Zhang, and Victoria Larriva. Cooperative hunting in a predator–prey system with Allee effects in the prey. Natural Resource Modeling, 31(4):e12194:1–e12194:??, November 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Miller:2018:IAD

[844] Adam D. Miller, Pen-Yuan Hsing, Stephen H. Roxburgh, Charles R. Fisher, and Katriona Shea. Impacts of altered disturbance regimes on community structure and biodiversity mediated by fecundity–tolerance interactions. Natural Resource Modeling, 31(4):e12199:1–e12199:??, November 2018. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2019:IIa

[845] Anonymous. Issue information. Natural Resource Modeling, 32(1): e12181:1–e12181:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ackleh:2019:SAR

[846] A. S. Ackleh, H. Caswell, R. A. Chiquet, T. Tang, and A. Veprauskas. Sensitivity analysis of the recovery time for a population under the impact of an environmental disturbance. Natural Resource Modeling, 32 (1):e12166:1–e12166:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ropero:2019:BNE

[847] Rosa F. Ropero, Rafael Rumí, and Pedro A. Aguilera. Bayesian networks for evaluating climate change influence in olive crops in Andalusia, Spain. Natural Resource Modeling, 32(1):e12169:1–e12169:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Devi:2019:DCD

[848] Sapna Devi and Nivedita Gupta. Dynamics of carbon dioxide gas (CO_{2}): Effects of varying capability of plants to absorb CO_{2}. Natural Resource Modeling, 32(1):e12174:1–e12174:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Li:2019:ELT

[849] Wenhua Li and Tsuyoshi Adachi. Evaluation of long-term silver supply shortage for c-Si PV under different technological scenarios. Natural Resource Modeling, 32(1):e12176:1–e12176:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Batabyal:2019:PAC

[850] Amitrajeet A. Batabyal and Hamid Beladi. Probabilistic approaches to cleaning the Ganges in Varanasi to attract tourists. Natural Resource Modeling, 32(1):e12177:1–e12177:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kucuker:2019:MPC

[851] Derya Mumcu Kucuker and Emin Zeki Baskent. Modeling the productivity of commercial Lactarius mushrooms: a case study in the Kızılcasu planning unit, Turkey. Natural Resource Modeling, 32(1):e12178:1– e12178:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Usoltsev:2019:CAM

[852] Vladimir Andreevich Usoltsev, Seyed Omid Reza Shobairi, and Viktor Petrovich Chasovskikh. Comparing of allometric models of singletree biomass intended for airborne laser sensing and terrestrial taxation of carbon pool in the forests of Eurasia. Natural Resource Modeling, 32 (1):e12187:1–e12187:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ranjan:2019:SPW

[853] Ram Ranjan. Shooting at the poachers while the rhinos drown: Managing short- and long-term threats to endangered wildlife in conservation reserves. Natural Resource Modeling, 32(1):e12188:1–e12188:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Oyebode:2019:NNM

[854] Oluwaseun Oyebode and Derek Stretch. Neural network modeling of hydrological systems: a review of implementation techniques. Natural Resource Modeling, 32(1):e12189:1–e12189:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Liu:2019:EIF

[855] Yiming Liu, Stuart W. Bunting, Shiming Luo, Kunzheng Cai, and Qiangqiang Yang. Evaluating impacts of fish stock enhancement and biodiversity conservation actions on the livelihoods of small-scale fishers on the Beijiang River, China. Natural Resource Modeling, 32(1):e12195:1– e12195:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cruz-Rivera:2019:CSD

[856] Erica Cruz-Rivera, Héctor Ramírez C., and Olga Vasilieva. Catchto-stock dependence: The case of small pelagic fishery with bounded harvesting effort. Natural Resource Modeling, 32(1):e12193:1–e12193:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Khan:2019:CML

[857] Syed S. Khan, Shawn Arita, Richard Howitt, and PingSun Leung. A calibrated model of local food system of Hawaii: What are the economic implications of the state’s food goals and policies? Natural Resource Modeling, 32(1):e12196:1–e12196:??, February 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2019:IIb

[858] Anonymous. Issue information. Natural Resource Modeling, 32(2): e12182:1–e12182:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Soedomo:2019:SMB

[859] Sudarsono Soedomo. A simple model to bridge Hardin’s tragedy and Ostrom’s comity. Natural Resource Modeling, 32(2):e12205:1–e12205:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mesbahzadeh:2019:FFA

[860] Tayyebeh Mesbahzadeh, Farshad Soleimani Sardoo, and Shapour Kouhestani. Flood frequency analysis for the Iranian interior deserts using the method of L-moments: a case study in the Loot River Basin. Natural Resource Modeling, 32(2):e12208:1–e12208:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Pereira:2019:ALE

[861] Omar Cléo Neves Pereira, Tiago Peres da Silva Suguiura, Anaih Pastana Pereira, Altair Bertonha, and Isolde Previdelli. Analysis of lettuce evapotranspiration across soil water. Natural Resource Modeling, 32(2): e12197:1–e12197:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bergland:2019:SUE

[862] Harald Bergland, Pål Andreas Pedersen, and John Wyller. Stable and unstable equilibrium states in a fishery–aquaculture model. Natural Resource Modeling, 32(2):e12200:1–e12200:??, May 2019. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Peng:2019:FEE

[863] Li Peng, Wei Deng, Hao Zhang, Jian Sun, and Junnan Xiong. Focus on economy or ecology? a three-dimensional trade-off based on ecological carrying capacity in southwest China. Natural Resource Modeling, 32 (2):e12201:1–e12201:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Fattahpour:2019:DCR

[864] Haniyeh Fattahpour, Hamid R. Z. Zangeneh, and Hao Wang. Dynamics of coral reef models in the presence of parrotfish. Natural Resource Modeling, 32(2):e12202:1–e12202:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Chen:2019:DBF

[865] Zhijing Chen and Wenjun Liu. Dynamical behavior of fractional-order energy-saving and emission-reduction system and its discretization. Natural Resource Modeling, 32(2):e12203:1–e12203:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Brunetti:2019:RBB

[866] Ilaria Brunetti, Mabel Tidball, and Denis Couvet. Relationship between biodiversity and agricultural production. Natural Resource Modeling, 32 (2):e12204:1–e12204:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Pichika:2019:OHR

[867] Srinivasu D. N. Pichika and Simon D. Zawka. Optimal harvesting of a renewable resource in a polluted environment: an allocation problem of the sole owner. Natural Resource Modeling, 32(2):e12206:1–e12206:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hanley:2019:AEE

[868] Brenda Hanley and Brian Dennis. Analytical expressions for the eigenvalues, demographic quantities, and extinction criteria arising from a three-stage wildlife population matrix. Natural Resource Modeling, 32 (2):e12207:1–e12207:??, May 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2019:IIc

[869] Anonymous. Issue information. Natural Resource Modeling, 32(3): e12183:1–e12183:??, August 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Paudel:2019:ISI

[870] Krishna P. Paudel. Introduction to the special issue on “Economic modeling of natural resources for sustainable development”. Natural Resource Modeling, 32(3):e12238:1–e12238:??, August 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Peterson:2019:IPS

[871] Jeffrey M. Peterson. Innovation as a policy strategy for natural resource protection. Natural Resource Modeling, 32(3):e12231:1–e12231:??, August 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Susaeta:2019:OHS

[872] Andres Susaeta and Peichen Gong. Optimal harvest strategy for evenaged stands with price uncertainty and risk of natural disturbances. Natural Resource Modeling, 32(3):e12211:1–e12211:??, August 2019. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dahmouni:2019:FTC

[873] Ilyass Dahmouni, Baris Vardar, and Georges Zaccour. A fair and timeconsistent sharing of the joint exploitation payoff of a fishery. Natural Resource Modeling, 32(3):e12216:1–e12216:??, August 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Poudel:2019:TEG

[874] Biswo N. Poudel and Krishna P. Paudel. Transboundary extraction of groundwater in the presence of hydraulic fracturing. Natural Resource Modeling, 32(3):e12217:1–e12217:??, August 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Chen:2019:FHR

[875] Fengbo Chen, Qiuzhuo Ma, and Wenbin Huang. Farm households’ rice production behavior in China under a separability assumption: a metaheuristic optimization approach. Natural Resource Modeling, 32(3): e12221:1–e12221:??, August 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Zhang:2019:IER

[876] Yijing Zhang, Wen Luo, Xinyun Duan, and Lan Gao. The impact of environmental regulations on forest product trade in China. Natural Resource Modeling, 32(3):e12234:1–e12234:??, August 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2019:IId

[877] Anonymous. Issue information. Natural Resource Modeling, 32(4): e12184:1–e12184:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ben-Hasan:2019:PCM

[878] Abdul Ben-Hasan, Carl Walters, Villy Christensen, Brett van Poorten, Sarah Rajab, Mohsen Al-Husaini, and Hussein Al-Foudari. Predicting changes in mean length with an age-structured stock assessment model. Natural Resource Modeling, 32(4):e12218:1–e12218:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Peng:2019:MLU

[879] Li Peng, Xuxi Wang, and Tiantian Chen. Multifunctional land-use value mapping and space type classification: a case study of Puge County, China. Natural Resource Modeling, 32(4):e12212:1–e12212:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rahimzadeh:2019:EFE

[880] Omid Rahimzadeh, Abdolreza Bahremand, Nader Noura, and Micah Mukolwe. Evaluating flood extent mapping of two hydraulic models, 1D HEC-RAS and 2D LISFLOOD-FP in comparison with aerial imagery observations in Gorgan flood plain, Iran. Natural Resource Modeling, 32(4):e12214:1–e12214:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Rodriguez-Quinones:2019:LDB

[881] Leoncio Rodriguez-Quinones and Luis F. Gordillo. Low-density barriers for controlling plant crowd diseases: How far and fast can pathogens spread? Natural Resource Modeling, 34(1):e12222:1–e12222:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xing:2019:SMS

[882] Shiqi Xing and Amitrajeet A. Batabyal. A safe minimum standard, the elasticity of substitution, and the cleanup of the Ganges in Varanasi. Natural Resource Modeling, 34(1):e12223:1–e12223:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Amundsen:2019:GMW

[883] Eirik S. Amundsen and Frank Jensen. Groundwater management: Waiting for a drought*. Natural Resource Modeling, 34(1):e12209:1–e12209:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

vanderWerf:2019:LDI

[884] Edwin van der Werf, Yonky Indrajaya, Frits Mohren, and Ekko C. van Ierland. Logging damage and injured tree mortality in tropical forest management. Natural Resource Modeling, 34(1):e12210:1–e12210:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bergland:2019:PLD

[885] Harald Bergland, John Wyller, and Evgenii Burlakov. Pasture–livestock dynamics with density-dependent harvest and changing environment. Natural Resource Modeling, 34(1):e12213:1–e12213:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lin:2019:CII

[886] Ping Lin, Shanchao Jiang, Du Li, Zhiyong Zou, and Yongming Chen. Comprehending international important Ramsar wetland documents using latent semantic topic model in kernel space. Natural Resource Modeling, 34(1):e12215:1–e12215:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Soh:2019:STP

[887] Moonwon Soh and Seong-Hoon Cho. Spatial targeting of payments for ecosystem services to achieve conservation goals and promote social equity and economic impact. Natural Resource Modeling, 34(1):e12219:1– e12219:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tan:2019:OTD

[888] Yi Tan, Lijuan Ning, Sanyi Tang, and Robert A. Cheke. Optimal threshold density in a stochastic resource management model with pulse intervention. Natural Resource Modeling, 34(1):e12220:1–e12220:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cairns:2019:PIP

[889] Robert D. Cairns. A point-input, point-output forest as a microprototype for capital accounting. Natural Resource Modeling, 34(1):e12224:1– e12224:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Haq:2019:ANN

[890] Mohd Anul Haq, Abhijit Ghosh, Gazi Rahaman, and Prashant Baral. Artificial neural network-based modeling of snow properties using field data and hyperspectral imagery. Natural Resource Modeling, 34(1):e12229:1– e12229:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Djomegni:2019:CHC

[891] Patrick M. Tchepmo Djomegni, Emile F. Doungmo Goufo, Subrata K. Sahu, and Mohamed Mbehou. Coexistence and harvesting control policy in a food chain model with mutual defense of prey. Natural Resource Modeling, 34(1):e12230:1–e12230:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Blanchard:2019:BFR

[892] Fabian Blanchard, Christian Chaboud, and Olivier Thébaud. Back to the future: a retrospective assessment of model-based scenarios for the management of the shrimp fishery in French Guiana facing global change. Natural Resource Modeling, 34(1):e12232:1–e12232:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Loisel:2019:SPF

[893] Patrice Loisel. Stochastic perturbations and fisheries management. Natural Resource Modeling, 34(1):e12233:1–e12233:??, November 2019. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Grass:2019:OFC

[894] Dieter Grass, Hannes Uecker, and Thorsten Upmann. Optimal fishery with coastal catch. Natural Resource Modeling, 34(1):e12235:1– e12235:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2019:E

[895] Anonymous. Erratum. Natural Resource Modeling, 34(1):e12236:1– e12236:??, November 2019. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2020:IIa

[896] Anonymous. Issue information. Natural Resource Modeling, 33(1): e12225:1–e12225:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wrobel:2020:ASI

[897] Michał Wróbel, Kamil Mańk, and Anna Krysztofiak-Kaniewska. Applying the stocking index to the determination of the curve number parameter in the forest catchment area. Natural Resource Modeling, 33(1): e12241:1–e12241:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bazrafshan:2020:CBI

[898] Ommolbanin Bazrafshan, Hossein Zamani, and Marzieh Shekari. A copula-based index for drought analysis in arid and semi-arid regions of Iran. Natural Resource Modeling, 33(1):e12237:1–e12237:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Shen:2020:CRB

[899] Yongchang Shen, Yunyun Fu, and Malin Song. Coupling relationship between green production and green consumption: Case of the Yangtze River Delta area. Natural Resource Modeling, 33(1):e12239:1–e12239:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Meng:2020:MBW

[900] Yuan Meng, Xibin Dong, Weiguo Liu, and Wenshu Lin. Modeling biomass of white birch (Betula platyphylla) in the Lesser Khingan Range of China based on terrestrial 3D laser scanning system. Natural Resource Modeling, 33(1):e12240:1–e12240:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Meresa:2020:RFC

[901] Hadush Kidane Meresa. River flow characteristics and changes under the influence of varying climate conditions. Natural Resource Modeling, 33(1):e12242:1–e12242:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Menendez:2020:SLS

[902] Hector M. Menendez III, Melissa R. Wuellner, Benjamin L. Turner, Roger N. Gates, Barry H. Dunn, and Luis O. Tedeschi. A spatial landscape scale approach for estimating erosion, water quantity, and quality in response to South Dakota grassland conversion. Natural Resource Modeling, 33(1):e12243:1–e12243:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Baggio:2020:EAE

[903] Michele Baggio, Jean-Paul Chavas, Salvatore Di Falco, Andreas Hertig, and Francesco Pomati. The effect of anthropogenic and environmental factors in coupled human–natural systems: Evidence from Lake Zürich. Natural Resource Modeling, 33(1):e12245:1–e12245:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xie:2020:LFE

[904] Fangting Xie, Xiaolan Kang, Juan Du, Xiaojin Liu, and Shubin Zhu. Labor off-farm employment and household forest management investment in Jiangxi, China: a perspective from gender influence of rural labor. Natural Resource Modeling, 33(1):e12246:1–e12246:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

He:2020:EIG

[905] Yiming He and Heyuan Huang. Energy intensity in Guangdong of China (2006–2015): a spatial dynamic general equilibrium econometric model. Natural Resource Modeling, 33(1):e12247:1–e12247:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lin:2020:AIC

[906] Ping Lin, Qun Lu, Du Li, Yongming Chen, Zhiyong Zou, and Shanchao Jiang. Artificial intelligence classification of wetland vegetation morphology based on deep convolutional neural network. Natural Resource Modeling, 33(1):e12248:1–e12248:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ansink:2020:CPR

[907] Erik Ansink and Hans-Peter Weikard. Common pool resources with support. Natural Resource Modeling, 33(1):e12249:1–e12249:??, February 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2020:IIb

[908] Anonymous. Issue information. Natural Resource Modeling, 33(2): e12226:1–e12226:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Caplan:2020:MBI

[909] Arthur J. Caplan and Dong-Hun Go. Measuring the bioeconomic impacts of prolonged drought on a lake ecosystem: The case of the Great Salt Lake, Utah. Natural Resource Modeling, 33(2):e12251:1–e12251:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lyu:2020:CTC

[910] Jingjing Lyu, Pamela J. Schofield, Kristen M. Reaver, Matthew Beauregard, and Rana D. Parshad. A comparison of the Trojan Y Chromosome strategy to harvesting models for eradication of nonnative species. Natural Resource Modeling, 33(2):e12252:1–e12252:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Demir:2020:OSF

[911] Mahir Demir and Suzanne Lenhart. Optimal sustainable fishery management of the Black Sea anchovy with food chain modeling framework. Natural Resource Modeling, 33(2):e12253:1–e12253:??, May 2020. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mirakbari:2020:OPT

[912] Maryam Mirakbari, Tayyebeh Mesbahzadeh, Farshad Soleimani Sardoo, Mario M. Miglietta, Nir Y. Krakauer, and Nahid Alipour. Observed and projected trends of extreme precipitation and maximum temperature during 1992–2100 in Isfahan province, Iran using REMO model and copula theory. Natural Resource Modeling, 33(2):e12254:1–e12254:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Elzinga:2020:MCM

[913] David C. Elzinga, Shelby R. Stowe, and F. Leland Russell. Modeling control methods to manage the sylvatic plague in black-tailed prairie dog towns. Natural Resource Modeling, 33(2):e12255:1–e12255:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yuan:2020:PPM

[914] Huanjie Yuan, Yuanshi Wang, Hong Wu, and Shican Wu. Persistence of pollination-mutualisms under the effect of intermediary resource. Natural Resource Modeling, 33(2):e12259:1–e12259:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dawed:2020:CDT

[915] Mohammed Y. Dawed, Patrick M. Tchepmo Djomegni, and Harald E. Krogstad. Complex dynamics in a tritrophic food chain model with general functional response. Natural Resource Modeling, 33(2):e12260:1– e12260:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tandon:2020:DBT

[916] Abhinav Tandon. Dynamical behavior of tribal-forest system in the presence of development activities: Leslie–Gower-based nonlinear modeling study. Natural Resource Modeling, 33(2):e12261:1–e12261:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Haq:2020:ATL

[917] Mohd Anul Haq, Prashant Baral, Shivaprakash Yaragal, and Gazi Rahaman. Assessment of trends of land surface vegetation distribution, snow cover and temperature over entire Himachal Pradesh using MODIS datasets. Natural Resource Modeling, 33(2):e12262:1–e12262:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Devi:2020:CSE

[918] Sapna Devi and Nivedita Gupta. Comparative study of the effects of different growths of vegetation biomass on CO_{2} in crisp and fuzzy environments. Natural Resource Modeling, 33(2):e12263:1–e12263:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Progenio:2020:CDF

[919] Mayke F. Progênio and Claudio J. C. Blanco. Cumulative distribution function of daily rainfall in the Tocantins–Araguaia hydrographic region, Amazon, Brazil. Natural Resource Modeling, 33(2):e12264:1–e12264:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kovacs:2020:ILF

[920] Kent F. Kovacs and Alvaro Durand-Morat. The influence of lateral flows in an aquifer on the agricultural value of groundwater. Natural Resource Modeling, 33(2):e12266:1–e12266:??, May 2020. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dia:2020:OCI

[921] Ben M. Dia, Mamadou L. Diagne, and M. Samsidy Goudiaby. Optimal control of invasive species with economic benefits: Application to the Typha proliferation. Natural Resource Modeling, 33(2):e12268:1–e12268:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Khaleghi:2020:FHS

[922] Esmaeil Khaleghi, Amir Sadoddin, Ali Najafinejad, and Abdolreza Bahremand. Flood hydrograph simulation using the SWMM model: a semiarid zone watershed case study, Shiraz Khoshk River, Iran. Natural Resource Modeling, 33(2):e12269:1–e12269:??, May 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2020:IIc

[923] Anonymous. Issue information. Natural Resource Modeling, 33(3): e12227:1–e12227:??, August 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Breton:2020:ESI

[924] Michèle Breton and Georges Zaccour. Editorial: Special issue on resilience and viability of natural resources. Natural Resource Modeling, 33(3):e12280:1–e12280:??, August 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Long:2020:OHT

[925] Ngo Van Long, Mabel Tidball, and Georges Zaccour. Optimal harvesting and taxation when accounting for the marine environmental quality of the fishery. Natural Resource Modeling, 33(3):e12244:1–e12244:??, August 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Doyen:2020:SSM

[926] Luc Doyen and Pedro Gajardo. Sustainability standards, multicriteria maximin, and viability. Natural Resource Modeling, 33(3):e12250:1– e12250:??, August 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Filar:2020:RSB

[927] Jerzy A. Filar, Zhihao Qiao, and Sabrina Streipert. Risk sensitivity in Beverton–Holt fishery with multiplicative harvest. Natural Resource Modeling, 33(3):e12257:1–e12257:??, August 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wagener:2020:GMA

[928] Florian Wagener. Geometrical methods for analyzing the optimal management of tipping point dynamics. Natural Resource Modeling, 33(3): e12258:1–e12258:??, August 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ay:2020:DDS

[929] Jean-Sauveur Ay and Estelle Gozlan. Disease dispersion as a spatial interaction: The case of Flavescence Dorée. Natural Resource Modeling, 33(3):e12265:1–e12265:??, August 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Riquelme:2020:RUD

[930] Victor Riquelme, Terrance J. Quinn II, and Hector Ramirez C. The role of uncertainty in the design of sustainable and precautionary management strategies for fisheries. Natural Resource Modeling, 33(3):e12279:1– e12279:??, August 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2020:IId

[931] Anonymous. Issue information. Natural Resource Modeling, 33(4): e12228:1–e12228:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Koetke:2020:CRL

[932] Lisa J. Koetke, Adam Duarte, and Floyd W. Weckerly. Comparing the Ricker and θ-logistic models for estimating elk population growth. Natural Resource Modeling, 33(4):e12270:1–e12270:??, November 2020. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Peirce:2020:MMC

[933] J. P. Peirce, J. J. Pellett, and G. J. Sandland. A mathematical model for the control of swimmer’s itch. Natural Resource Modeling, 33(4): e12275:1–e12275:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ashrafi:2020:STD

[934] Tannaz Alizadeh Ashrafi, Arne Eide, and Øystein Hermansen. Spatial and temporal distributions in the Norwegian cod fishery. Natural Resource Modeling, 33(4):e12276:1–e12276:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Chen:2020:MSC

[935] Tiantian Chen, Li Peng, and Qiang Wang. From multifunctionality to sustainable cultivated land development? a three-dimensional tradeoff model tested in Panxi region of southwestern China. Natural Resource Modeling, 33(4):e12278:1–e12278:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tripathi:2020:PPM

[936] Jai Prakash Tripathi, Sarita Bugalia, Vandana Tiwari, and Yun Kang. A predator–prey model with Crowley–Martin functional response: a nonautonomous study. Natural Resource Modeling, 33(4):e12287:1–e12287:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Elzinga:2020:ABM

[937] David C. Elzinga, Erin Boggess, Jordan Collignon, Alanna Riederer, and Alex Capaldi. An agent-based model determining a successful reintroduction of the extinct passenger pigeon. Natural Resource Modeling, 33(4):e12292:1–e12292:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Salajegheh:2020:MIS

[938] Sosan Salajegheh, Hamid R. Jafari, and Sharareh Pourebrahim. Modeling the impact of social network measures on institutional adaptive capacity needed for sustainable governance of water resources. Natural Resource Modeling, 33(4):e12277:1–e12277:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Borland:2020:UMC

[939] Lauren K. Borland, Collin J. Mulcahy, Barbara A. Bennie, Douglas D. Baumann, Roger J. Haro, Molly Van Appledorn, Kathi Jo Jankowski, Aaron R. Cupp, and Richard A. Erickson. Using Markov chains to quantitatively assess movement patterns of invasive fishes impacted by a carbon dioxide barrier in outdoor ponds. Natural Resource Modeling, 33(4):e12281:1–e12281:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hamzaoui-Azaza:2020:HMG

[940] Fadoua Hamzaoui-Azaza, Mounira Zammouri, Meriem Ameur, Mohamedou Baba Sy, Moncef Gueddari, and Rachida Bouhlila. Hydrogeochemical modeling for groundwater management in arid and semiarid regions using MODFLOW and MT3DMS: a case study of the Jeffara of Medenine coastal aquifer, South-Eastern Tunisia. Natural Resource Modeling, 33(4):e12282:1–e12282:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Paul:2020:AND

[941] Pranesh K. Paul, Babita Kumari, Srishti Gaur, Ashok Mishra, Niranjan Panigrahy, and Rajendra Singh. Application of a newly developed large-scale conceptual hydrological model in simulating streamflow for credibility testing in data scarce condition. Natural Resource Modeling, 33(4):e12283:1–e12283:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mai:2020:LDA

[942] Lina Mai, Qiying Ran, and Haitao Wu. A LMDI decomposition analysis of carbon dioxide emissions from the electric power sector in Northwest China. Natural Resource Modeling, 33(4):e12284:1–e12284:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Batabyal:2020:PEM

[943] Amitrajeet A. Batabyal and Hamid Beladi. A political economy model of the Ganges pollution cleanup problem. Natural Resource Modeling, 33(4):e12285:1–e12285:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Aieb:2020:SMM

[944] Amir Aieb, Khalef Lefsih, Marco Scarpa, Brunella Bonaccorso, Nicola Cicero, Omar Mimeche, and Khodir Madani. Statistical modeling of monthly rainfall variability in Soummam watershed of Algeria, between 1967 and 2018. Natural Resource Modeling, 33(4):e12288:1–e12288:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sihvonen:2020:MLN

[945] Matti Sihvonen, Jussi Lintunen, Elena Valkama, and Kari Hyytiäinen. Management of legacy nutrient stores through nitrogen and phosphorus fertilization, catch crops, and gypsum treatment. Natural Resource Modeling, 33(4):e12289:1–e12289:??, November 2020. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2021:IIa

[946] Anonymous. Issue information. Natural Resource Modeling, 34(1): e12271:1–e12271:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wei:2021:ESI

[947] Yu Wei. Editorial: Special issue for systems analysis in forest resources. Natural Resource Modeling, 34(1):e12302:1–e12302:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Henderson:2021:LHI

[948] Eric B. Henderson and Howard M. Hoganson. A learning heuristic for integrating spatial and temporal detail in forest planning. Natural Resource Modeling, 34(1):e12299:1–e12299:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

She:2021:MRR

[949] Ji She, Woodam Chung, and Hector Vergara. Multiobjective record-to-record travel metaheuristic method for solving forest supply chain management problems with economic and environmental objectives. Natural Resource Modeling, 34(1):e12256:1–e12256:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bushaj:2021:OSM

[950] Sabah Bushaj, İ Esra Büyüktahtakın, Denys Yemshanov, and Robert G. Haight. Optimizing surveillance and management of emerald ash borer in urban environments. Natural Resource Modeling, 34(1):e12267:1– e12267:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Scholz:2021:SFF

[951] Johannes Scholz, Florian Breitwieser, and Peter Mandl. Simulating the forest fuel market as a socio-ecological system with spatial agent-based methods: a case study in Carinthia, Austria. Natural Resource Modeling, 34(1):e12291:1–e12291:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wei:2021:CCF

[952] Yu Wei, Matthew P. Thompson, Erin Belval, Benjamin Gannon, David E. Calkin, and Christopher D. O’Connor. Comparing contingency fire containment strategies using simulated random scenarios. Natural Resource Modeling, 34(1):e12295:1–e12295:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yemshanov:2021:PWH

[953] Denys Yemshanov, Robert G. Haight, Rob Rempel, Ning Liu, and Frank H. Koch. Protecting wildlife habitat in managed forest landscapes — how can network connectivity models help? Natural Resource Modeling, 34(1):e12286:1–e12286:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Suksavate:2021:SPS

[954] Warong Suksavate, Yu Wei, and John Lundquist. Studying the probability of spruce beetle caused mortality in Colorado’s spruce forests using Bayesian hierarchical models. Natural Resource Modeling, 34(1): e12290:1–e12290:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xiao:2021:IRL

[955] Hui Xiao, Caiwang Ning, Fangting Xie, Xiaolan Kang, and Shubin Zhu. Influence of rural labor migration behavior on the transfer of forestland. Natural Resource Modeling, 34(1):e12293:1–e12293:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Akpalu:2021:CCE

[956] Wisdom Akpalu and Jesper Stage. Connectivity at a cost: Economic dynamics of restoring habitat connectivity. Natural Resource Modeling, 34(1):e12294:1–e12294:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Tunnell:2021:CJM

[957] Bolorchimeg N. Tunnell, James A. Conder, Ken B. Anderson, and Marek Locmelis. A cycle-jumping method for multicyclic Hubbert modeling of resource production. Natural Resource Modeling, 34(1):e12296:1– e12296:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2021:E

[958] Anonymous. Erratum. Natural Resource Modeling, 34(1):e12297:1– e12297:??, February 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2021:IIb

[959] Anonymous. Issue information. Natural Resource Modeling, 34(2): e12272:1–e12272:??, May 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sheikh:2021:LUO

[960] Vahedberdi Sheikh, Hossein Salmani, Abdolrassoul Salman Mahiny, Majid Ownegh, and Abolhasan Fathabadi. Land use optimization through bridging multiobjective optimization and multicriteria decision-making models (case study: Tilabad Watershed, Golestan Province, Iran). Natural Resource Modeling, 34(2):e12301:1–e12301:??, May 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Liu:2021:RWS

[961] Yuan Liu. Relationships of wind speed and precipitable water vapor with regional PM_{2.5} based on WRF–Chem model. Natural Resource Modeling, 34(2):e12306:1–e12306:??, May 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cordeiro:2021:ASP

[962] Adria Lorena de Moraes Cordeiro and Claudio José Cavalcante Blanco. Assessment of satellite products for filling rainfall data gaps in the Amazon region. Natural Resource Modeling, 34(2):e12298:1–e12298:??, May 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Verma:2021:EPG

[963] Maitri Verma and Alok Kumar Verma. Effect of plantation of genetically modified trees on the control of atmospheric carbon dioxide: a modeling study. Natural Resource Modeling, 34(2):e12300:1–e12300:??, May 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ranjan:2021:CCT

[964] Ram Ranjan. Can combining two environmental services under a single PES program result in better environmental outcomes and lower costs? Natural Resource Modeling, 34(2):e12303:1–e12303:??, May 2021. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Williams:2021:ABM

[965] Benjamin C. Williams, Keith R. Criddle, and Gordon H. Kruse. An agent-based model to optimize transboundary management for the walleye pollock (Gadus chalcogrammus) fishery in the Gulf of Alaska. Natural Resource Modeling, 34(2):e12305:1–e12305:??, May 2021. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2021:IIc

[966] Anonymous. Issue information. Natural Resource Modeling, 34(3): e12273:1–e12273:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Blackwood:2021:E

[967] Julie Blackwood. Editorial. Natural Resource Modeling, 34(3):e12328:1– e12328:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Duan:2021:MEM

[968] Junyan Duan, Mykhaylo M. Malakhov, Jordan J. Pellett, Ishan S. Phadke, Jackson Barber, and Julie C. Blackwood. Management efficacy in a metapopulation model of white-nose syndrome. Natural Resource Modeling, 34(3):e12304:1–e12304:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Erickson:2021:PCF

[969] Richard A. Erickson, Jessica L. Burnett, Mark T. Wiltermuth, Edward A. Bulliner, and Leslie Hsu. Paths to computational fluency for natural resource educators, researchers, and managers. Natural Resource Modeling, 34(3):e12318:1–e12318:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ackleh:2021:FDE

[970] Azmy S. Ackleh and Amy Veprauskas. Frequency-dependent evolution in a predator–prey system. Natural Resource Modeling, 34(3):e12308:1– e12308:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Grey:2021:OCH

[971] Skylar Grey, Suzanne Lenhart, Frank M. Hilker, and Daniel Franco. Optimal control of harvest timing in discrete population models. Natural Resource Modeling, 34(3):e12321:1–e12321:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Zimmerman:2021:EAT

[972] Mark P. Zimmerman, David M. Chan, Karen M. Kester, Rosalyn C. Rael, and Suzanne L. Robertson. The effects of allelochemical transfer on the dynamics of hosts, parasitoids, and competing hyperparasitoids. Natural Resource Modeling, 34(3):e12311:1–e12311:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Dawed:2021:CHO

[973] Mohammed Y. Dawed and Kiros G. Kebedow. Coexistence and harvesting optimal policy in three species food chain model with general Holling type functional response. Natural Resource Modeling, 34(3):e12316:1– e12316:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hansel:2021:TIE

[974] Martin C. Hänsel and Jeroen C. J. M. van den Bergh. Taxing interacting externalities of ocean acidification, global warming, and eutrophication. Natural Resource Modeling, 34(3):e12317:1–e12317:??, August 2021. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Nurhan:2021:CSS

[975] Yosia I. Nurhan and Shandelle M. Henson. Cannibalism and synchrony in seabird egg-laying behavior. Natural Resource Modeling, 34(3):e12325:1– e12325:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ghonchepour:2021:MFH

[976] Diba Ghonchepour, Amir Sadoddin, Abdolreza Bahremand, Barry Croke, Anthony Jakeman, and Abdolrassoul Salmanmahiny. A methodological framework for the hydrological model selection process in water resource management projects. Natural Resource Modeling, 34(3): e12326:1–e12326:??, August 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2021:IId

[977] Anonymous. Issue information. Natural Resource Modeling, 34(4): e12274:1–e12274:??, November 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gajardo:2021:E

[978] Pedro Gajardo and Héctor Ramírez. Editorial. Natural Resource Modeling, 34(4):e12335:1–e12335:??, November 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

McKelvey:2021:BMI

[979] Steven C. McKelvey, Frank H. Koch, William D. Smith, and Kelly R. Hawley. A Bayesian model identifying locations at risk from humantransported exotic pathogens. Natural Resource Modeling, 34(4): e12307:1–e12307:??, November 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Duffy:2021:MAS

[980] Kevin J. Duffy and Obiora C. Collins. Model analyses show how biodiversity conservation could reduce infectious diseases in an ecosystem. Natural Resource Modeling, 34(4):e12319:1–e12319:??, November 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Vallejos:2021:ESS

[981] Ronny Vallejos and Jonathan Acosta. The effective sample size for multivariate spatial processes with an application to soil contamination. Natural Resource Modeling, 34(4):e12322:1–e12322:??, November 2021. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bhatta:2021:PEI

[982] Dependra Bhatta, Krishna P. Paudel, and Bin Li. Potential economic impacts of groundwater conservation in the Mississippi River Alluvial Aquifer (MRAA), Louisiana, USA. Natural Resource Modeling, 34(4): e12330:1–e12330:??, November 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Campillay-Llanos:2021:FCM

[983] William Campillay-Llanos, Victor Saldaña-Núñez, Fernado Córdova-Lepe, and Felipe N. Moreno-Gómez. Fish catch management strategies: Evaluating the interplay between body size and global warming. Natural Resource Modeling, 34(4):e12331:1–e12331:??, November 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Prieto:2021:CFH

[984] Kernel Prieto and Jhoana P. Romero-Leiton. Current forecast of HIV/ AIDS using Bayesian inference. Natural Resource Modeling, 34(4): e12332:1–e12332:??, November 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gajardo:2021:SRS

[985] Pedro Gajardo, Cristopher Hermosilla, and Athena Picarelli. On the set of robust sustainable thresholds. Natural Resource Modeling, 34(4): e12334:1–e12334:??, November 2021. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2022:IIa

[986] Anonymous. Issue information. Natural Resource Modeling, 35(1): e12312:1–e12312:??, February 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gompil:2022:MCM

[987] Battur Gompil, Batchuluun Tseveen, and Janerke Almasbek. Modeling and control of Mongolian forest utilization: Impact of illegal logging. Natural Resource Modeling, 35(1):e12333:1–e12333:??, February 2022. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Falco:2022:OSM

[988] Carles Falcó and Holly V. Moeller. Optimal spatial management in a multiuse marine habitat: Balancing fisheries and tourism. Natural Resource Modeling, 35(1):e12309:1–e12309:??, February 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yakubu:2022:SAE

[989] Abdul-Aziz Yakubu and Najat Ziyadi. Strong Allee effect and basins of attraction in a discrete-time zoonotic infectious disease model. Natural Resource Modeling, 35(1):e12310:1–e12310:??, February 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Meresa:2022:RIH

[990] Hadush Meresa, Bernhard Tischbein, Josephine Mendela, Rediet Demoz, Tarikua Abreha, Metsihat Weldemichael, and Kingsley Ogbu. The role of input and hydrological parameters uncertainties in extreme hydrological simulations. Natural Resource Modeling, 35(1):e12320:1–e12320:??, February 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bogmans:2022:CGA

[991] Christian W. J. Bogmans and Daan van Soest. Can global aquaculture growth help to conserve wild fish stocks? Theory and empirical analysis. Natural Resource Modeling, 35(1):e12323:1–e12323:??, February 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Jin:2022:RAC

[992] Di Jin, Porter Hoagland, and Andrew D. Ashton. Risk averse choices of managed beach widths under environmental uncertainty. Natural Resource Modeling, 35(1):e12324:1–e12324:??, February 2022. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Liu:2022:FTB

[993] Xiaojin Liu, Longjunjiang Huang, Juan Du, Fangting Xie, and Shubin Zhu. Forestland transfer between rural households in Jiangxi, China: Differentiated effects of actual and perceived tenure security. Natural Resource Modeling, 35(1):e12327:1–e12327:??, February 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mu:2022:CSA

[994] Xianzhong Mu, Li Kong, Chuang Tu, Jian Chen, and Guangwen Hu. Correlation and synergy analysis of urban economy–energy–environment system — a case study of Beijing. Natural Resource Modeling, 35(1): e12329:1–e12329:??, February 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Voss:2022:FMT

[995] Rudi Voss and Martin Quaas. Fisheries management and tipping points: Seeking optimal management of Eastern Baltic cod under conditions of uncertainty about the future productivity regime. Natural Resource Modeling, 35(1):e12336:1–e12336:??, February 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Baghizadeh:2022:SAS

[996] Komeyl Baghizadeh, Naoufel Cheikhrouhou, Kannan Govindan, and Mahboubeh Ziyarati. Sustainable agriculture supply chain network design considering water–energy–food nexus using queuing system: a hybrid robust possibilistic programming. Natural Resource Modeling, 35 (1):e12337:1–e12337:??, February 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2022:IIb

[997] Anonymous. Issue information. Natural Resource Modeling, 35(2): e12313:1–e12313:??, May 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Goshu:2022:MMC

[998] Masitawal Demsie Goshu and Mehari Fentahun Endalew. Mathematical modeling on conservation of depleted forestry resources. Natural Resource Modeling, 35(2):e12338:1–e12338:??, May 2022. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Saeedi:2022:SSM

[999] Iman Saeedi, Ali Reza Mikaeili Tabrizi, Abdolreza Bahremand, and Abdolrassoul Salmanmahiny. A soft systems methodology and interpretive structural modeling framework for Green infrastructure development to control runoff in Tehran metropolis. Natural Resource Modeling, 35(2): e12339:1–e12339:??, May 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

vanLangevelde:2022:MVM

[1000] Frank van Langevelde, Lennart Suselbeek, and Joel S. Brown. Modeling vigilance in mixed-species groups. Natural Resource Modeling, 35(2): e12340:1–e12340:??, May 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2022:IIc

[1001] Anonymous. Issue information. Natural Resource Modeling, 35(3): e12314:1–e12314:??, August 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Wu:2022:MMT

[1002] Taosuo Wu, Hongmei Bai, Feng Feng, and Qian Lin. Multi-month timelag effects of regional vegetation responses to precipitation in arid and semi-arid grassland: a case study of Hulunbuir, Inner Mongolia. Natural Resource Modeling, 35(3):e12342:1–e12342:??, August 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mohammadlou:2022:OEG

[1003] Mohammad Mohammadlou, Abdolreza Bahremand, Daniel Princz, Nicholas Kinar, Amin Haghnegahdar, and Saman Razavi. Objective evaluation of the Global Environmental Multiscale Model (GEM) with precipitation and temperature for Iran. Natural Resource Modeling, 35 (3):e12343:1–e12343:??, August 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Xiong:2022:SIE

[1004] Hongbin Xiong, Tianxin Liu, Haiyun Wang, and Chenxiao Feng. Simulation of the improving effect of graphene visible-light photocatalysis using the MIKE11 model of an urban landscape river in the Chaohu Lake Basin, China. Natural Resource Modeling, 35(3):e12344:1–e12344:??, August 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2022:IId

[1005] Anonymous. Issue information. Natural Resource Modeling, 35(4): e12315:1–e12315:??, November 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Meier:2022:E

[1006] Felix Meier and Hanna Schenk. Editorial. Natural Resource Modeling, 35(4):e12359:1–e12359:??, November 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Blomquist:2022:JMM

[1007] Johan Blomquist, Frank Jensen, Staffan Waldo, Ola Flaaten, and Maija K. Holma. Joint management of marine mammals and a fish species: The case of cod and grey seals in the Nordic–Baltic Sea countries. Natural Resource Modeling, 35(4):e12341:1–e12341:??, November 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Stecher:2022:SMS

[1008] Michael Stecher and Stefan Baumgärtner. A stylized model of stochastic ecosystems with alternative stable states. Natural Resource Modeling, 35(4):e12345:1–e12345:??, November 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Riekhof:2022:TTW

[1009] Marie-Catherine Riekhof, Lotta Clara Kluger, Renato Salvatteci, Lotta Siebert, and Rudi Voss. To tip or not to tip: The Window of Tipping Point Analysis for social-ecological systems. Natural Resource Modeling, 35(4):e12357:1–e12357:??, November 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kourantidou:2022:MEF

[1010] Melina Kourantidou and Di Jin. Mesopelagic–epipelagic fish nexus in viability and feasibility of commercial-scale mesopelagic fisheries. Natural Resource Modeling, 35(4):e12350:1–e12350:??, November 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sun:2022:OCA

[1011] Changyou Sun, Bin Mei, and Yanshu Li. Optimal contract arrangements for conservation on working forests. Natural Resource Modeling, 35(4): e12351:1–e12351:??, November 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Audinot:2022:MPI

[1012] Timothée Audinot, Holger Wernsdörfer, Gilles Le Moguédec, and Jean-Daniel Bontemps. Modeling and propagating inventory-based sampling uncertainty in the large-scale forest demographic model “MARGOT”. Natural Resource Modeling, 35(4):e12352:1–e12352:??, November 2022. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2023:IIa

[1013] Anonymous. Issue information. Natural Resource Modeling, 36(1): e12346:1–e12346:??, February 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Toma:2023:HFD

[1014] Meseret B. Toma, Mulugeta D. Belete, and Mihret D. Ulsido. Historical and future dynamics of land use land cover and its drivers in Ajora-Woybo watershed, Omo-Gibe basin, Ethiopia. Natural Resource Modeling, 36(1):e12353:1–e12353:??, February 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hashem:2023:PRE

[1015] Israa Hashem and Walid Marrouch. Polluting resource extraction and climate risk. Natural Resource Modeling, 36(1):e12354:1–e12354:??, February 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Upmann:2023:SOS

[1016] Thorsten Upmann and Dmitry Gromov. The structure of optimal solutions for harvesting a renewable resource. Natural Resource Modeling, 36(1):e12355:1–e12355:??, February 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Cho:2023:UDB

[1017] Seong-Hoon Cho, James C. Mingie, Nawon Kang, Gengping Zhu, and Sreedhar Upendram. Understanding the differences between single- and multiobjective optimization for the conservation of multiple species. Natural Resource Modeling, 36(1):e12356:1–e12356:??, February 2023. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lin:2023:RRM

[1018] Chao Lin and Lan Gao. Reserve regulation and multidimensional relative poverty of farmers: Evidence from the Panda Nature Reserves in China. Natural Resource Modeling, 36(1):e12358:1–e12358:??, February 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Singh:2023:MRE

[1019] Rajesh Singh and Quinn Weninger. A model of rational equilibrium in quota-regulated multiple-species fisheries. Natural Resource Modeling, 36(1):e12361:1–e12361:??, February 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2023:E

[1020] Anonymous. Erratum. Natural Resource Modeling, 36(1):e12360:1– e12360:??, February 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2023:IIb

[1021] Anonymous. Issue information. Natural Resource Modeling, 36(2): e12347:1–e12347:??, May 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Farzi:2023:STZ

[1022] Parisa Farzi, Seyed Hamidreza Sadeghi, and Mahmoud Jomehpour. Spatial and temporal zoning of watershed resilience using a multidimensional composition approach. Natural Resource Modeling, 36(2): e12362:1–e12362:??, May 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bergland:2023:TDP

[1023] Harald Bergland, Purnedu Mishra, Pål Andreas Pedersen, Arkadi Ponossov, and John Wyller. Time delays and pollution in an open-access fishery. Natural Resource Modeling, 36(2):e12363:1–e12363:??, May 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ghonchepour:2023:PAA

[1024] Diba Ghonchepour, Abdolreza Bahremand, and Nicholas Kinar. Parameter allocation approach for runoff simulation in an arid catchment using the KINEROS2 hydrological model. Natural Resource Modeling, 36(2): e12364:1–e12364:??, May 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Villavicencio-Pulido:2023:CSS

[1025] J. G. Villavicencio-Pulido, V. Vázquez-Hipólito, and G. J. García-Cruz. Catastrophic or sustainable scenarios might occur when the carrying capacities of a tourism-based socioecological system vary. Natural Resource Modeling, 36(2):e12365:1–e12365:??, May 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Zhang:2023:AMR

[1026] Xingyuan Zhang, Fawen Li, and Ximin Yuan. Assessment model of rainwater resource utilization and influencing factors in arid and semiarid areas. Natural Resource Modeling, 36(2):e12366:1–e12366:??, May 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ni:2023:GGH

[1027] Yuanming Ni, Leif K. Sandal, and Sturla F. Kvamsdal. Greed is good: Heuristic adaptations for resilience in renewable resource management.

Natural Resource Modeling, 36(2):e12367:1–e12367:??, May 2023. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2023:IIc

[1028] Anonymous. Issue information. Natural Resource Modeling, 36(3): e12348:1–e12348:??, August 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gebre:2023:OBI

[1029] Agegnehu M. Gebre, Mulugeta D. Belete, and Moltot Z. Belayneh. Object-based image analysis (OBIA)-based gully erosion dynamics, sediment loading rate and sediment yield study in Lake Hawassa Sub-basin, Ethiopia. Natural Resource Modeling, 36(3):e12368:1–e12368:??, August 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Shen:2023:DEF

[1030] Yongchang Shen and Shujing Yue. Does ecological footprint affect biocapacity? Evidence from the experiences of G20 countries. Natural Resource Modeling, 36(3):e12369:1–e12369:??, August 2023. CODEN NR-MOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Batabyal:2023:CCR

[1031] Amitrajeet A. Batabyal, Karima Kourtit, and Peter Nijkamp. Climate change and river water pollution: an application to the Ganges in Kanpur. Natural Resource Modeling, 36(3):e12370:1–e12370:??, August 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Sadeghi:2023:WHE

[1032] Seyed Hamidreza Sadeghi, Mostafa Zabihi Silabi, Hossein Sarvi Sadrabad, Mohammadreza Riahi, and Sedigheh Modarresi Tabatabaei. Watershed health and ecological security modeling using anthropogenic, hydrologic, and climatic factors. Natural Resource Modeling, 36(3): e12371:1–e12371:??, August 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Ali:2023:TIN

[1033] Kishwar Ali, Du Jianguo, Dervis Kirikkaleli, Zoltán Bács, and Judit Oláh. Technological innovation, natural resources, financial inclusion, and environmental degradation in BRI economies. Natural Resource Modeling, 36(3):e12373:1–e12373:??, August 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2023:IId

[1034] Anonymous. Issue information. Natural Resource Modeling, 36(4): e12349:1–e12349:??, November 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Badger:2023:WPO

[1035] Lee Badger. World population outpacing 2003 model predictions. Natural Resource Modeling, 36(4):e12383:1–e12383:??, November 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gars:2023:AOF

[1036] Johan Gars and Daniel Spiro. Approximately optimal forest rotation in a nonstationary environment. Natural Resource Modeling, 36(4):e12372:1– e12372:??, November 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Saeedi:2023:POG

[1037] Iman Saeedi, Ali Reza Mikaeili Tabrizi, Abdolreza Bahremand, and Abdolrassoul Salmanmahiny. Planning and optimization of green infrastructures for stormwater management: The case of Tehran West Bus Terminal. Natural Resource Modeling, 36(4):e12378:1–e12378:??, November 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Mathewos:2023:SEV

[1038] Markos Mathewos, Misgena Tsegaye, and Nigatu Wondrade. Soil erosion variations along land use and land cover dynamics in Matenchose watershed, Rift Valley Basin, Southern Ethiopia. Natural Resource Modeling, 36(4):e12379:1–e12379:??, November 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Khiavi:2023:CAM

[1039] Ali Nasiri Khiavi, Mehdi Vafakhah, and Seyed Hamidreza Sadeghi. Comparative applicability of MCDM-SWOT based techniques for developing integrated watershed management framework. Natural Resource Modeling, 36(4):e12380:1–e12380:??, November 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Heiderman:2023:PNC

[1040] Ryan R. Heiderman and Mark J. Kimsey, Jr. Pacific Northwest conifer forest stand carrying capacity under future climate scenarios. Natural Resource Modeling, 36(4):e12381:1–e12381:??, November 2023. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2024:IIa

[1041] Anonymous. Issue information. Natural Resource Modeling, 37(1): e12374:1–e12374:??, February 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Stigter:2024:CPI

[1042] J. D. Stigter. Computing parameter identifiability and other structural properties for natural resource models. Natural Resource Modeling, 37 (1):e12382:1–e12382:??, February 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Zobel:2024:MEE

[1043] John M. Zobel. Measurement error effects on estimates from linear and nonlinear regression whole-stand yield models. Natural Resource Modeling, 37(1):e12384:1–e12384:??, February 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Chamani:2024:FOW

[1044] Reza Chamani, Seyed Hamidreza Sadeghi, Somaye Zare, Hengameh Shekohideh, Azam Mumzaei, Hamed Amini, Laila Hemmati, and Reza Zarei. Flood-oriented watershed health and ecological security conceptual modeling using pressure, state, and response (PSR) approach for the Sharghonj Watershed, South Khorasan Province, Iran. Natural Resource Modeling, 37(1):e12385:1–e12385:??, February 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Lopes:2024:EFD

[1045] Adrian A. Lopes. Evaluating feedback dynamics between poaching and population with an application to Indian tigers. Natural Resource Modeling, 37(1):e12386:1–e12386:??, February 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Indrajaya:2024:PFC

[1046] Yonky Indrajaya, Hans-Peter Weikard, Frits Mohren, and Edwin van der Werf. Paying for forest carbon: Cost-effectiveness of the Verified Carbon Standard (VCS) remuneration scheme. Natural Resource Modeling, 37 (1):e12387:1–e12387:??, February 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Gillis:2024:MSY

[1047] Darren M. Gillis, Jonah Koscielny, and Benjamin Blanz. Maximum sustainable yield as a reference point in the presence of fishing effort that follows an ideal free distribution. Natural Resource Modeling, 37(1): e12390:1–e12390:??, February 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Anonymous:2024:IIb

[1048] Anonymous. Issue information. Natural Resource Modeling, 37(2): e12375:1–e12375:??, May 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Huber:2024:WPM

[1049] Celine Huber, Luc Doyen, and Sylvie Ferrari. When profitability meets conservation objectives through biodiversity offsets. Natural Resource Modeling, 37(2):e12389:1–e12389:??, May 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Kibira:2024:OCP

[1050] Gerald Kibira, Edwin Muchapondwa, and Herbert Ntuli. The optimal combination of pastoral activities and wildlife conservation in the Serengeti ecosystem. Natural Resource Modeling, 37(2):e12391:1– e12391:??, May 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Hong:2024:HDG

[1051] Jiaxuan Hong and Yanling Mao. How does the green efficiency of urban land use evolve in the urban agglomeration of China’s middle Yangtze River? Natural Resource Modeling, 37(2):e12392:1–e12392:??, May 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Bang:2024:BAS

[1052] Rasmus Noss Bang and Stein Ivar Steinshamn. Beyond age-structured single-species management: Optimal harvest selectivity in the face of predator–prey interactions. Natural Resource Modeling, 37(2):e12393:1– e12393:??, May 2024. CODEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).

Yoshioka:2024:OHP

[1053] Hidekazu Yoshioka. Optimal harvesting policy for biological resources with uncertain heterogeneity for application in fisheries management.

REFERENCES 176 Natural Resource Modeling, 37(2):e12394:1–e12394:??, May 2024. CO-DEN NRMOEU. ISSN 0890-8575 (print), 1939-7445 (electronic).